### v1

Abstract
The deployment of autonomous agents in environments involving human interaction has increasingly raised security concerns. Consequently, understanding the circumstances behind an event becomes critical, requiring the development of capabilities to justify their behaviors to non-expert users. Such explanations are essential in enhancing trustworthiness and safety, acting as a preventive measure against failures, errors, and misunderstandings. Additionally, they contribute to improving communication, bridging the gap between the agent and the user, thereby improving the effectiveness of their interactions. This work presents an accountability and explainability architecture implemented for ROS-based mobile robots. The proposed solution consists of two main components. Firstly, a black box-like element to provide accountability, featuring anti-tampering properties achieved through blockchain technology. Secondly, a component in charge of generating natural language explanations by harnessing the capabilities of Large Language Models (LLMs) over the data contained within the previously mentioned black box. The study evaluates the performance of our solution in three different scenarios, each involving autonomous agent navigation functionalities. This evaluation includes a thorough examination of accountability and explainability metrics, demonstrating the effectiveness of our approach in using accountable data from robot actions to obtain coherent, accurate and understandable explanations, even when facing challenges inherent in the use of autonomous agents in real-world scenarios.
keywords:
[…]
{highlights}
Security improvements for autonomous agents in public spaces.
Review of blockchain trends for immutable logs, EDRs, XARs, and LLMs.
Accountability and Explainability architecture for autonomous agents.
Comprehensive assessment across diverse navigation scenarios and metrics.
Future research directions for accountable and explainable autonomous agents.
1 Introduction
The rising presence of mobile robots alongside humans has increased growing security and safety concerns. One of the main challenges in this context is the ability of robots to justify their actions and decisions, particularly in safety-critical scenarios (Sado et al., 2023). Therefore, it is essential to ensure that the events performed when developing an action or behavior are reliable and transparent (Borgo et al., 2018). Equally important is the capacity to render such explanations comprehensibly to non-expert users, offering appropriate, effective, and natural answers to their queries (Rosenfeld and Richardson, 2019; Cruz et al., 2023).
The most substantial evidence in post-event analysis relates to the logging and auditing capabilities, which are essential to mitigate technical issues and provide forensic evidence to detect any data tampering or unauthorized leakage. Therefore, a trustworthy and secure log system is needed when generating a complete trace of a robot’s behavior (Al-Slais and Ali, 2023). The system must ensure the integrity, confidentiality, and availability of the agent and the data. Additionally, it should possess the capability of incident detection, facilitating investigations, and ensuring compliance, ultimately improving the overall security of the robot.
Regarding log systems vulnerabilities, covering tracks has become a common practice for attackers, erasing any traces that could lead to their identification or assisting in the analysis of their actions. As a result, it is necessary to strengthen system properties against unauthorized modifications (Ahmad et al., 2019). This fact not only includes safeguarding the integrity of log data but also implementing measures to detect and counteract attempts at tampering. Accountability is equally required, enabling detection and isolation of faulty behaviors and their origin. It must not only provide tamper resistance, but also to offer verifiability for checking the integrity of log entries against tampering. Moreover, it should guarantee search accessibility, enabling access to specific records whose integrity can be verified (Chiu et al., 2021).
Integrating accountability solutions into robotic systems often demands a trade-off with system performance. Specifically, features such as computational power and networking bandwidth can be negatively affected. The inclusion of anti-tampering techniques, such as full disk and communication encryption, may be restrictive due to resource limitations (Fernández-Becerra et al., 2022). Additionally, high-cost tamper proof storage devices may prove economically unviable, especially when dealing with continuously generated log data.
Therefore, it becomes necessary to establish an efficient and verifiable proof of integrity to guarantee the reproducibility and identification of the events that could arise during an autonomous agent action. For this evidence to be significant, it should belong to a persistent chain of custody, totally reliable and verifiable, main requirements for an auditable and transparent logging system (Putz et al., 2019; Hajer et al., 2024).
Blockchain solutions emerge as a mechanism to ensure data integrity, protection against unauthorized changes, and traceability. The consensus protocols established by all blockchain network members enable the detection and rejection of any destruction, manipulation, or non-permitted editing of data through illicit channels. The nature of this technology helps to avoid tampering risk, safeguarding data (Iqbal and Matulevičius, 2019) or even .
Nevertheless, the potential obtained evidence poses significant challenges for analysis due to its diverse and varied nature, terminology, and the presence of substantial volumes of similar messages unrelated to anomalous events. These facts may result in limited usefulness and misalignment with the expectations of non-expert users (Shehu and Harper, 2022). Regarding autonomous agents, the raw log messages produced by these systems, are usually noisy and possess semi-structured format. In addition to this, it is necessary to deal with specific features such as large quantities of domain-specific language messages (Ott et al., 2021).
In this particular context, explaining the reasons that make a robot carry out a specific and unexpected action is essential for diagnosing failures in goal achievement. Besides, making these explanations accessible to humans is equally important. This requirement has led to the development of eXplainable Autonomous Robots (XAR), which aims to provide mechanisms for presenting clear and understandable explanations for the actions and decisions of autonomous agents. XAR enhances the transparency and interpretability of these agents’ decision-making processes, often incorporating techniques like natural language explanations to improve clarity (Anjomshoae et al., 2019; Langley et al., 2017).
Large Language Models (LLMs) have gained significant attention due to their exceptional performance in diverse natural language processing tasks, including translation, coding or text summarization (Kocmi and Federmann, 2023; Zhang et al., 2024; Sarsa et al., 2022). One approach to improve the results of large language models for a specific domain is through the use of retrieval-based methods. By retrieving relevant information from a large corpus of text, these models can provide explanations grounded in existing knowledge (Ren et al., 2023).
This work presents an accountability and explainability architecture implemented for ROS-based mobile robots. This solution included two main components. Firstly, a black box-like component to provide accountability that includes anti-tampering features by distributing integrity proofs from the events recorded by the black box component. The second component generates natural language explanations from data in the black box, including the use of interactive natural language conversations as a means to provide comprehensive explainability for the recorded events. Both components are connected through an additional component responsible for processing raw data collected by the first one, eliminating non-essential information and building connections and relations between data. These functionalities release software developers from the need to make calls to the log API to build explanations for the code generated and the behaviors developed by the autonomous agent. Additionally, we have conducted a thorough evaluation of our solution, including its impact on system performance and the correctness, accuracy, helpfulness and coherence of the generated natural language explanations.
The remainder of this paper is organized as follows: Section 2 provides an overview of previous works and findings related to the study, focusing on the enhancement of immutable data recorders and explainability solutions for robots. Section 3 delves into the system architecture, offering a detailed description of the designed approaches. Section 4 details the technical elements of the designed system attending to the proposed architecture. Section 5 discusses the experimental process, the scenarios and the variables employed for assessment. Section 6 outlines the obtained results, particularly addressing accountability and explainability metrics. Finally, Section 7 summarises the obtained conclusions.
2 Related works
In this section, we explore key works related to various domains relevant to our approach. It includes progress in understanding how blockchain ensures the immutability of logs, research on the application of event data recorders in robotics, and significant contributions Explainable Autonomous Robots (XAR) field. Lastly, we examine the evolution of Large Language Models (LLMs) to provide natural language explanations.
2.1 Blockchain for immutable logs
Blockchain is a decentralized solution that eliminates the necessity for trust in a central authority through its immutable and distributed ledger, composed of timestamped transaction blocks. These blocks are linked through hashing and storage of previous block content, ensuring the integrity of transactions. Any attempt to alter a transaction in one block requires the modification of all subsequent blocks, resulting in a computationally expensive process. As a result, data stored in the blockchain is immutable (Rajasekaran et al., 2022; Guo and Yu, 2022).
There are two primary types of blockchains based on access controls: permissionless blockchains allow open access and public transaction visibility, while permissioned blockchains limit participation and transaction visibility. Nevertheless, closed networks may pose security risks, as authorized users could recalculate subsequent blocks in the event of a block modification (Solat et al., 2021). While permissionless blockchains have traditionally related to scalability and throughput challenges, approaches like Ethereum have transitioned to Proof of Stake (PoS), a more energy-efficient, secure, accessible and scalable consensus protocol (Nguyen et al., 2019; Sanka and Cheung, 2021).
Blockchain technology ensures log integrity and auditability, offering unique advantages over alternatives such as distributed storage systems or IPFS (InterPlanetary File System), which may require central authorities for storage and verification (Omar et al., 2024). Blockchain offers consensus and immutability, coupled with the capability of integrating logic mechanisms such as smart contracts (SCs). SCs contain functions that either write values onto the blockchain, thus altering its state, or read data from it (Mohanta et al., 2018). These features make blockchain the preferred choice for applications where data integrity and trust are required (Nyaletey et al., 2019).
Different secure logging systems rely on permissioned blockchain frameworks for log immutability (Shekhtman and Waisbard, 2019), introducing accountability even in embedded systems (Chiu et al., 2021). To get a balance between throughput and cost-effectiveness, some works combine both categories (Zhao et al., 2023), developing approaches that include a permissioned blockchain to ensure per-entry immutability and public non-repudiation solution (Pourmajidi et al., 2023). Nevertheless, inherent features of permissionless blockchain such as transparency and trust, have also boosted the development of secure logging systems (Ge et al., 2019).
2.2 Robotic Event Data Recorders
Event Data Recorders (EDRs), often referred to as black boxes, have traditionally played a crucial role in in capturing and analyzing events within robotics systems. These recorders support the collection and storage of event data, helping to comprehend system behavior, detecting anomalies, and ensuring safety (Winfield et al., 2021, 2022). Therefore, these solutions are needed in forensic investigations, ethical assessments and transparency regarding autonomous agents (Taurer et al., 2018).
The integration of black box modules to enhance robots’ introspection capabilities contributes to improving system transparency by attributing responsibilities (Fernández-Becerra et al., 2022). These solutions typically include functionalities such as hashing, signing, and encryption (Schmidt et al., 2022). However, to ensure the integrity and authenticity of recorded events, anti-tamper features must also be in place (Srinivas Aditya et al., 2021).
Using blockchain technology in event data recorders can guarantee the integrity of a transparent and auditable record of events. These features enhance accountability and trust in the recorded data, as it becomes difficult for any party to manipulate or tamper with the recorded events without detection (Salah et al., 2019). Additionally, blockchain involves decentralized and distributed storage of event data, eliminating the necessity for central authorities or intermediaries and reducing the risk of data loss.
Further research explores the application of blockchain technology to improve robotic capabilities, particularly event data recording and management. The Black Block Recorder (BBR) system proposed in (White et al., 2019), employs permissioned blockchain technology for immutable black box logging in robotic applications. The BBR system combines Rosbag2 and distributed ledger technologies (DLTs) to ensure data integrity and immutability. The potential applications of blockchain in robot swarms, including secure communication, data logging, and consensus agreement are discussed in (Strobel et al., 2020). In (Lopes et al., 2021), a time-segmented consortium blockchain for robotic event registration is proposed, organizing and storing event records efficiently to allow for retrieval and analysis of specific events within specific time frames. Authors in (Zhang et al., 2022) present a tool to bridge the Robot Operating System (ROS) and the Ethereum blockchain, evaluating encryption efficiency and stability, transactions response time and packet loss rate. Nevertheless, these works lack capabilities to save integrity proofs in the blockchain at selected intervals, to enhance granularity, or including service definitions or smart contracts to improve the verifiability process.
2.3 Explainable Autonomous Robots
The need for explainability in autonomous robots arises from increasing complexity and autonomy of these systems, making it challenging for users to comprehend their behavior. Explainable Autonomous Robots (XARs) are designed to provide explanations for their actions and decisions to non-expert users. Such explanations enrich aspects like security and the trust level, preventing failures, errors, or conflicts in achieving objectives or undesired interactions (Han et al., 2019).
Despite the growing demand for explainability in autonomous agents, their use reveals a lack of practical and effective usability or interpretability in real environments (Abdul et al., 2018). Therefore, it is a problem that requires special attention, being necessary to explain the performed decisions. This explanation should include alternatives, choices made, expectations of each option, decision criteria, and any unexpected events that may have led to the modification of the initially planned behavior (Langley et al., 2017; Sadeghi et al., 2021).
As described in (Sakai and Nagai, 2022), an explanation regarding XAR, is an "abductive reasoning that bridges the gap between the individual explaining, which in this case may be a robot, and the person to whom they are explaining". Self-explanatory capabilities in social robots have been provided through the development of architectures and frameworks (Stange et al., 2022; Han et al., 2021).
Nevertheless, the essence of explainability regarding XAR lies in effective communication (Sakai and Nagai, 2022). Progress in natural language processing (NLP) technology through large language models (LLMs), based on the estimation of human statements and the relevance and position of the words contained in them, could improve the efficiency of human-robot interactions. This enhancement can create a sense that these systems truly comprehend and respond to individuals meaningfully, significantly benefiting their usefulness and applicability (Sado et al., 2023).
2.4 Large Language Models
Large Language Models (LLMs) have recently played a significant role in natural language processing, thanks to replacing recurrent neural networks with self-attention mechanisms (Vaswani et al., 2017). The design and implementation of the first Transformers models constitute the foundation of most LLM architectures. Trained on massive text data, LLMs include strong capabilities in natural language understanding and complex task solving (Zhao et al., 2023).
However, LLMs face two primary challenges. First, they tend to produce a significant amount of hallucinations due to their reliance on internal knowledge (weights) (Andriopoulos and Pouwelse, 2023). Second, their large parameter sizes and the associated high update costs, make traditional pre-training and fine-tuning methods unfeasible. As a solution, Retrieval-Augmented Generation (RAG) methods emerge as a solution for LLMs to effectively interact with the external world (Zhao et al., 2023). This strategy is particularly notable in question-answering systems, which represent the most widely adopted use cases for harnessing explanatory and reasoning capabilities of large language models. In this domain, RAG-based approaches are emerging as the standard for developing solutions that seamlessly integrate the advantages of a knowledge base (utilizing a vector store) and generative models.
LLMs display emergent abilities, such as in-context learning (ICL), where LLMs make predictions based on contexts augmented with a few examples, helping the model to learn from analogy (Wei et al., 2022). This paradigm simplifies inclusion of human knowledge into LLMs by modifying the prompt templates (Dong et al., 2022). ICL does not require altering the model’s parameters or training the model on a specific dataset; instead, it relies on guiding the LLM through the prompting interface to align its results with human values and preferences.
Nevertheless, evaluating large language models (LLMs) and the impact of previously described approaches, poses several challenges due to their complexity and the nature of their outputs. These difficulties can be attributed to several factors, including the lack of suitable evaluation metrics, biases in generated text, and the cost of human assessment (Chang et al., 2024). Nevertheless, recently, employing LLMs as evaluators for LLMs-based solutions has been proven to be nearly as effective as human judges (Zheng et al., 2023).
Previous works have demonstrated the potential of using LLMs to generate natural language descriptions of logging systems (Setianto et al., 2022) or robots’ behaviors (González-Santamarta et al., 2023), indicating room for improvements to deliver more accurate and complete interpretations. Nevertheless, to our knowledge, no previous research has investigated the capabilities of LLMs in generating natural language explanations within the context of an accountability approach for record interpretation.
3 System architecture
Our approach comprises to main components: a tamper-proof black box-like device, and a module responsible for generating natural language explanations for the data stored by the former. The first component offers services for building, storing, and verifying integrity proofs through smart contract specifications. The explainability component provides natural language explanations by implementing Retrieval-Augmented Generation (RAG), thereby enhancing the obtained results using the data recorded by the first component as an external source to enrich the context in the answer formulation process. Both components are integrated through the existence of two main asynchronous tasks focused on cleaning and processing the raw data generated by the black box-like device to increase the effectiveness of the explainability engine.
Figure 1 depicts a generic architecture of our proposal, highlighting in blue color the functionalities related to accountability, and in green those related to explainability. Below we provide a deeper description of the approaches designed and followed in order to develop and integrate the above mentioned components.
Figure 1: System architecture.
3.1 Black Box Recorder
Our first component focuses on recording accountable information that captures essential messages generated during the navigation task of an autonomous agent. The selected data will guarantee the reproducibility of the mentioned task, including details like the robot’s poses, the employed map, odometry data, linear and angular velocity, and laser scan readings.
The immutability of the previously described information is achieved through blockchain. Blockchain technology usually faces limitations regarding the size of the data that can be written into a single transaction. Efficiently storing autonomous agent messages becomes challenging due to these constraints. Our proposed solution involves selectively storing the hash of messages at intervals, providing in addition the chance of using different storage frequencies depending on the topic and its relevance. This fact not only enhances the granularity of stored message proofs but also provides a more adaptable solution.
Our anti-tampering approach for this accountability proposal comprises two key components: a chain of integrity proofs coupled with the smart contract specification. This section delves into their design and justifies how both align with the constraints of autonomous agents.
Key features such as consistency are guaranteed by the blockchain’s consensus protocol, where the content of each block is hashed and stored in the next one. Additionally, availability is also assured through the distributed and decentralized nature of the blockchain, ensuring accessibility across the entire network.
3.1.1 Chain of integrity proofs
To preserve the integrity of the messages without compromising system performance, our solution entails building a chain where each message hash is linked with the previous hash value. This approach ensures that any alteration of the selected messages will result in a different final hash, easing tampering detection. The verification process is also simplified since only the last hash in the chain needs to be recomputed and compared with the last stored value. A single modification of a hash will disrupt the entire chain from that point onward, simplifying the identification of tampering.
We define a chain entry ($CE$) with a counter $i\in 1..m$ to be linked with the previous one by concatenating the prior digest $h_{i-1}$ with the current log message $Msg_{i}$. hash() denotes the cryptographic hash function, with SHA-256 being the chosen one, commonly used for data integrity verification and digital signatures.
The initial hash value $h_{0}$ of the chain is calculated by applying the selected hash function to a set of $l=32$ random bytes ($RB$), to generate a 256-bit value. Introducing initial entropy into the chain of integrity proofs enhances privacy, ensuring that integrity proofs for similar initial log messages from different records are not identical. Data privacy is further prioritized by storing in the blockchain only the hashes, rather than the contents of the messages.
$CE_{i}$ $=h_{i}$ (1)
$\text{where}~{}RB$$\leftarrow\{0..F\}^{l},$
$h_{0}$ $\leftarrow\text{hash}(RB),$
$h_{i}$ $=\text{hash}(h_{i-1}||Msg_{i})$
By chaining every hash with the previous one, we optimize the validation of stored messages due the last linked digest should correspond to the final proof saved in the blockchain. Additionally, a service has been developed to generate the chain of hashes using the recorded data, ensuring partial validation of the integrity proofs without modifying the stored messages or adding non-essential information to the transactions saved in the blockchain.
The number of proofs included in a blockchain transaction depends on the transaction cost limit set by the blockchain framework. Therefore, being $j\in 1..m$ a transaction counter, each newly arriving chain entry is added in a transaction payload $P_{j}$ until the limit is reached. Being sign() a shorthand for a signature function and $PrivKey$ the private key of the account, transaction $T_{j}$ is transmitted from the service in charge of recording the accountable data to the blockchain.
$T_{j}$ $=(P_{j},S_{j})$ (2)
$\text{where}~{}P_{j}$$\leftarrow\{CE_{i},CE_{i+1},CE_{i+2},\dots\},$
$S_{j}$ $=\text{sign}(P_{j},PrivKey)$
A service interruption or a disruption in the behavior of the autonomous agent will lead to the aggregation of the remaining hash values into a single transaction, even if there is capacity for additional values. This practice guarantees the immutability of the last messages sent by a robot upon completing a task or achieving a goal.
3.1.2 Smart contract
Calling a smart contract function intended for blockchain writing involves the execution of a transaction that changes the state of the storage data on the blockchain. In our approach, this transaction is signed with the account’s private key via the Elliptic Curve Digital Signature Algorithm (ECDSA), effectively identifying the signer.
The defined smart contract also determines its ownership, implementing access control to ensure that only the contract owner can add or update message proofs. These features guarantee authentication and non-repudiation.
Moreover, our smart contract includes a function to verify the existence of a message proof on the blockchain. This fact allows external parties to independently confirm whether a specific message hash has been stored in the contract and when it was stored. This functionality contributes to the system’s transparency, enabling users or auditors to check the validity of message proofs without relying solely on the contract owner, thereby promoting public accountability.
3.2 Explainability Component
Effective communication is essential for achieving explainability in autonomous robots. Our approach is focused on improving this feature through the use of Natural Language Processing (NLP) technology, enabling non-expert users to interact with agents in a meaningful way.
Our solution involves use of RAG to retrieve data from outside the language model (non-parametric) and augment the prompts by adding the information recorded from our accountability component in context. Before this integration, the information undergoes processing by a dedicated module responsible for acquiring and processing the raw data generated by the autonomous agent, serving as contextual input in the question-answering system.
Given the volume of messages, their varied nature, their length, the presence of non-meaningful records generated by the autonomous agent, and the need of establishing specific relationships between messages to enhance the context of the engine, this component extracts relevant information. Its aim is to improve the effectiveness of the explanations provided to non-expert users.
After processing the raw data stored by our accountability component, the ingestion process performs the chunking task to break down extensive text segments into more manageable units. To ensure comparability during relevancy searches at inference time, both this data and the user query are converted into numerical representations through embedding language models. Next, based on the embedding of the user query, its relevant text is identified in the document collection by a similarity search in the embedding space. The user-provided prompt is then extended with relevant information obtained in the search and added to the context. This enhanced prompt is sent to the LLM. As the context contains pertinent external data along with the original prompt, the model output is relevant and accurate.
The addition of new accountable data to our explainability component requires asynchronous updates of the accountable data and its embedding representation. These updates ensure that the model remains sensitive to changes, addressing limitations introduced by offline training.
4 Implementation
The solution presented in Section 3 can be implemented using different technologies. This section details the specific solutions chosen for developing our proposal.
4.1 Black Box Recorder
Robot Operating System (ROS) is the most commonly used framework in robotics, consisting of a set of software libraries and tools for building robot applications. However, challenges in meeting the requirements of real-time embedded systems, and the need to improve its multi-platform capabilities led to a major upgrade to ROS 2, whose main change lies in using the Data Distribution Service (DDS) (Maruyama et al., 2016).
Nodes constitute the fundamental building block in both ROS and ROS 2 applications. They consist of individual processes, each performing specific tasks and communicating with each other via messages. Data can be moved between nodes using different methods, including topics, services, and actions. Topics are used for message passing, while services provide synchronous request-reply interactions. Actions, on the other hand, are needed for long-running tasks that require feedback.
This work focuses on the navigation functionality. To this end, we have used Nav2, a collection of tools for ROS 2, which is considered the successor of the ROS Navigation Stack. Nav2 provides the robot with the ability to perform complex navigation tasks.
Its main functionalities include planning, control, localization, visualization and other additional features. Furthermore, it uses behavior trees to create customized navigation behavior via orchestrating many independent modular servers, which implies high flexibility in the navigation task and the specification of complex robot behaviors (Macenski et al., 2020). Different independent task servers, intended to compute a path, control effort, or recovery, are communicated with the behavior tree (BT) over a ROS 2 interface, such as an action server or service. The use of BTs to generate explanations for robot behavior has been previously successfully evaluated, allowing seeking causal information answer questions (Han et al., 2021).
Previous authors’ work describes the steps involved in the accountable information recording process (Fernández-Becerra et al., 2023). ROS 2 nodes publish data that other nodes consume through ROS 2 topics. Such data include perception information from sensors or commands to actuators from decision-making nodes. Navigation reproducibility of autonomous robots in ROS 2 involves storing essential data, such as their poses, maps, odometry, velocity, and laser scan readings. To this aim, we have used Rosbag files, a binary file format designed to capture and store data generated during the execution of a ROS system. These files can reproduce the exact conditions and inputs from a previous run, making them particularly valuable for understanding system behavior. The set of topics included in the Rosbag files generated are detailed in Table 1.
Table 1: Topics name included in the Rosbag file their description.
Topic name 1 Information provided
map 2-D grid map, in which each cell represents the probability of occupancy.
tf_static Static transform information; fixed relationships between coordinate frames over time.
robot_description URDF (Unified Robot Description Format) or robot model description.
navigate_to_pose/_action/status Information about the current state or progress of the navigation action.
global_costmap/costmap Representation of the environment in terms of costs associated with different areas.
plan Representation of the planned path for the robot to follow as a sequence of waypoints.
rosout Logs and messages from ROS nodes.
local_costmap/costmap Representation of the immediate surroundings of the robot and their cost.
amcl_pose Position and orientation of the robot.
behavior_tree_log Nav2 Behavior Tree nodes status.
cmd_vel Linear and angular velocity commands to the robot’s motion control system.
camera/image_raw Raw image data captured by a camera sensor on the robot.
scan Distances from the sensor to obstacles in the robot’s surroundings.
odom Estimated motion and position of a robot based on sensor data.
tf Coordinate frame transformations in the robotic system.
To ensure the immutability of generated Rosbag files, Ethereum has been selected as the blockchain framework. Ethereum, an opensource public blockchain, introduces the cryptocurrency Ether and a programming platform known as Smart Contract. Employing Proof of Stake (PoS) as a consensus protocol, Ethereum offers advantages in energy efficiency, scalability, security, decentralization, accessibility, and economic efficiency. The high-level algorithms 1 and 2 depict the main steps in recording and proof verification services.
Algorithm 1 Accountable information recording
1:topicsSet = [$selectedTopics$], $serviceCallValue$, $contractAddress$, $contractAbi$, BLOCKS_LIMIT, $privateKey$
2:$rosbag$
3:Boolean $recording\leftarrow$ false
4:String $initialNonce\leftarrow$ sha256($randomBytes$)
5:// Recording process attending to a service call.
6:$recording\leftarrow serviceCallValue$
7:if $recording$ = $true$ then
8:     String $previousHash\leftarrow$$initialNonce$
9:     for $message$ in $topicsSet$ do
10:         if $messageNumber$ $topicRate$ == 0  then
11:              $chainedHash\leftarrow$ sha256($message$,$previousHash$)
12:              $previousHash\leftarrow$ $chainedHash$
13:         end if
14:         // Building chain of hashes.
15:         if $payloadSize$ < BLOCKS_LIMIT then
16:              $payload\leftarrow$ $payload\cup chainedHash$
17:         else
18:              // Load Solitidy contract.
19:              loadBcContract($contractAddress$, $contractAbi$)
20:              // Saving chain of hashes into blockchain.
21:              $transaction\leftarrow$buildTransaction($payload$)
22:              $signedTransaction\leftarrow$ signTransaction($transaction$, $privateKey$)
23:              sendTransaction($signed\_transaction$)
24:              for $hash$ in $payload$ do
[…]
26:              end for
27:         end if
28:     end for
29:end if
30:return $rosbag$
Algorithm 2 Proof verification
1:$hashValue$
2:$BlockNumber$
3:// Load Solitidy contract.
4:loadBcContract($contractAddress$, $contractAbi$)
5:$transaction\leftarrow$buildTransaction($hashValue$)
6:$signedTransaction\leftarrow$ signTransaction($transaction$, $privateKey$)
7:sendTransaction($signed\_transaction$)
8:$blockNumber\leftarrow$readBC($hashValue$)
9:if $blockNumber$<>0 then
10:     $msg\leftarrow$ “The hash value is stored in $blockNumber$."
11:else
12:     $msg\leftarrow$ “The hash value is not stored."
13:end if
14:return $msg$
Our approach addresses challenges such as the drop in robot performance during the Rosbag recording process and costs associated with storing blockchain data. We accomplish this by saving chained hashes from messages at different intervals. These intervals can be established based on the relevance of messages from a topic, regarding their required immutability features. Furthermore, message proofs are stored in real-time on the Ethereum blockchain, eliminating the need to modify the content of the Rosbag file. Detection of any alteration to the messages within this file is facilitated by a service that generates the hash chain from the file at intervals set for each topic during the recording process. To verify the existence of any of these hashes in the blockchain, the blockchain logic has been extended by developing a Smart Contract (SC) using the Solidity language. Solidity is an open-source programming language that facilitates the creation of decentralized applications, contracts, protocols, and agreements. This contract provides functionalities for storing and verifying proof information for hashes stored on the Ethereum blockchain by associating each hash value with the number of block where it is stored. Verification of transaction ownership is also conducted during the storage of these proofs.
Therefore, this approach combines immutability, decentralization, traceability, and cost efficiency to address challenges related to the real-time recording and storage of Rosbag files securely and transparently.
4.2 Log Curation Component
To filter relevant information from the Rosbag file messages, a raw data processing node has been implemented. Given the size, variety, and intricate nature of messages contained in the Rosbag file, the inclusion of this raw information might be challenging for non-expert users. Therefore, it is necessary to include a component that processes this data, eliminating non-essential content in the natural language explanation generation process. By analyzing messages from the recorded Rosbag topics, this service can identify key circumstances in the navigation process, such as the appearance of an obstacle or a change in the planned trajectory. As a result, through this component, ROS 2 developers are relieved from having to call the logging API to generate explanations. This functionality streamlines the development process, allowing developers to focus on core functionalities without need for managing log calls. This approach reduces debugging efforts and boosts efficient resource usage, scalability, and adaptability in the development lifecycle, contributing to a more productive natural language explanation generation process. An example of how raw ROS messages are transformed through this component is included in Figure 2.
Rosbag Raw Messages action_msgs.msg.GoalStatusArray(status_list=action_msgs.msg.GoalStatus(goal_info= action_msgs.msg.GoalInfo( goal_id=unique_identifier_msgs.msg.UUID(uuid=array([190, 81, 48, 208, 248, 147, 78, 82, 188, 116, 202, 152, 203, 225, 149, 158], dtype=uint8)), stamp=builtin_interfaces.msg.Time(sec=27, nanosec=933000000)), status=2)]) nav2_msgs.msg.BehaviorTreeStatusChange(timestamp=builtin_interfaces.msg.Time(sec=1701354686, nanosec=130104968), node_name=’ComputePathToPose’, previous_status=’IDLE’, current_status=’RUNNING’) Processed Data 1701354686 Navigation to the goal number 1 has started. 1701354686 Navigation to the goal number 1 is in progress. 1701354686 Nav2 Behavior Tree node NavigateRecovery that recovers from unexpected situations, is running.
Figure 2: Example of raw Rosbag messages and information obtained from them after the Log Curation process.
4.3 Explainability Component
The output of the above described node serves as an external source in our implementation of the RAG approach. This development is based on the use and extension of localGPT  [^1] https://github.com/PromtEngineer/localGPT, an open-source question-answering system designed to interact with documents conversationally while safeguarding the privacy of the underlying information. Running locally, it ensures data privacy and secure interactions with the interpreted and recorded data, facilitating the generation of natural language explanations. Its functionality involves loading a set of selected documents, breaking them into smaller chunks, augmenting search capabilities with LangChain tools, and generating local embeddings. The outcomes are stored in a local vector database using Chroma vector storage. Once this information is saved, and given a user input, relevant splits are retrieved from storage via a similarity search, which finds the appropriate piece of information from the documents. The retrieved data is used by the selected LLM when producing an answer. Figure 3 depicts the main components in the RAG architecture used.
Figure 3: Retrieval Augmented Generation arquitecture.
Therefore, this implementation combines the strengths of robotics frameworks, blockchain technology, and natural language processing to provide a comprehensive solution for real-time recording, secure storage, and user-friendly interpretation of robotic system data. It addresses challenges related to performance, storage costs, and user comprehension, making it an effective approach when understanding the behavior of autonomous robots.
4.4 Software and hardware artifacts
Our solution has been deployed in a Linux server running Ubuntu 22.04 as Operating System (OS). The server is equipped with 6-core 2.6 GHz CPU, 64 GB of memory, and a 4 GB GPU.
The chosen ROS 2 version is Humble. Details regarding the topics incorporated into the Rosbag file, along with their average rates, are provided in Table 2. These topics play a crucial role in ensuring the reproducibility of the navigation process, facilitating data analysis and debugging purposes.
The experimentation was executed entirely in Gazebo, an open-source solution for 3D robotics simulations. To provide a realistic scenario, we used a hospital simulation environment made available by Amazon Web Services (AWS) [^2] https://github.com/aws-robotics/aws-robomaker-hospital-world. Additionally, we selected RB-1, an autonomous mobile robot based on the ROS 2 platform produced by Robotnik (Guzmán et al., 2017). This combination of tools and platforms enabled us to assess the impact and effectiveness of our approach.
The interaction with Ethereum has been developed through web3.py library (version 6.6.1). Ganache (version 2.7.1) has been used as the Ethereum network simulator.
RAG solution holds Langchain (version 0.325), ChromaDB (version 0.4.6). We used the open source model Llama 2 7B GGUF (Touvron et al., 2023) as the LLM responsible for generating natural language explanations for end users. This LLM integrates RLHF and shows versatility in handling general-purpose tasks, ranging from answering questions, text generation, and language translation to more demanding tasks such as reasoning or coding. Preliminary tests of Llama 2, compared to other LLMs such as Mistral (Jiang et al., 2023) or Zephyr (Tunstall et al., 2023), led us to choose this model for evaluating our data, due to the promising results obtained in these assessments conducted during the last quarter of 2023, when this work was in its development phase. As text embedding model we have chosen Instructor-xl (Su et al., 2022), a high-performance model designed to generate text embeddings tailored to tasks such as classification, clustering, text evaluation or retrieval, fact particularly important in the processes performed for any RAG solution.
Table 2: Robag topics rate and average rate per message
Topic Avg. Rate (Hz) Avg. Size Msg. (B)
map 0 1007308.8
tf_static 0 9011.2
robot_description 0 25497.6
navigate_to_pose/_action/status 0.028 2389.333
global_costmapcostmap 0.500 996147.2
plan 0.884 3809.28
rosout 1.014 746.057
local_costmapcostmap 1.646 3986.99
amcl_pose 1.925 1054.72
behavior_tree_log 2.74 313.224
cmd_vel 18.151 147.110
camera/image_raw 25.57 1037110.93
scan 47.173 8846.012
odom 92.946 779.618
tf 387.278 195.413
5 Evaluation
To evaluate our proposal, we conducted an experimental study comparing the performance of an accountability system built in accordance with Section 4. This study covered several approaches for storing integrity proofs across various scenarios. Additionally, we explored and adapted the capabilities of RAG solutions and LLMs in question-answering systems, using as source the data gathered by our black box-like component, through a comprehensive set of questions. This work aimed to asses the effectiveness of these technologies in delivering natural language explanations to individuals who are not experts.
The experiments conducted within this proposal focus on navigation tasks included in the Navigation Functionality of the European Robotics League Consumer Service Robots Challenge (ERL Consumer) (ERL, 2023).
This functionality evaluates the capabilities of a robot to navigate accurately and autonomously within a given scenario, containing furniture, objects, and people. Details like the initial position and the number of waypoints the robot needs to visit are provided beforehand. Furthermore, the coordinates of the waypoints are communicated sequentially to the robot during runtime. Each waypoint is characterized by its $X$, $Y$, and $\Theta$ coordinates. Existing elements in the environment, such as furniture, doors, or walls, must be mapped. The obstacles, with varying shapes and sizes, remain unknown to the teams beforehand and may differ between runs.
Our experiments are structured around three distinct assessment scenarios in line with the navigation functionality previously described. In the initial scenario, the robot must navigate to three waypoints, encountering no obstacles along the way. The second scenario presents the robot with the challenge of adjusting its pre-planned path to reach the second goal. This adjustment becomes necessary due to the appearance of an obstacle obstructing the original planned route, requiring a deviation to achieve the goal. The third scenario introduces a new obstacle that blocks the only door needed to reach the first goal. Consequently, this obstacle prompts the cancellation of the first waypoint objective. The second and third goals are achieved with the same restrictions as detailed in the second scenario.
The three scenarios are represented in Figure 4.
The entire source code is accessible online on GitHub [^3] https://github.com/laurafbec/immutable_explainable_BBR.git, and is licensed under GPLv3.
Figure 4: Scenario maps with trajectories, goals, and obstacles. (a) Scenario 1: No obstacles scenario. (b) Scenario 2: One obstacle scenario. (c) Scenario 3: Two obstacles scenario.
5.1 Black Box Recorder evaluation
In evaluating the accountability component, we conducted six experiments across three previously described scenarios. These experiments involved integrity proofs storage in the blockchain, with hash calculations occurring at intervals of one every 10, 25, 50, or 100 messages. Moreover, our solution introduces the flexibility to save proofs of evidence at different rates based on the topic, ensuring non-tampering for critical topics or those with lower publishing rates where message loss could be crucial.
Therefore, our fifth recording method is designed to store a hash for every message within a topic with a rate below 0.5 Hz, one hash for every 5 messages for topics with rates below 1 Hz, a hash for every 10 messages if the rate is below 3 Hz, one hash for every 15 messages if the topic’s rate is below 25 Hz, a hash for every 50 messages for topics with rates below 45 Hz, and one hash for every 100 and 1000 messages when the rate falls below 100 Hz and 400 Hz, respectively. This approach, together with those operating at fixed intervals, has been compared with the generation of a Rosbag file that covers identical topics but lacks anti-tampering properties. Each experiment has been repeated six times, totalling 108 runs, to derive insights into the characteristics of the black box-like device.
5.1.1 Metrics
We employed different metrics to evaluate the impact of the accountability system in the chosen scenarios. First, through the topics’ publication rate and the number of messages stored per topic in each Rosbag file, we analyzed the average rate of lost messages based on the employed blockchain storage approach.
Next, regarding system performance, we studied variables such as CPU, memory and disk usage. Specifically, we examined the following variables: system load 1-, 5- and 15-minute average, CPU usage (%), memory usage (GB), sent network traffic (MB/s) and disk writing rate (MB/s). These variables are highly significant for system-overhead analysis given that elevated load conditions can potentially influence the robot’s behavior and event-recording capabilities. In particular, we have obtained a full report about system load, CPU, RAM, network traffic and disk write operations updated every second, facilitating the tracking of overall system activity and an easy comparison with preceding values. To describe the overall distribution and characteristics of these results, we calculated their average as statistical measure. This information is included in Section 6.
5.2 Explainability evaluation
Despite the remarkable performance of LLMs across diverse language tasks, evaluating these models presents challenges, especially in ensuring factual consistency. The extensive capabilities of LLMs and the limitations of current benchmarks in capturing human preferences, have given rise to emerging trends in LLM evaluations, such as employing strong LLMs as judges (Zheng et al., 2023). Traditional similarity-based metrics, such as BLEU and ROUGE, have proven to misalign with human evaluation. While human evaluation remains the primary standard for assessing human preferences, it is exceptionally time-consuming and expensive. To automate the evaluation process, several works have adopted the ’LLM-as-a-judge’ approach, which focuses on using strong language models (LLMs) like GPT-4. These models are often trained with reinforcement learning from human feedback (RLHF), resulting in strong human alignment (Lin, 2024). Previous studies have evidenced the consistent agreement of these models with human grading in over 80% of judgments (Bubeck et al., 2023; Helgesson Hallström, 2023; Leng et al., 2023). Given these characteristics, we opted for GPT-4 to assess the answers generated by our explainability component. Through LLM-as-a-judge, we are able to obtain not only scores but also explanations, making their outputs interpretable. The whole process is represented in Figure 5.
Figure 5: Explainability evaluation.
LLMs exhibit a certain degree of indeterminism, defined as the model’s ability to offer different answers when asked the same question or prompt several times. This behavior derived from their training, in which the models learn to generate outputs by sampling from probabilistic distributions over possible word sequences, introducing randomness to their response. The effect of this indeterminism must be considered in any LLM evaluation.
In addition, RAG for question-answering systems is one of the most difficult LLM use cases to evaluate. A basic RAG setup includes two main sources of error: the retriever and the LLM. The evaluation of the retriever aims at determining the relevance of its document selection in relation to a query. To this end, we conducted an experimental study running with 16 different questions over the data generated and processed during the navigation task in the three previously described scenarios. This set of questions has been evaluated six times in each scenario, giving a total of 288 questions and answers to perform the analysis. This approach ensures a thorough examination of the performance of the RAG system in various scenarios, showing potential strengths and weaknesses in our solution.
Using the context provided by the Rosbag data and an evaluation question, our approach includes an evaluation component for the generated explanations that contrasts the prediction made by the evaluated LLM and, thanks to the retrieved selection, compares it to an annotated ground truth. This process yields a correctness metric that scores whether the retriever returned no relevant documents (0) or whether it returned a perfect selection of documents (1). The final score, an average of all the scores in the dataset, is used to judge the retriever’s overall performance.
A comprehensive evaluation of a LLM should consider multiple metrics that include correctness as well as other aspects. A language model should produce text that is not only correct but also logically structured and easy to understand. Produced response must be appropriately detailed without unnecessary verbosity. Furthermore, the LLM must not only provide correct information but also delivers in a meaningful and valuable way to the user.
Therefore, to assess the performance of the selected LLM on specific attributes, our approach also includes a criteria-based evaluation covering accuracy, conciseness, coherence, and helpfulness. Selected criteria metrics are describe in Table 3. This structured assessment offers a means to assess the model’s effectiveness in generating responses, providing a broader understanding of the model’s capabilities and its alignment with the used scenarios. Employing a scorecard defined for each of these features, our solution assigns scores between 0 and 10 to the responses it generates. The use of this scorecard facilitates a detailed and quantitative analysis of the LLM’s output, ensuring a comprehensive understanding of both its strengths and areas for improvement. A complete example of the answers and evaluations obtained for a selected question in the three selected scenarios is included in Figure 6.
Table 3: Criteria-based evaluation metrics
Criteria metric
Description
Accuracy
Accuracy evaluates the extent to which the response conforms to the reference, ranging from being completely unrelated to being completely accurate and in line with the reference.
Helpfulness
Helpfulness evaluates the degree to which the response provides valuable information, ranging from providing no helpful information to being extremely helpful and offering comprehensive information.
Conciseness
Conciseness assesses the brevity of the response, ranging from being extremely verbose to highly concise with no unnecessary words.
Coherence
Coherence examines the organizational structure and logical flow of the response, from extremely inconsistent to highly coherent, with a clear and logical organization.
Question: How many goals have been reached by the robot? Scenario 1 Answer: Based on the provided context, we know that the robot has reached 3 goals. The context mentions that the goal number 1, 2 and 3 have been reached successfully. Therefore, the answer to your question is 3. GPT-4 Judgement Predicted Grade: CORRECT Reasoning: The AI assistant’s response is accurate, concise, coherent, and helpful. It directly answers the user’s question by confirming that the robot has reached three goals. The assistant also provides clarification by referencing the context, which adds to the response’s accuracy and helpfulness. There are no unnecessary words or unclear statements, making the response concise and coherent. Accuracy: [[10]] Conciseness: [[10]] Coherence: [[10]] Helpfulness: [[10]]", score: 1.0 Scenario 2 Answer: Based on the context provided, the robot has reached 3 goals. GPT-4 Judgement Predicted Grade: CORRECT Reasoning: The assistant’s response is completely accurate as it matches the information given in the ground truth. The response is concise, with no unnecessary words or information. It is highly coherent and has a clear structure. The response is extremely helpful as it gives the exact information asked for by the user. Therefore, the ratings are as follows: Accuracy: [[10]] Conciseness: [[10]] Coherence: [[10]] Helpfulness: [[10]]", score: 1.0 Scenario 3 Answer: Based on the provided context,the robot has successfully reached 2 out of 3 goals. The successful goals are: * Goal number 2: Navigation to the goal number 2 has succeeded. Position: -9.14068582195674, -25.977117025927537. Orientation: -0.6580310307304338,0.7529908117605705. * Goal number 3: Navigation to the goal number 3 has succeeded. Position: -1.8811226231679194, -28.773672979732826. Orientation: -0.5946646510401259,0.803973850820613. The aborted goal is: * Goal number 1: Navigation to the goal number 1 has been aborted. Position: -5.062160931312302, -8.330878461767831. Orientation: -0.8760794639388502,0.4821667479872674. Therefore, the answer to the user’s question is: The robot has successfully reached 2 out of 3 goals. Goal numbers 2 and 3 have been successfully achieved, while goal number 1 has been aborted. GPT-4 Judgement Predicted Grade: CORRECT Reasoning: The AI assistant’s response is comprehensive and accurate. It correctly states that the robot has successfully reached goals number 2 and 3 while goal number 1 has been aborted. The assistant also provides additional details about the positions and orientations of the goals. While these extra details are not explicitly asked for in the user’s question, they may be useful for the user’s understanding. Therefore, the assistant’s response is also helpful. However, the response is not concise because it includes more information than necessary, leading to some redundancy in the explanation. The assistant’s response is coherent as it logically organizes the information about each goal. Accuracy: [[10]] Conciseness: [[7]] Coherence: [[10]] Helpfulness: [[10]]", score: 1.0
Figure 6: Examples of correct answers, RAG, and criteria-based evaluations obtained across the three scenarios for the question ’How many goals have been reached by the robot?’
5.2.1 Categories and evaluation questions
To evaluate our explainability proposal in the context of ROS 2 messages interpretation, we have defined a set of questions covering diverse aspects of the navigation process over the how, the what, the which, the when and the where dimensions. These questions provide insights into different stages and components of the three scenarios. To structure and improve the understanding and analysis of this set of questions, we have organized them into five different categories. These categories range from high-level overviews of navigation to detailed inquiries about trajectory planning, goal completion, and the functioning of the behavior tree. The categories also delve into the causes of unexpected behavior or errors in navigation, improving the system explainability whenever an unexpected action occurs. Through these questions, we aim to address key concerns in XAI, including understanding why decisions are made, establishing causality between actions, determining success or failure, identifying timing, and disguising the methods by which achievements are realized (Gunning and Aha, 2019).
The set of categories, their descriptions and the questions included in each one are described next.
Navigation Process Overview
These questions provide an overview of the navigation process captured in the Rosbag file. Understanding what has happened and how the navigation task has proceeded provides a high-level context that sets the stage for more detailed questions in later categories. This category comprises the following questions:
  1. 1.
    What has happened in this ROS2 log regarding navigation?
  2. 2.
    How has the navigation task proceeded?
Trajectory Planning and Replanning
These questions delve into the specifics of trajectory planning, a critical aspect of navigation. They address whether the robot has re-planned its trajectory, the reasons behind this behavior, and whether it has encountered any obstacles. These details are necessary for understanding how the robot dynamically adjusts its path. The questions included in this category are:
  1. 3.
    Has the robot re-planned an alternative trajectory during navigation?
  2. 4.
    Why did the robot re-plan the route?
  3. 5.
    Did the robot find any obstacle during the navigation?
Goal Completion and Navigation Task Status
These questions focus on the completion status of the navigation goals. Knowing how many targets have been reached, whether the robot has completed the navigation task, and when the navigation task ended provides insights into the overall progress and success of the robot’s task. This category includes the following questions:
  1. 6.
    How many goals have been reached by the robot?
  2. 7.
    Has the robot completed the navigation task?
  3. 8.
    When has the robot ended the navigation task?
  4. 9.
    Have all objectives been successfully achieved or have any been cancelled or aborted?
Specifics about Goals and Locations
These questions examine the specifics of goals and locations. Understanding the location of goals, the robot’s linear velocity during navigation, its initial position and orientation, and the velocity after receiving a goal contribute to a detailed understanding of the robot’s movements and performance. Questions included in this category are:
  1. 10.
    Where is the second location or goal located?
  2. 11.
    What was the linear velocity when navigating to goal pose number 2?
  3. 12.
    What is the initial position and orientation of the robot?
  4. 13.
    What was the linear velocity of the robot after receiving the goal number 1?
Nav2 Behavior Tree and Node Status
These questions provide information about the underlying behavior tree and the state of its nodes in the navigation system. Knowing the specific node responsible for determining a viable path, identifying any failed nodes during navigation, and understanding the use of behavior tree nodes, contribute to a deeper understanding of the inner workings of the navigation system. This category includes:
  1. 14.
    What is Nav2 Behavior Tree’s node to determine a viable path from a starting point to a specified target pose or location?
  2. 15.
    Did any node from the Nav2 Behavior Tree fail during navigation?
  3. 16.
    Which Behavior Tree’s nodes were used during navigation?
6 Experimental results and discussion
This section summarizes the experiments’ results involving both the black box-like engine and the explainability component when using our solution in the previously depicted assessment scenarios.
6.1 Message loss rates
Figure 7 shows the percentage of lost messages in the Rosbag file for experiments detailed in Section 5, capturing the topics outlined in Table 2, attending to their publication rates and to the approaches selected to store integrity proofs in blockchain. The results across the three scenarios demonstrate a consistent trend, showing uniform message loss patterns in the creation of immutable Rosbag files.
Choosing to store a single message hash for every 100 messages results in a loss of less than 5% compared to avoiding anti-tampering measures in the Rosbag recording process, even for topics with rates exceeding 300 Hz. This approach stands out as the most efficient option for integrating anti-tampering techniques into our accountability solution.
Alternative strategies, such as saving one proof every 50 messages or adjusting proof intervals based on topic rates, achieve losses below 10% for each topic. This is particularly noteworthy, ensuring the integrity of messages based on their significance in a given scenario or task.
Introducing an integrity proof every 25 messages leads to an overall reduction in message recording, especially evident when topic rates surpass 100 Hz. Storing one hash every 10 messages could lead to losses exceeding 30% for topics with rates higher or around 100 Hz. Beyond the consistent linear trend of message loss with high topic rates, a slight deviation is observed in this approach when topics fall below 50 Hz. This variation signifies the influence on message recording when computing hash functions for messages nearing 1 MB in size, as observed in messages from the /camera/image_raw topic, given that the computational complexity of common hash functions, such as SHA-256, increases with input size.
Figure 7: Message loss rates in Rosbag Files under several anti-tampering techniques. Comparative visualizations for: (a) Scenario 1, (b) Scenario 2, and (c) Scenario 3.
6.2 System benchmarking
The system load or, the running thread demand on the system as the average number of running and idle threads, is shown in Figure 8. The average of 1 minute is higher than the average of 5 or 15 minutes in all scenarios, showing a load increase. However, the values of this metric do not exceed the CPU count, which is 6 in our experimental setup, indicating no performance issues or resource saturation. The inclusion of runnable and uninterruptible tasks in this metric means that the load average may increase due to a disk I/O workload, not just CPU demand, giving a more general view of system overhead. Other more specific metrics are described next to clarify the possible existence of a bottleneck in our approach.
Figure 8: Load average under several anti-tampering techniques. Comparative visualizations for: (a) Scenario 1, (b) Scenario 2, and (c) Scenario 3.
The CPU consumption by user processes maintains a consistent linear trend across all approaches and scenarios, remaining around 30%. This level of consumption is generally considered acceptable for resource-intensive processes such as the recording of a Rosbag file, providing a substantial margin in the event of introducing other critical tasks into the system.
Concerning memory consumption, all approaches within the three scenarios exhibit comparable values, differing by no more than 0.5 GB. Consequently, the inclusion of blockchain approaches into our black box recorder does not yield a remarkable impact on RAM comsuption in comparison to avoiding the use of anti-tampering techniques in this recording process.
Regarding network traffic, approaches with higher process storage in the blockchain increase this metric, but do not exceed 0,20 MB/s, which is a very affordable quantity. Furthermore, approaches such as storing one hash every 100 messages nearly match the avoidance of storing any proof into the blockchain, indicating that including this approach do not have a remarkable impact in the system performance.
Disk write operations do not form a bottleneck in system performance, with all approaches differing by no more than 0,40 MB/s, which is not a very significant impact. In addition, all the approaches that perform integrity proof storage on the blockchain achieve lower or similar results to the approach avoiding anti-tampering operations.
Figure 9 shows the results described.
Figure 9: System benchmarking results: (a) CPU usage of user processes. (b) RAM usage. (c) Sent network traffic. (d) Disk writes traffic.
These findings may encourage the adoption of our approach to ensure integrity in messages from autonomous agents.
6.3 Correctness evaluation
Figure 10 shows the results of the correctness of RAG evaluation, presenting both the overall results and assessments based on predefined categories. The results demonstrate remarkable consistency across the three scenarios, getting a significant number of correct answers, even though there has not been a prior tuning in the ROS 2 messages for autonomous agents navigation tasks. The slight percentage of incorrect answers remains within an acceptable threshold, given the specific context employed and the inclusion of general questions.
Figure 10: (a) Overall correctness in the three scenarios. (b) Correctness by category in the three scenarios.
Regarding individual categories, the correctness in the category ’Navigation Process Overview’ is relatively consistent across all scenarios, indicating a good understanding of events and progress related to navigation. However, the broad nature of the question "What has happened in this ROS2 log regarding navigation?" initially yielded a correctness below 65% in the third scenario. In this instance, the cancellation of a goal resulted in the execution and failures of Nav2 Behavior Tree nodes, not included in the other two scenarios and focused on navigation recovery. Therefore, the incorporation in a dataset of information related to the cancellation of a goal, results in LLM asnwers that usually diverge from the ground truth provided, as the model occasionally prioritized Behavior Trees information, other times focused on locations or timestamps. This variability made it difficult to align with the ground truth provided, even if the obtained answers could be considered partially accurate. Introducing the prompt extension "Describe the sequence of events regarding navigation goals and include information about each of them." in question 1 significantly increased correctness in the first category, achieving results around and exceeding the 75% in all three scenarios. This fact underscores the effectiveness of prompt engineering when tackling ambiguous or broad queries.
The results also show consistently high correctness across all scenarios concerning questions related to trajectory planning and re-planning. The model is able to determine whether the modification of a predicted trajectory has been caused by the occurrence of an obstacle or other circumstance. Scenario 1 and Scenario 2 also exhibit relatively high correctness in questions related to goal completion and task status. However, the model encounters challenges in Scenario 3, occasionally failing to identify if a navigation that starts correctly concludes with an abortion or cancellation. This issue may be attributed to misalignment during the retrieval process, specifically in the selection of the best chunk associated with a particular goal associated with numerous embeddings within a single collection from the vector database.
Correctness in identifying specific targets and locations lacks some consistency, with results exceeding 50% in the first scenario, around 70% in the second and over 75% in the third. In particular, three scenarios encounter difficulties in identifying the initial position of the robot before starting the navigation, due to the model misaligns the queried position with the one assumed by the robot when starting the navigation process. This fact may also be attributed to deficiencies in the retrieval process. In Scenario 1, understanding the sequence of events or the order of goals results challenging for the model, leading to inaccuracies in specifying the second location or goal. Nevertheless, this challenge is not present in Scenario 3 and absent or reduced in Scenario 2, demonstrating how differences in data distribution can introduce varied responses even when using specific and unambiguous questions.
Through questions included in the Nav2 Behavior Tree and node status category, we assessed the effectiveness of the model’s formal linguistic competence compared to its functional linguistic competence, a feature of present LLMs (Mahowald et al., 2023). While the model excels in identifying Behavior Tree Nodes given its description, it faces challenges in pinpointing these components under specific conditions. As a result, the category correctness suffers a drop in Scenario 3, although overall results remain above 75%. In this scenario, the navigation recovery actions caused by the cancellation of the first goal introduced a challenge in the interpretation of the information by the LLM due to the execution of several Nav2 Behavior Tree nodes not present in the others scenarios. Particularly, the mistakes stem from the model’s confusion in determining whether a previously failed Nav2 Behavior Tree node has recovered or finished its execution, as well as identifying all nodes that were executed during navigation. This can be attributed to the fact that the complexities of the recovery actions and the dynamic nature of Behavior Tree executions during navigation recovery introduce complexities that the model may not fully generalize from the training data, underscoring the need for additional fine-tuning in such scenario-specific contexts in case the obtained correctness needs to be improved.
6.4 Criteria-based evaluation
Figure 11 shows the remarkable performance of the Large Language Model’s (LLM) in the evaluated metrics, demonstrating consistency with slight variations in the scores. On average, the answers provided are accurate and in line with the reference, though there might be room for improvement. While responses generally provide useful information, there are instances where additional depth could enhance their utility. In addition, most answers usually include a remarkable balance between conciseness and effectiveness, showing a proficient level of brevity, as well as clear and logical organization. The overall results show no notable differences in these metrics across the three scenarios. This suggests that messages within the Rosbag related to the introduction of obstacles or cancellation of goals do not notably influence the general criteria evaluation outcomes.
Figure 11: (a) Overall criteria evaluation in the three scenarios.
The language model’s adaptability to challenges like obstacles and goal cancellations regarding the described categories, is shown in Figure 12.
Figure 12: Large Language Model criteria-based evaluation by categories. Comparative visualizations for: (a) Scenario 1, (b) Scenario 2, and (c) Scenario 3.
Concerning general questions about about events and the progression of the navigation task, we observe variations in scores across Scenarios 2 and 3. While coherence and conciseness are even improved, Scenario 1, involving direct navigation to three goals, outperforms in accuracy and helpfulness. This fact suggests that effective communication on general queries benefits from the absence of obstacles or goal cancellations, enhancing overall understanding.
In the category ’Trajectory Planning and Replanning’, the questions are effectively addressed in all the three scenarios, showing scores consistently above 8 in every evaluated metric, even when obstacles are introduced or a goal is cancelled. These results underlines the system’s proficiency in handling trajectory-related questions despite potential new challenges.
Regarding the category ’Goal Completion and Navigation Task Status’, Scenarios 1 and 2 demonstrate the effectiveness of the system in conveying goal completion and navigation task status information. Nevertheless, the slightly lower scores for Scenario 3 indicate potential challenges in effectively communicating completion status when a goal is aborted.
Concerning the ’Specifics about Goals and Locations’ category, the system consistently provide valuable goal and location information, even in the presence of obstacles or goal cancellations. This suggests a robust capacity to understand and localize specific details regardless of challenges introduced in each scenario.
The results regarding questions included in ’Nav2 Behavior Tree and Node Status’ demonstrate effectiveness in communication across all scenarios and metrics. This emphasizes the potential for Behavior Trees to complement language models in explainability, offering insight into the decision-making processes during navigation tasks.
In summary, although results can be improved in accuracy regarding general questions, the language model exhibits adaptability to different navigation scenarios, facing obstacles and changes in navigation tasks. The consistency in conveying information related to Nav2 Behavior Trees suggests a robust understanding of underlying navigation processes. The model also performs well when providing information about trajectory planning, goal completion, and specifics about goals and locations, regardless of the specific challenges introduced in each scenario. Its robustness in effective communication, despite technical vocabulary, details and the introduced challenges, is demonstrated through the achievement of responses that are not only generally accurate but also easily comprehensible to users without expertise, which enhances the model’s practical utility.
7 Conclusions
This work presents an accountability and explicability architecture, implemented for ROS-based mobile robots. Our approach comprises a black box component for accountability that effectively addresses real-time recording challenges, ensuring message integrity and verifiability through blockchain technology, and an explainability component. This last component is based on the use of RAG and LLM through a private question-answering system designed to use the information recorded by the first component as context. This solution demonstrates the feasibility of obtaining accurate, coherent, concise and helpful explanations for robot actions from raw data stored in Rosbag files, thereby enhancing human-robot interaction. Additionally, both main components have been integrated through the development of a third main functionality in charge of filtering non-relevant information and making messages closer to the final user, due to their specific nature can introduce challenges in their comprehension process. Furthermore, this last component releases software developers from making direct calls to the logging API, easing the development process and improving the scalability and adaptability of our solution for every navigation task performed by a ROS-based robot.
The results achieved highlight the capability of our solution to guarantee the immutability of information generated by an agent without compromising its performance, a feature particularly important when trying to identify the causes behind an event. Moreover, our system improves efficiency and effectiveness in the deployment of autonomous agents, providing natural language explanations across key navigation aspects including goal achievement, obstacle appearance, trajectory re-planning, and objective cancellation or abortion. This fact strengthens trust and confidence in the use of autonomous agents across diverse applications and contexts.
While our approach has shown remarkable results in navigation functionalities regarding several scenarios and circumstances, future work should involve adapting our solution to other applications by simply adapting the recording and the processing components to the relevant topics of the area or domain of use. Furthermore, the consistency and robustness of responses generated by our explainability approach could benefit from refinement across different scenarios, especially when categorizing questions under specific subjects. Therefore, future work should explore the impact of In-Context Learning (ICL) techniques on prompt generation, in addition to fine-tuning the model for the specific context of autonomous agent navigation tasks.
Additionally, our RAG approach could improve in terms of retrieving the most relevant chunk of information from the context. This enhancement may involve exploring the use of separate collections within the vector database or employing a hybrid approach, such as integrating knowledge graphs (KG) with vector indexes. These approaches can help to improve the accuracy, quality and overall consistency of the answers obtained across several scenarios, thereby enhancing the performance of our proposed system.
[…]
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
Data and tools are available in a public repository.
Acknowledgments
This research has been partially funded by the Recovery, Transformation, and Resilience Plan, financed by the European Union (Next Generation) thanks to the TESCAC project (Traceability and Explainability in Autonomous Systems for improved Cybersecurity) granted by INCIBE to the University of León; and by grant PID2021-126592OB-C21 funded by MCIN/AEI/10.13039/501100011033 EDMAR (Explainable Decision Making in Autonomous Robots) project, PID2021-126592OB-C21 funded by
MCIN/AEI/10.13039/501100011033 and by ERDF ”A way of making Europe”.

### v2

@@
 Consequently, understanding the circumstances behind an event becomes critical, requiring the development of capabilities to justify their behaviors to non-expert users.
-Such explanations are essential in enhancing trustworthiness and safety, acting as a preventive measure against failures, errors, and misunderstandings.
+Such explanations are essential in enhancing trustworthiness and safety.
 Additionally, they contribute to improving communication, bridging the gap between the agent and the user, thereby improving the effectiveness of their interactions.
-This work presents an accountability and explainability architecture implemented for ROS-based mobile robots.
+This work presents an accountability and explainability architecture implemented for mobile Robot Operating System (ROS)-based robots.
 The proposed solution consists of two main components.
-Firstly, a black box-like element to provide accountability, featuring anti-tampering properties achieved through blockchain technology.
-Secondly, a component in charge of generating natural language explanations by harnessing the capabilities of Large Language Models (LLMs) over the data contained within the previously mentioned black box.
+First, a black box-like element to provide accountability, featuring anti-tampering properties achieved through blockchain technology.
+Second, a component in charge of generating natural language explanations by harnessing the capabilities of Large Language Models (LLMs) over the data contained within the previously mentioned black box.
+The integration of blockchain and generative Artificial Intelligence (AI) solutions ensures the deployment of responsible, transparent, interpretable, and trustworthy agents.
 The study evaluates the performance of our solution in three different scenarios, each involving autonomous agent navigation functionalities.
@@
 1 Introduction
-The rising presence of mobile robots alongside humans has increased growing security and safety concerns.
-One of the main challenges in this context is the ability of robots to justify their actions and decisions, particularly in safety-critical scenarios (Sado et al., 2023).
-Therefore, it is essential to ensure that the events performed when developing an action or behavior are reliable and transparent (Borgo et al., 2018).
-Equally important is the capacity to render such explanations comprehensibly to non-expert users, offering appropriate, effective, and natural answers to their queries (Rosenfeld and Richardson, 2019; Cruz et al., 2023).
+The rising presence of mobile robots alongside humans has increased security and safety concerns.
+One of the main challenges in this context is the ability of robots to justify their actions and decisions, particularly in safety-critical scenarios [1].
+Therefore, it is essential to ensure that the events performed when developing an action or behavior are reliable and transparent [2].
+Equally important is the capacity to render these explanations comprehensible to non-expert users, offering them appropriate, effective, and natural responses to their queries [3, 4].
 The most substantial evidence in post-event analysis relates to the logging and auditing capabilities, which are essential to mitigate technical issues and provide forensic evidence to detect any data tampering or unauthorized leakage.
-Therefore, a trustworthy and secure log system is needed when generating a complete trace of a robot’s behavior (Al-Slais and Ali, 2023).
+Therefore, a trustworthy and secure log system is needed when generating a complete trace of a robot’s behavior [5].
 The system must ensure the integrity, confidentiality, and availability of the agent and the data.
@@
 Regarding log systems vulnerabilities, covering tracks has become a common practice for attackers, erasing any traces that could lead to their identification or assisting in the analysis of their actions.
-As a result, it is necessary to strengthen system properties against unauthorized modifications (Ahmad et al., 2019).
+As a result, it is necessary to strengthen system properties against unauthorized modifications [6].
 This fact not only includes safeguarding the integrity of log data but also implementing measures to detect and counteract attempts at tampering.
-Accountability is equally required, enabling detection and isolation of faulty behaviors and their origin.
-It must not only provide tamper resistance, but also to offer verifiability for checking the integrity of log entries against tampering.
-Moreover, it should guarantee search accessibility, enabling access to specific records whose integrity can be verified (Chiu et al., 2021).
+Accountability is equally required, enabling the detection and isolation of faulty behaviors and their origin.
+It must not only provide tamper resistance but also offer verifiability for checking the integrity of log entries against tampering.
+Moreover, it should guarantee search accessibility, enabling access to specific records whose integrity can be verified [7].
 Integrating accountability solutions into robotic systems often demands a trade-off with system performance.
 Specifically, features such as computational power and networking bandwidth can be negatively affected.
-The inclusion of anti-tampering techniques, such as full disk and communication encryption, may be restrictive due to resource limitations (Fernández-Becerra et al., 2022).
-Additionally, high-cost tamper proof storage devices may prove economically unviable, especially when dealing with continuously generated log data.
-Therefore, it becomes necessary to establish an efficient and verifiable proof of integrity to guarantee the reproducibility and identification of the events that could arise during an autonomous agent action.
-For this evidence to be significant, it should belong to a persistent chain of custody, totally reliable and verifiable, main requirements for an auditable and transparent logging system (Putz et al., 2019; Hajer et al., 2024).
+The inclusion of anti-tampering techniques, such as full disk and communication encryption, may be restrictive due to resource limitations [8].
+Additionally, high-cost tamper-proof storage devices may prove economically unviable, especially when dealing with continuously generated log data.
+Therefore, it is necessary to establish an efficient and verifiable proof of integrity to guarantee the reproducibility and identification of the events that could arise during an autonomous agent action.
+For this evidence to be significant, it should belong to a persistent chain of custody, totally reliable and verifiable, main requirements for an auditable and transparent logging system [9].
 Blockchain solutions emerge as a mechanism to ensure data integrity, protection against unauthorized changes, and traceability.
 The consensus protocols established by all blockchain network members enable the detection and rejection of any destruction, manipulation, or non-permitted editing of data through illicit channels.
-The nature of this technology helps to avoid tampering risk, safeguarding data (Iqbal and Matulevičius, 2019) or even .
-Nevertheless, the potential obtained evidence poses significant challenges for analysis due to its diverse and varied nature, terminology, and the presence of substantial volumes of similar messages unrelated to anomalous events.
-These facts may result in limited usefulness and misalignment with the expectations of non-expert users (Shehu and Harper, 2022).
-Regarding autonomous agents, the raw log messages produced by these systems, are usually noisy and possess semi-structured format.
-In addition to this, it is necessary to deal with specific features such as large quantities of domain-specific language messages (Ott et al., 2021).
+The nature of this technology helps to avoid tampering risk, safeguarding data [10].
+Nevertheless, the potential evidence obtained poses significant challenges for analysis due to its diverse and varied nature, terminology, and the presence of substantial volumes of similar messages unrelated to anomalous events.
+These facts may result in limited usefulness and misalignment with the expectations of non-expert users [11].
+Regarding autonomous agents, the raw log messages produced by these systems are usually noisy and possess a semi-structured format.
+In addition to this, it is necessary to deal with specific features such as large quantities of domain-specific language messages [12].
 In this particular context, explaining the reasons that make a robot carry out a specific and unexpected action is essential for diagnosing failures in goal achievement.
-Besides, making these explanations accessible to humans is equally important.
+Besides, making these explanations accessible to humans is equally important [13].
 This requirement has led to the development of eXplainable Autonomous Robots (XAR), which aims to provide mechanisms for presenting clear and understandable explanations for the actions and decisions of autonomous agents.
-XAR enhances the transparency and interpretability of these agents’ decision-making processes, often incorporating techniques like natural language explanations to improve clarity (Anjomshoae et al., 2019; Langley et al., 2017).
-Large Language Models (LLMs) have gained significant attention due to their exceptional performance in diverse natural language processing tasks, including translation, coding or text summarization (Kocmi and Federmann, 2023; Zhang et al., 2024; Sarsa et al., 2022).
-One approach to improve the results of large language models for a specific domain is through the use of retrieval-based methods.
-By retrieving relevant information from a large corpus of text, these models can provide explanations grounded in existing knowledge (Ren et al., 2023).
+XAR enhances the transparency and interpretability of these agents’ decision-making processes, often incorporating techniques like natural language explanations to improve clarity [14, 15].
+Large Language Models (LLMs) have gained significant attention due to their exceptional performance in diverse natural language processing tasks, including translation, coding or text summarization [16, 17, 18].
+One approach to improve the results of LLMs for a specific domain is through the use of retrieval-based methods.
+By retrieving relevant information from a large corpus of text, these models can provide explanations grounded in existing knowledge [19].
+The integration of Artificial Intelligence (AI) and blockchain has received significant attention in recent years, with researchers exploring various ways of combining these two areas to provide efficiency, security, reliability and transparency to different domains [20, 21].
+Our contribution demonstrates how bringing these technologies together can improve the behavioral analysis of autonomous agents, enhancing their interactions with humans and thus increasing their acceptance.
 This work presents an accountability and explainability architecture implemented for ROS-based mobile robots.
-This solution included two main components.
-Firstly, a black box-like component to provide accountability that includes anti-tampering features by distributing integrity proofs from the events recorded by the black box component.
+This solution includes two main components.
+First, a black box-like component to provide accountability that includes anti-tampering features by distributing integrity proofs from the events recorded by the black box component.
 The second component generates natural language explanations from data in the black box, including the use of interactive natural language conversations as a means to provide comprehensive explainability for the recorded events.
 Both components are connected through an additional component responsible for processing raw data collected by the first one, eliminating non-essential information and building connections and relations between data.
-These functionalities release software developers from the need to make calls to the log API to build explanations for the code generated and the behaviors developed by the autonomous agent.
+These functionalities release software developers from the need to make calls to the logging Application Programming Interface (API) to build explanations for the code generated and the behaviors developed by the autonomous agent.
 Additionally, we have conducted a thorough evaluation of our solution, including its impact on system performance and the correctness, accuracy, helpfulness and coherence of the generated natural language explanations.
@@
 In this section, we explore key works related to various domains relevant to our approach.
-It includes progress in understanding how blockchain ensures the immutability of logs, research on the application of event data recorders in robotics, and significant contributions Explainable Autonomous Robots (XAR) field.
-Lastly, we examine the evolution of Large Language Models (LLMs) to provide natural language explanations.
+It includes progress in understanding how blockchain ensures the immutability of logs, research on the application of event data recorders in robotics, and significant contributions XAR field.
+Lastly, we examine the evolution of LLMs to provide natural language explanations.
 2.1 Blockchain for Immutable Logs
@@
 Any attempt to alter a transaction in one block requires the modification of all subsequent blocks, resulting in a computationally expensive process.
-As a result, data stored in the blockchain is immutable (Rajasekaran et al., 2022; Guo and Yu, 2022).
+As a result, data stored in the blockchain is immutable [22, 23].
 There are two primary types of blockchains based on access controls: permissionless blockchains allow open access and public transaction visibility, while permissioned blockchains limit participation and transaction visibility.
-Nevertheless, closed networks may pose security risks, as authorized users could recalculate subsequent blocks in the event of a block modification (Solat et al., 2021).
-While permissionless blockchains have traditionally related to scalability and throughput challenges, approaches like Ethereum have transitioned to Proof of Stake (PoS), a more energy-efficient, secure, accessible and scalable consensus protocol (Nguyen et al., 2019; Sanka and Cheung, 2021).
-Blockchain technology ensures log integrity and auditability, offering unique advantages over alternatives such as distributed storage systems or IPFS (InterPlanetary File System), which may require central authorities for storage and verification (Omar et al., 2024).
+Nevertheless, closed networks may pose security risks, as authorized users could recalculate subsequent blocks in the event of a block modification [24].
+While permissionless blockchains have traditionally related to scalability and throughput challenges, approaches like Ethereum have transitioned to Proof of Stake (PoS), a more energy-efficient, secure, accessible and scalable consensus protocol [25, 26].
+Blockchain technology ensures log integrity and auditability, offering unique advantages over alternatives such as distributed storage systems or InterPlanetary File Systems (IPFSs), which may require central authorities for storage and verification [27, 28].
 Blockchain offers consensus and immutability, coupled with the capability of integrating logic mechanisms such as Smart Contracts (SCs).
-SCs contain functions that either write values onto the blockchain, thus altering its state, or read data from it (Mohanta et al., 2018).
-These features make blockchain the preferred choice for applications where data integrity and trust are required (Nyaletey et al., 2019).
-Different secure logging systems rely on permissioned blockchain frameworks for log immutability (Shekhtman and Waisbard, 2019), introducing accountability even in embedded systems (Chiu et al., 2021).
-To get a balance between throughput and cost-effectiveness, some works combine both categories (Zhao et al., 2023), developing approaches that include a permissioned blockchain to ensure per-entry immutability and public non-repudiation solution (Pourmajidi et al., 2023).
-Nevertheless, inherent features of permissionless blockchain such as transparency and trust, have also boosted the development of secure logging systems (Ge et al., 2019).
+SCs contain functions that either write values onto the blockchain, thus altering its state, or read data from it [29].
+These features make blockchain the preferred choice for applications where data integrity and trust are required [30].
+Different secure logging systems rely on permissioned blockchain frameworks for log immutability [31], introducing accountability even in embedded systems [7].
+To get a balance between throughput and cost-effectiveness, some works combine both categories [32], developing approaches that include a permissioned blockchain to ensure per-entry immutability and public non-repudiation solution [33].
+Nevertheless, inherent features of permissionless blockchain such as transparency and trust, have also boosted the development of secure logging systems [34].
 2.2 Robotic Event Data Recorders
-Event Data Recorders (EDRs), often referred to as black boxes, have traditionally played a crucial role in in capturing and analyzing events within robotics systems.
-These recorders support the collection and storage of event data, helping to comprehend system behavior, detecting anomalies, and ensuring safety (Winfield et al., 2021, 2022).
-Therefore, these solutions are needed in forensic investigations, ethical assessments and transparency regarding autonomous agents (Taurer et al., 2018).
-The integration of black box modules to enhance robots’ introspection capabilities contributes to improving system transparency by attributing responsibilities (Fernández-Becerra et al., 2022).
-These solutions typically include functionalities such as hashing, signing, and encryption (Schmidt et al., 2022).
-However, to ensure the integrity and authenticity of recorded events, anti-tamper features must also be in place (Srinivas Aditya et al., 2021).
+Event Data Recorders (EDRs), often referred to as black boxes, have traditionally played a crucial role in capturing and analyzing events within robotics systems.
+These recorders support the collection and storage of event data, helping to comprehend system behavior, detecting anomalies, and ensuring safety [35, 36].
+Therefore, these solutions are needed in forensic investigations, ethical assessments and transparency regarding autonomous agents [37].
+The integration of black box modules to enhance robots’ introspection capabilities contributes to improving system transparency by attributing responsibilities [38].
+These solutions typically include functionalities such as hashing, signing, and encryption [39].
+However, to ensure the integrity and authenticity of recorded events, anti-tamper features must also be in place [40].
 Using blockchain technology in event data recorders can guarantee the integrity of a transparent and auditable record of events.
-These features enhance accountability and trust in the recorded data, as it becomes difficult for any party to manipulate or tamper with the recorded events without detection (Salah et al., 2019).
+These features enhance accountability and trust in the recorded data, as it becomes difficult for any party to manipulate or tamper with the recorded events without detection [41].
 Additionally, blockchain involves decentralized and distributed storage of event data, eliminating the necessity for central authorities or intermediaries and reducing the risk of data loss.
 Further research explores the application of blockchain technology to improve robotic capabilities, particularly event data recording and management.
-The Black Block Recorder (BBR) system proposed in (White et al., 2019), employs permissioned blockchain technology for immutable black box logging in robotic applications.
+The Black Block Recorder (BBR) system proposed in [42], employs permissioned blockchain technology for immutable black box logging in robotic applications.
 The BBR system combines Rosbag2 and Distributed Ledger Technologies (DLTs) to ensure data integrity and immutability.
-The potential applications of blockchain in robot swarms, including secure communication, data logging, and consensus agreement are discussed in (Strobel et al., 2020).
-In (Lopes et al., 2021), a time-segmented consortium blockchain for robotic event registration is proposed, organizing and storing event records efficiently to allow for retrieval and analysis of specific events within specific time frames.
-Authors in (Zhang et al., 2022) present a tool to bridge the Robot Operating System (ROS) and the Ethereum blockchain, evaluating encryption efficiency and stability, transactions response time and packet loss rate.
-Nevertheless, these works lack capabilities to save integrity proofs in the blockchain at selected intervals, to enhance granularity, or including service definitions or smart contracts to improve the verifiability process.
+The potential applications of blockchain in robot swarms, including secure communication, data logging, and consensus agreement are discussed in [43].
+In [44], a time-segmented consortium blockchain for robotic event registration is proposed, organizing and storing event records efficiently to allow for retrieval and analysis of specific events within specific time frames.
+Authors in [45] present a tool to bridge Robot Operating System (ROS) and the Ethereum blockchain, evaluating encryption efficiency and stability, transactions response time and packet loss rate.
+Nevertheless, these works lack capabilities to save integrity proofs in the blockchain at selected intervals, to enhance granularity, or to include service definitions or SCs to improve the verifiability process.
 2.3 Explainable Autonomous Robots
-The need for explainability in autonomous robots arises from increasing complexity and autonomy of these systems, making it challenging for users to comprehend their behavior.
+Explainable Artificial Intelligence (XAI) is a research area focused on improving the transparency and interpretability of AI systems, thereby promoting their responsible deployment.
+The need for explainability in autonomous robots arises from the increasing complexity and autonomy of these systems, making it challenging for users to comprehend their behavior.
 Explainable Autonomous Robots (XARs) are designed to provide explanations for their actions and decisions to non-expert users.
-Such explanations enrich aspects like security and the trust level, preventing failures, errors, or conflicts in achieving objectives or undesired interactions (Han et al., 2019).
-Despite the growing demand for explainability in autonomous agents, their use reveals a lack of practical and effective usability or interpretability in real environments (Abdul et al., 2018).
+Such explanations enrich aspects like security and the trust level, preventing failures, errors, or conflicts in achieving objectives or undesired interactions [46].
+Despite the growing demand for explainability in autonomous agents, their use reveals a lack of practical and effective usability or interpretability in real environments [47].
 Therefore, it is a problem that requires special attention, being necessary to explain the performed decisions.
-This explanation should include alternatives, choices made, expectations of each option, decision criteria, and any unexpected events that may have led to the modification of the initially planned behavior (Langley et al., 2017; Sadeghi et al., 2021).
-As described in (Sakai and Nagai, 2022), an explanation regarding XAR, is an "abductive reasoning that bridges the gap between the individual explaining, which in this case may be a robot, and the person to whom they are explaining".
-Self-explanatory capabilities in social robots have been provided through the development of architectures and frameworks (Stange et al., 2022; Han et al., 2021).
-Nevertheless, the essence of explainability regarding XAR lies in effective communication (Sakai and Nagai, 2022).
-Progress in natural language processing (NLP) technology through large language models (LLMs), based on the estimation of human statements and the relevance and position of the words contained in them, could improve the efficiency of human-robot interactions.
-This enhancement can create a sense that these systems truly comprehend and respond to individuals meaningfully, significantly benefiting their usefulness and applicability (Sado et al., 2023).
+This explanation should include alternatives, choices made, expectations of each option, decision criteria, and any unexpected events that may have led to the modification of the initially planned behavior [15, 48].
+As described in [49], an explanation regarding XAR is an "abductive reasoning that bridges the gap between the individual explaining, which in this case may be a robot, and the person to whom they are explaining".
+Self-explanatory capabilities in social robots have been provided through the development of architectures and frameworks [50, 51].
+Nevertheless, the essence of explainability regarding XAR lies in effective communication [49].
+Progress in natural language processing (NLP) technology through LLMs, based on the estimation of human statements and the relevance and position of the words contained in them, could improve the efficiency of human-robot interactions.
+This enhancement can create a sense that these systems truly comprehend and respond to individuals meaningfully, significantly benefiting their usefulness and applicability [1].
 2.4 Large Language Models
-Large Language Models (LLMs) have recently played a significant role in natural language processing, thanks to replacing recurrent neural networks with self-attention mechanisms (Vaswani et al., 2017).
+LLMs have recently played a significant role in natural language processing, thanks to replacing recurrent neural networks with self-attention mechanisms [52].
 The design and implementation of the first Transformers models constitute the foundation of most LLM architectures.
-Trained on massive text data, LLMs include strong capabilities in natural language understanding and complex task solving (Zhao et al., 2023).
+Trained on massive text data, LLMs include strong capabilities in natural language understanding and complex task solving [53].
 However, LLMs face two primary challenges.
-First, they tend to produce a significant amount of hallucinations due to their reliance on internal knowledge (weights) (Andriopoulos and Pouwelse, 2023).
+First, they tend to produce a significant amount of hallucinations due to their reliance on internal knowledge (weights) [54].
 Second, their large parameter sizes and the associated high update costs, make traditional pre-training and fine-tuning methods unfeasible.
-As a solution, Retrieval-Augmented Generation (RAG) methods emerge as a solution for LLMs to effectively interact with the external world (Zhao et al., 2023).
-This strategy is particularly notable in question-answering systems, which represent the most widely adopted use cases for harnessing explanatory and reasoning capabilities of large language models.
+As a solution, Retrieval-Augmented Generation (RAG) methods emerge as a solution for LLMs to effectively interact with the external world [55].
+This strategy is particularly notable in question-answering systems, which represent the most widely adopted use cases for harnessing the explanatory and reasoning capabilities of LLMs.
 In this domain, RAG-based approaches are emerging as the standard for developing solutions that seamlessly integrate the advantages of a knowledge base (utilizing a vector store) and generative models.
-LLMs display emergent abilities, such as in-context learning (ICL), where LLMs make predictions based on contexts augmented with a few examples, helping the model to learn from analogy (Wei et al., 2022).
-This paradigm simplifies inclusion of human knowledge into LLMs by modifying the prompt templates (Dong et al., 2022).
+LLMs display emergent abilities, such as In-Context Learning (ICL), where LLMs make predictions based on contexts augmented with a few examples, helping the model to learn from analogy [56].
+This paradigm simplifies the inclusion of human knowledge into LLMs by modifying the prompt templates [57].
 ICL does not require altering the model’s parameters or training the model on a specific dataset; instead, it relies on guiding the LLM through the prompting interface to align its results with human values and preferences.
-Nevertheless, evaluating large language models (LLMs) and the impact of previously described approaches, poses several challenges due to their complexity and the nature of their outputs.
-These difficulties can be attributed to several factors, including the lack of suitable evaluation metrics, biases in generated text, and the cost of human assessment (Chang et al., 2024).
-Nevertheless, recently, employing LLMs as evaluators for LLMs-based solutions has been proven to be nearly as effective as human judges (Zheng et al., 2023).
-Previous works have demonstrated the potential of using LLMs to generate natural language descriptions of logging systems (Setianto et al., 2022) or robots’ behaviors (González-Santamarta et al., 2023), indicating room for improvements to deliver more accurate and complete interpretations.
+Nevertheless, evaluating LLMs and the impact of previously described approaches poses several challenges due to their complexity and the nature of their outputs.
+These difficulties can be attributed to several factors, including the lack of suitable evaluation metrics, biases in generated text, and the cost of human assessment [58].
+Nevertheless, recently, employing LLMs as evaluators for LLMs-based solutions has been proven to be nearly as effective as human judges [59].
+Previous works have demonstrated the potential of using LLMs to generate natural language descriptions of logging systems [60] or robots’ behaviors [61], indicating room for improvements to deliver more accurate and complete interpretations.
 Nevertheless, to our knowledge, no previous research has investigated the capabilities of LLMs in generating natural language explanations within the context of an accountability approach for record interpretation.
+2.5 Artificial Intelligence and Blockchain
+XAI and blockchain technology serve as complementary tools to enhance transparency, accountability, trust and acceptance of AI systems.
+In recent years, there has been a growing recognition of the importance of understanding and interpreting the decisions made by AI systems, particularly within critical domains such as healthcare, finance, and autonomous vehicles [62, 63, 64].
+XAI aims to address this need by providing human-understandable explanations for AI predictions and decisions, enabling users to trust and effectively utilize AI systems.
+However, achieving transparency and interpretability in AI models poses significant challenges, especially when dealing with complex algorithms and large datasets.
+Through blockchain, a decentralized and immutable ledger can record and verify the decision-making processes of AI models in a transparent and auditable manner.
+By integrating blockchain, XAI initiatives can store explanations for AI predictions on a distributed ledger, allowing stakeholders to validate and verify the reasoning behind AI decisions.
+A method for auditing legal decisions using blockchain technology alongside AI tools is proposed in [21].
+Using IPFS and blockchain technology, this approach offers a secure and tamper-proof record to store details, judgements, and reasoning.
+The use of Generative AI over this data helps to build trust in the system, by analyzing and finding relevant information.
+Data privacy concerns are also addressed, helping to build a more transparent, accountable, and efficient legal system.
+A framework for auditing XAI decisions using blockchain technology is proposed in [28].
+This framework includes features for both data storage and verification, with data stored in IPFS and the Ethereum blockchain preserving a unique integrity proof (hash) of the data.
+This setup allows anyone to verify the integrity of the data, thereby providing transparency and accountability in XAI decision-making processes.
+However, these approaches do not directly benefit the autonomous agents’ domain.
+Instead of relying on IPFS for storing contextual information for explanations, the use of Rosbag2 files, a binary file format, ensures the reproducibility of robots’ behaviors [42].
+Rosbags, capable of recording thousands of messages per second, offer a comprehensive and standardized format for capturing robotic data, guaranteeing the recording of each interaction and facilitating accurate analysis.
+Furthermore, given the inherent indeterminism of LLM algorithms, ensuring the integrity of the messages contained in the Rosbag file through a permissionless blockchain solution offers public transparency and accountability.
+By storing a chain of integrity proofs of these messages on the blockchain, their immutability is guaranteed, and their verification is simplified in every analysis of the Rosbag file, thereby improving the reliability of the data utilized in the explanation generation process.
 3 System Architecture
-Our approach comprises to main components: a tamper-proof black box-like device, and a module responsible for generating natural language explanations for the data stored by the former.
-The first component offers services for building, storing, and verifying integrity proofs through smart contract specifications.
-The explainability component provides natural language explanations by implementing Retrieval-Augmented Generation (RAG), thereby enhancing the obtained results using the data recorded by the first component as an external source to enrich the context in the answer formulation process.
+Our approach comprises two main components: a tamper-proof black box-like device, and a module responsible for generating natural language explanations for the data stored by the former.
+The first component offers services for building, storing, and verifying integrity proofs through SC specifications.
+The explainability component provides natural language explanations by implementing RAG, thereby enhancing the obtained results using the data recorded by the first component as an external source to enrich the context in the answer formulation process.
 Both components are integrated through the existence of two main asynchronous tasks focused on cleaning and processing the raw data generated by the black box-like device to increase the effectiveness of the explainability engine.
 Figure 1 depicts a generic architecture of our proposal, highlighting in blue color the functionalities related to accountability, and in green those related to explainability.
-Below we provide a deeper description of the approaches designed and followed in order to develop and integrate the above mentioned components.
+Below we provide a deeper description of the approaches designed and followed in order to develop and integrate the above-mentioned components.
 Figure 1: System architecture.
@@
 This fact not only enhances the granularity of stored message proofs but also provides a more adaptable solution.
-Our anti-tampering approach for this accountability proposal comprises two key components: a chain of integrity proofs coupled with the smart contract specification.
+Our anti-tampering approach for this accountability proposal comprises two key components: a chain of integrity proofs coupled with the SC specification.
 This section delves into their design and justifies how both align with the constraints of autonomous agents.
@@
 $h_{i}$ $=\text{hash}(h_{i-1}||Msg_{i})$
-By chaining every hash with the previous one, we optimize the validation of stored messages due the last linked digest should correspond to the final proof saved in the blockchain.
+By chaining every hash with the previous one, we optimize the validation of stored messages because the last linked digest should correspond to the final proof saved in the blockchain.
 Additionally, a service has been developed to generate the chain of hashes using the recorded data, ensuring partial validation of the integrity proofs without modifying the stored messages or adding non-essential information to the transactions saved in the blockchain.
@@
 3.1.2 Smart Contract
-Calling a smart contract function intended for blockchain writing involves the execution of a transaction that changes the state of the storage data on the blockchain.
+Calling an SC function intended for blockchain writing involves the execution of a transaction that changes the state of the storage data on the blockchain.
 In our approach, this transaction is signed with the account’s private key via the Elliptic Curve Digital Signature Algorithm (ECDSA), effectively identifying the signer.
-The defined smart contract also determines its ownership, implementing access control to ensure that only the contract owner can add or update message proofs.
+The defined SC also determines its ownership, implementing access control to ensure that only the contract owner can add or update message proofs.
 These features guarantee authentication and non-repudiation.
-Moreover, our smart contract includes a function to verify the existence of a message proof on the blockchain.
+Moreover, our SC includes a function to verify the existence of a message proof on the blockchain.
 This fact allows external parties to independently confirm whether a specific message hash has been stored in the contract and when it was stored.
@@
 Our approach is focused on improving this feature through the use of Natural Language Processing (NLP) technology, enabling non-expert users to interact with agents in a meaningful way.
-Our solution involves use of RAG to retrieve data from outside the language model (non-parametric) and augment the prompts by adding the information recorded from our accountability component in context.
+Our solution involves the use of RAG to retrieve data from outside the language model (non-parametric) and augment the prompts by adding the information recorded from our accountability component in context.
 Before this integration, the information undergoes processing by a dedicated module responsible for acquiring and processing the raw data generated by the autonomous agent, serving as contextual input in the question-answering system.
-Given the volume of messages, their varied nature, their length, the presence of non-meaningful records generated by the autonomous agent, and the need of establishing specific relationships between messages to enhance the context of the engine, this component extracts relevant information.
-Its aim is to improve the effectiveness of the explanations provided to non-expert users.
+Given the volume of messages, their varied nature, their length, the presence of non-meaningful records generated by the autonomous agent, and the need to establish specific relationships between messages to enhance the context of the engine, the log curation component extracts relevant information.
+This functionality serves to enhance the effectiveness of explanations provided to non-expert users.
+A more comprehensive description of this component is provided in Section 4.
 After processing the raw data stored by our accountability component, the ingestion process performs the chunking task to break down extensive text segments into more manageable units.
@@
 4.1 Black Box Recorder
-Robot Operating System (ROS) is the most commonly used framework in robotics, consisting of a set of software libraries and tools for building robot applications.
-However, challenges in meeting the requirements of real-time embedded systems, and the need to improve its multi-platform capabilities led to a major upgrade to ROS 2, whose main change lies in using the Data Distribution Service (DDS) (Maruyama et al., 2016).
+ROS is the most commonly used framework in robotics, consisting of a set of software libraries and tools for building robot applications.
+However, challenges in meeting the requirements of real-time embedded systems, and the need to improve its multi-platform capabilities led to a major upgrade to ROS 2, whose main change lies in using the Data Distribution Service (DDS) [65].
 Nodes constitute the fundamental building block in both ROS and ROS 2 applications.
@@
 Its main functionalities include planning, control, localization, visualization and other additional features.
-Furthermore, it uses behavior trees to create customized navigation behavior via orchestrating many independent modular servers, which implies high flexibility in the navigation task and the specification of complex robot behaviors (Macenski et al., 2020).
-Different independent task servers, intended to compute a path, control effort, or recovery, are communicated with the behavior tree (BT) over a ROS 2 interface, such as an action server or service.
-The use of BTs to generate explanations for robot behavior has been previously successfully evaluated, allowing seeking causal information answer questions (Han et al., 2021).
-Previous authors’ work describes the steps involved in the accountable information recording process (Fernández-Becerra et al., 2023).
+Furthermore, it uses behavior trees to create customized navigation behavior via orchestrating many independent modular servers, which implies high flexibility in the navigation task and the specification of complex robot behaviors [66].
+Different independent task servers, intended to compute a path, control effort, or recovery, are communicated with the behavior tree over a ROS 2 interface, such as an action server or service.
+The use of behavior trees to generate explanations for robot behavior has been previously successfully evaluated, allowing seeking causal information to answer questions [67].
 ROS 2 nodes publish data that other nodes consume through ROS 2 topics.
@@
 To ensure the immutability of generated Rosbag files, Ethereum has been selected as the blockchain framework.
-Ethereum, an opensource public blockchain, introduces the cryptocurrency Ether and a programming platform known as Smart Contract.
-Employing Proof of Stake (PoS) as a consensus protocol, Ethereum offers advantages in energy efficiency, scalability, security, decentralization, accessibility, and economic efficiency.
+Ethereum, an open-source public blockchain, introduces the cryptocurrency Ether and self-executing programs (SCs) for code execution on the blockchain.
+Ethereum’s open-source nature ensures decentralization and accessibility, allowing participation without needing approval, a feature crucial for maintaining the integrity of recorded data.
+Additionally, Ethereum’s use of Proof of Stake (PoS) as a consensus protocol enhances energy efficiency and throughput compared to traditional protocols such as Proof of Work (PoW).
+In addition, Ethereum’s scalability, security and economic efficiency make it a very secure option for ensuring data immutability.
 The high-level algorithms 1 and 2 depict the main steps in recording and proof verification services.
@@
 Our approach addresses challenges such as the drop in robot performance during the Rosbag recording process and costs associated with storing blockchain data.
-We accomplish this by saving chained hashes from messages at different intervals.
-These intervals can be established based on the relevance of messages from a topic, regarding their required immutability features.
+We accomplish this by saving chained hashes from messages at specific intervals, ensuring the integrity of data while preserving system performance.
+These intervals define how often an integrity proof is stored and can be established based on the relevance of messages from a topic, regarding their required immutability features.
 Furthermore, message proofs are stored in real-time on the Ethereum blockchain, eliminating the need to modify the content of the Rosbag file.
 Detection of any alteration to the messages within this file is facilitated by a service that generates the hash chain from the file at intervals set for each topic during the recording process.
-To verify the existence of any of these hashes in the blockchain, the blockchain logic has been extended by developing a Smart Contract (SC) using the Solidity language.
+To verify the existence of any of these hashes in the blockchain, the blockchain logic has been extended by developing an SC using the Solidity language.
 Solidity is an open-source programming language that facilitates the creation of decentralized applications, contracts, protocols, and agreements.
-This contract provides functionalities for storing and verifying proof information for hashes stored on the Ethereum blockchain by associating each hash value with the number of block where it is stored.
+This contract provides functionalities for storing and verifying proof information for hashes stored on the Ethereum blockchain by associating each hash value with the block number where it is stored.
 Verification of transaction ownership is also conducted during the storage of these proofs.
@@
 By analyzing messages from the recorded Rosbag topics, this service can identify key circumstances in the navigation process, such as the appearance of an obstacle or a change in the planned trajectory.
+The status of the navigation process is checked through the study of the ROS 2 action server NavigateToPose, used for commanding a robot to move to a specified pose.
+By default, Nav2 guarantees to find the shortest path under all conditions.
+The /plan topic includes the set of poses computed to reach a specific goal pose.
+If this initial path changes due to an external circumstance, the Euclidean distance between all the poses included in the pre-computed path will increase.
+This fact, along with the values obtained through the /scan topic, allows us to determine if the appearance of an obstacle has caused the route modification.
+These reasonings have been successfully checked in previous authors’ work [68] with the aim of building explanations for autonomous agents based on the use of ROS 2 topics.
+Additionally, the present approach uses the information provided by Nav2 behavior trees due to their potential in designing, visualizing, and understanding the behavior of autonomous robots.
+Other information such as the present position of the robot and its velocity is also processed through this component.
+A general overview of its behavior is described through the algorithm 3.
 As a result, through this component, ROS 2 developers are relieved from having to call the logging API to generate explanations.
-This functionality streamlines the development process, allowing developers to focus on core functionalities without need for managing log calls.
+This functionality streamlines the development process, allowing developers to focus on core functionalities without the need for managing log calls.
 This approach reduces debugging efforts and boosts efficient resource usage, scalability, and adaptability in the development lifecycle, contributing to a more productive natural language explanation generation process.
 An example of how raw ROS messages are transformed through this component is included in Figure 2.
-Rosbag Raw Messages action_msgs.msg.GoalStatusArray(status_list=action_msgs.msg.GoalStatus(goal_info= action_msgs.msg.GoalInfo( goal_id=unique_identifier_msgs.msg.UUID(uuid=array([190, 81, 48, 208, 248, 147, 78, 82, 188, 116, 202, 152, 203, 225, 149, 158], dtype=uint8)), stamp=builtin_interfaces.msg.Time(sec=27, nanosec=933000000)), status=2)]) nav2_msgs.msg.BehaviorTreeStatusChange(timestamp=builtin_interfaces.msg.Time(sec=1701354686, nanosec=130104968), node_name=’ComputePathToPose’, previous_status=’IDLE’, current_status=’RUNNING’) Processed Data 1701354686 Navigation to the goal number 1 has started.
+Algorithm 3 Log Curation Process
+1:$rosbag$, $btNodeDesc$
+2:$curationLog$
+3:for $message$ in $rosbag$ do
+4:     $msgTopic\leftarrow$getMsgTopic($message$)
+5:     $LogMsg\leftarrow message.data$
+6:     // Check a change in the planned path.
+7:     if $msgTopic$ == ’/plan’  then
+8:         $distance\leftarrow$ getEuclideanDistance($message.poses$)
+9:         if $distance>previousDistance$ DISTANCE_THRESHOLD then
+10:              $logMsg\leftarrow$ “Planned path has changed when trying to achieve goal pose number $nGoal$"
+11:              $previousDistance\leftarrow distance$
+12:              if min($last\_scan\_msg.ranges$) $<$ OBSTACLE_DISTANCE_THRESHOLD then
+13:                  $logMsg\mathrel{+}=$“ because there was an obstacle."
+14:              end if
+15:         end if
+16:     end if
+17:     // Check the navigation status.
+18:     if $msgTopic$ == ’/navigate_to_pose/_action/status’  then
+19:         $logMsg\leftarrow$ “No navigation is running."
+20:         $statusList\leftarrow$ /navigate_to_pose/_action/status
+21:         for $status$ in $statusList$ do
+22:              $currentGoal\leftarrow statusList.id$
+23:              if $currentGoal$ != $previousGoal$ and $currentGoal$ not in $finishedGoals$ then
+24:                  $logMsg\leftarrow$ “Navigation to a new goal has started."
+25:              else
+26:                  $logMsg\leftarrow$ “Navigation to the goal "
+27:                  if $status$ = EXECUTING then
+28:                       $logMsg\mathrel{+}=$ “is in progress.”
+29:                  else if $status$ = SUCCEEDED then
+30:                       $logMsg\mathrel{+}=$ “has succeeded."
+31:                  else if $status$ = CANCELED then
+32:                       $logMsg\mathrel{+}=$ “was cancelled."
+33:                  else if $status$ = ABORTED then
+34:                       $logMsg\mathrel{+}=$ “has been aborted."
+35:                  end if
+36:                  if $status$ = SUCCEEDED or $status$ = CANCELED or $status$ = ABORTED  then
+37:                       $finishedGoals\cup=currentGoal$
+38:                  end if
+39:              end if
+40:              // Check the executed Behavior Tree nodes and their status.
+41:              if $msgTopic$ == ’/behavior_tree_log’  then
+42:                  $logMsg\leftarrow$“Node " $message.NodeName$ “performing "$btNodeDesc$“is in status " $message.nodeStatus$
+43:              end if
+44:              // Gets robot’s position and orientation.
+45:              if $msgTopic$ == ’/amcl_pose’  then
+46:                  $logMsg\leftarrow$ “Position: " $message.x,message.y$ “Orientation: " $message.z,message.w$
+47:              end if
+48:              if $msgTopic$ == ’/cmd_vel’  then
+49:                  $logMsg\leftarrow$ “Linear velocity: " $message.x,message.y$ “Angular velocity: " $message.z$
+50:              end if
+51:         end for
+52:     end if
+53:     $logMsg\leftarrow message.timestamp$ + $logMsg$
+54:     $curationLog\leftarrow curationLog$ + $logMsg$
+55:end for
+56:return $curationLog$
+{mdframed} [hidealllines=true,backgroundcolor=gray!10,innerleftmargin=3pt,innerrightmargin=3pt,leftmargin=-3pt,rightmargin=-3pt] Rosbag Raw Messages action_msgs.msg.GoalStatusArray(status_list=action_msgs.msg.GoalStatus(goal_info= action_msgs.msg.GoalInfo( goal_id=unique_identifier_msgs.msg.UUID(uuid=array([190, 81, 48, 208, 248, 147, 78, 82, 188, 116, 202, 152, 203, 225, 149, 158], dtype=uint8)), stamp=builtin_interfaces.msg.Time(sec=27, nanosec=933000000)), status=2)]) nav2_msgs.msg.BehaviorTreeStatusChange(timestamp=builtin_interfaces.msg.Time(sec=1701354686, nanosec=130104968), node_name=’ComputePathToPose’, previous_status=’IDLE’, current_status=’RUNNING’){mdframed} [hidealllines=true,backgroundcolor=gray!10,innerleftmargin=3pt,innerrightmargin=3pt,leftmargin=-3pt,rightmargin=-3pt] Processed Data 1701354686 Navigation to the goal number 1 has started.
 1701354686 Navigation to the goal number 1 is in progress.
@@
 4.3 Explainability Component
-The output of the above described node serves as an external source in our implementation of the RAG approach.
+The output of the previously described node serves as an external source in our implementation of the RAG approach.
 This development is based on the use and extension of localGPT  [^1] https://github.com/PromtEngineer/localGPT, an open-source question-answering system designed to interact with documents conversationally while safeguarding the privacy of the underlying information.
@@
 Figure 3 depicts the main components in the RAG architecture used.
-Figure 3: Retrieval Augmented Generation arquitecture.
+Figure 3: Retrieval Augmented Generation architecture.
 Therefore, this implementation combines the strengths of robotics frameworks, blockchain technology, and natural language processing to provide a comprehensive solution for real-time recording, secure storage, and user-friendly interpretation of robotic system data.
@@
 4.4 Software and Hardware Artifacts
-Our solution has been deployed in a Linux server running Ubuntu 22.04 as Operating System (OS). The server is equipped with 6-core 2.6 GHz CPU, 64 GB of memory, and a 4 GB GPU.
+Our solution has been deployed in a Linux server running Ubuntu 22.04 as the Operating System (OS). The server is equipped with a 6-core 2.6 GHz CPU, 64 GB of memory, and a 4 GB GPU.
 The chosen ROS 2 version is Humble.
@@
 To provide a realistic scenario, we used a hospital simulation environment made available by Amazon Web Services (AWS) [^2] https://github.com/aws-robotics/aws-robomaker-hospital-world.
-Additionally, we selected RB-1, an autonomous mobile robot based on the ROS 2 platform produced by Robotnik (Guzmán et al., 2017).
+Additionally, we selected RB-1, an autonomous mobile robot based on the ROS 2 platform produced by Robotnik [69].
 This combination of tools and platforms enabled us to assess the impact and effectiveness of our approach.
-The interaction with Ethereum has been developed through web3.py library (version 6.6.1). Ganache (version 2.7.1) has been used as the Ethereum network simulator.
-RAG solution holds Langchain (version 0.325), ChromaDB (version 0.4.6).
-We used the open source model Llama 2 7B GGUF (Touvron et al., 2023) as the LLM responsible for generating natural language explanations for end users.
-This LLM integrates RLHF and shows versatility in handling general-purpose tasks, ranging from answering questions, text generation, and language translation to more demanding tasks such as reasoning or coding.
-Preliminary tests of Llama 2, compared to other LLMs such as Mistral (Jiang et al., 2023) or Zephyr (Tunstall et al., 2023), led us to choose this model for evaluating our data, due to the promising results obtained in these assessments conducted during the last quarter of 2023, when this work was in its development phase.
-As text embedding model we have chosen Instructor-xl (Su et al., 2022), a high-performance model designed to generate text embeddings tailored to tasks such as classification, clustering, text evaluation or retrieval, fact particularly important in the processes performed for any RAG solution.
+The interaction with Ethereum has been developed through the web3.py library (version 6.6.1). Ganache (version 2.7.1) has been used as the Ethereum network simulator.
+RAG solution holds Langchain (version 0.325), and ChromaDB (version 0.4.6).
+We used the open-source model Llama 2 7B GGUF [70] as the LLM responsible for generating natural language explanations for end users.
+This LLM integrates Reinforcement Learning from Human Feedback (RLHF) and shows versatility in handling general-purpose tasks, ranging from answering questions, text generation, and language translation to more demanding tasks such as reasoning or coding.
+Preliminary tests of Llama 2, compared to other LLMs such as Mistral [71] or Zephyr [72], led us to choose this model for evaluating our data, due to the promising results obtained in these assessments conducted during the last quarter of 2023, when this work was in its development phase.
+As a text embedding model we have chosen Instructor-xl [73], a high-performance model designed to generate text embeddings tailored to tasks such as classification, clustering, text evaluation or retrieval, a fact particularly important in the processes performed for any RAG solution.
 Table 2: Robag topics rate and average rate per message
@@
 This study covered several approaches for storing integrity proofs across various scenarios.
-Additionally, we explored and adapted the capabilities of RAG solutions and LLMs in question-answering systems, using as source the data gathered by our black box-like component, through a comprehensive set of questions.
-This work aimed to asses the effectiveness of these technologies in delivering natural language explanations to individuals who are not experts.
-The experiments conducted within this proposal focus on navigation tasks included in the Navigation Functionality of the European Robotics League Consumer Service Robots Challenge (ERL Consumer) (ERL, 2023).
+Additionally, we explored and adapted the capabilities of RAG solutions and LLMs in question-answering systems, using as a source the data gathered by our black box-like component, through a comprehensive set of questions.
+This work aimed to assess the effectiveness of these technologies in delivering natural language explanations to individuals who are not experts.
+The experiments conducted within this proposal focus on navigation tasks included in the Navigation Functionality of the European Robotics League Consumer Service Robots Challenge (ERL Consumer) [74].
 This functionality evaluates the capabilities of a robot to navigate accurately and autonomously within a given scenario, containing furniture, objects, and people.
@@
 In the initial scenario, the robot must navigate to three waypoints, encountering no obstacles along the way.
-The second scenario presents the robot with the challenge of adjusting its pre-planned path to reach the second goal.
+The second scenario presents the robot with the challenge of adjusting its previously planned path to reach the second goal.
 This adjustment becomes necessary due to the appearance of an obstacle obstructing the original planned route, requiring a deviation to achieve the goal.
@@
 In evaluating the accountability component, we conducted six experiments across three previously described scenarios.
-These experiments involved integrity proofs storage in the blockchain, with hash calculations occurring at intervals of one every 10, 25, 50, or 100 messages.
+These experiments involved storing integrity proofs on the blockchain, with hash calculations occurring at intervals of one every 10, 25, 50, or 100 messages.
 Moreover, our solution introduces the flexibility to save proofs of evidence at different rates based on the topic, ensuring non-tampering for critical topics or those with lower publishing rates where message loss could be crucial.
@@
 In particular, we have obtained a full report about system load, CPU, RAM, network traffic and disk write operations updated every second, facilitating the tracking of overall system activity and an easy comparison with preceding values.
-To describe the overall distribution and characteristics of these results, we calculated their average as statistical measure.
+To describe the overall distribution and characteristics of these results, we calculated their average as a statistical measure.
 This information is included in Section 6.
@@
 Despite the remarkable performance of LLMs across diverse language tasks, evaluating these models presents challenges, especially in ensuring factual consistency.
-The extensive capabilities of LLMs and the limitations of current benchmarks in capturing human preferences, have given rise to emerging trends in LLM evaluations, such as employing strong LLMs as judges (Zheng et al., 2023).
+The extensive capabilities of LLMs and the limitations of current benchmarks in capturing human preferences have given rise to emerging trends in LLM evaluations, such as employing strong LLMs as judges [59].
 Traditional similarity-based metrics, such as BLEU and ROUGE, have proven to misalign with human evaluation.
 While human evaluation remains the primary standard for assessing human preferences, it is exceptionally time-consuming and expensive.
-To automate the evaluation process, several works have adopted the ’LLM-as-a-judge’ approach, which focuses on using strong language models (LLMs) like GPT-4.
-These models are often trained with reinforcement learning from human feedback (RLHF), resulting in strong human alignment (Lin, 2024).
-Previous studies have evidenced the consistent agreement of these models with human grading in over 80% of judgments (Bubeck et al., 2023; Helgesson Hallström, 2023; Leng et al., 2023).
+To automate the evaluation process, several works have adopted the ’LLM-as-a-judge’ approach, which focuses on using strong LLMs like GPT-4.
+These models are often trained with RLHF, resulting in strong human alignment [75].
+Previous studies have evidenced the consistent agreement of these models with human grading in over 80% of judgments [76, 77, 78].
 Given these characteristics, we opted for GPT-4 to assess the answers generated by our explainability component.
-Through LLM-as-a-judge, we are able to obtain not only scores but also explanations, making their outputs interpretable.
+Through LLM-as-a-judge, we can obtain not only scores but also explanations, making their outputs interpretable.
 The whole process is represented in Figure 5.
@@
 LLMs exhibit a certain degree of indeterminism, defined as the model’s ability to offer different answers when asked the same question or prompt several times.
-This behavior derived from their training, in which the models learn to generate outputs by sampling from probabilistic distributions over possible word sequences, introducing randomness to their response.
+This behavior derives from their training, in which the models learn to generate outputs by sampling from probabilistic distributions over possible word sequences, introducing randomness to their response.
 The effect of this indeterminism must be considered in any LLM evaluation.
@@
 A basic RAG setup includes two main sources of error: the retriever and the LLM.
-The evaluation of the retriever aims at determining the relevance of its document selection in relation to a query.
-To this end, we conducted an experimental study running with 16 different questions over the data generated and processed during the navigation task in the three previously described scenarios.
+The evaluation of the retriever aims to determine the relevance of its document selection concerning a query.
+To this end, we conducted an experimental study with 16 different questions on the data generated and processed during the navigation task in the three previously described scenarios.
 This set of questions has been evaluated six times in each scenario, giving a total of 288 questions and answers to perform the analysis.
@@
 The final score, an average of all the scores in the dataset, is used to judge the retriever’s overall performance.
-A comprehensive evaluation of a LLM should consider multiple metrics that include correctness as well as other aspects.
+A comprehensive evaluation of an LLM should consider multiple metrics that include correctness as well as other aspects.
 A language model should produce text that is not only correct but also logically structured and easy to understand.
-Produced response must be appropriately detailed without unnecessary verbosity.
-Furthermore, the LLM must not only provide correct information but also delivers in a meaningful and valuable way to the user.
+Produced responses must be appropriately detailed without unnecessary verbosity.
+Furthermore, the LLM must not only provide correct information but also deliver it in a meaningful and valuable way to the user.
 Therefore, to assess the performance of the selected LLM on specific attributes, our approach also includes a criteria-based evaluation covering accuracy, conciseness, coherence, and helpfulness.
-Selected criteria metrics are describe in Table 3.
+Selected criteria metrics are described in Table 3.
 This structured assessment offers a means to assess the model’s effectiveness in generating responses, providing a broader understanding of the model’s capabilities and its alignment with the used scenarios.
@@
 Table 3: Criteria-based evaluation metrics
-Criteria metric
-Description
-Accuracy
-Accuracy evaluates the extent to which the response conforms to the reference, ranging from being completely unrelated to being completely accurate and in line with the reference.
-Helpfulness
-Helpfulness evaluates the degree to which the response provides valuable information, ranging from providing no helpful information to being extremely helpful and offering comprehensive information.
-Conciseness
-Conciseness assesses the brevity of the response, ranging from being extremely verbose to highly concise with no unnecessary words.
-Coherence
-Coherence examines the organizational structure and logical flow of the response, from extremely inconsistent to highly coherent, with a clear and logical organization.
-Question: How many goals have been reached by the robot?
-Scenario 1 Answer: Based on the provided context, we know that the robot has reached 3 goals.
+Criteria metric Description
+Accuracy Accuracy evaluates the extent to which the response conforms to the reference, ranging from being completely unrelated to being completely accurate and in line with the reference.
+Helpfulness Helpfulness evaluates the degree to which the response provides valuable information, ranging from providing no helpful information to being extremely helpful and offering comprehensive information.
+Conciseness Conciseness assesses the brevity of the response, ranging from being extremely verbose to highly concise with no unnecessary words.
+Coherence Coherence examines the organizational structure and logical flow of the response, from extremely inconsistent to highly coherent, with a clear and logical organization.
+{mdframed} [hidealllines=true,backgroundcolor=blue!20,innerleftmargin=3pt,innerrightmargin=3pt,leftmargin=-3pt,rightmargin=-3pt, skipabove=-3pt,skipbelow=-10pt] Question: How many goals have been reached by the robot?
+{mdframed}[hidealllines=true,backgroundcolor=blue!10,innerleftmargin=3pt,innerrightmargin=3pt,leftmargin=-3pt,rightmargin=-3pt] Scenario 1 Answer: Based on the provided context, we know that the robot has reached 3 goals.
 The context mentions that the goal number 1, 2 and 3 have been reached successfully.
-Therefore, the answer to your question is 3.
-GPT-4 Judgement Predicted Grade: CORRECT Reasoning: The AI assistant’s response is accurate, concise, coherent, and helpful.
+Therefore, the answer to your question is 3.{mdframed} [hidealllines=true,backgroundcolor=gray!10,innerleftmargin=3pt,innerrightmargin=3pt,leftmargin=-3pt,rightmargin=-3pt] GPT-4 Judgement Predicted Grade: CORRECT Reasoning: The AI assistant’s response is accurate, concise, coherent, and helpful.
 It directly answers the user’s question by confirming that the robot has reached three goals.
@@
 There are no unnecessary words or unclear statements, making the response concise and coherent.
-Accuracy: [[10]] Conciseness: [[10]] Coherence: [[10]] Helpfulness: [[10]]", score: 1.0 Scenario 2 Answer: Based on the context provided, the robot has reached 3 goals.
-GPT-4 Judgement Predicted Grade: CORRECT Reasoning: The assistant’s response is completely accurate as it matches the information given in the ground truth.
+Accuracy: [[10]] Conciseness: [[10]] Coherence: [[10]] Helpfulness: [[10]]", score: 1.0{mdframed} [hidealllines=true,backgroundcolor=blue!10,innerleftmargin=3pt,innerrightmargin=3pt,leftmargin=-3pt,rightmargin=-3pt] Scenario 2 Answer: Based on the context provided, the robot has reached 3 goals.{mdframed} [hidealllines=true,backgroundcolor=gray!10,innerleftmargin=3pt,innerrightmargin=3pt,leftmargin=-3pt,rightmargin=-3pt] GPT-4 Judgement Predicted Grade: CORRECT Reasoning: The assistant’s response is completely accurate as it matches the information given in the ground truth.
 The response is concise, with no unnecessary words or information.
@@
 The response is extremely helpful as it gives the exact information asked for by the user.
-Therefore, the ratings are as follows: Accuracy: [[10]] Conciseness: [[10]] Coherence: [[10]] Helpfulness: [[10]]", score: 1.0 Scenario 3 Answer: Based on the provided context,the robot has successfully reached 2 out of 3 goals.
+Therefore, the ratings are as follows: Accuracy: [[10]] Conciseness: [[10]] Coherence: [[10]] Helpfulness: [[10]]", score: 1.0{mdframed} [hidealllines=true,backgroundcolor=blue!10,innerleftmargin=3pt,innerrightmargin=3pt,leftmargin=-3pt,rightmargin=-3pt] Scenario 3 Answer: Based on the provided context,the robot has successfully reached 2 out of 3 goals.
 The successful goals are: * Goal number 2: Navigation to the goal number 2 has succeeded.
@@
 Therefore, the answer to the user’s question is: The robot has successfully reached 2 out of 3 goals.
-Goal numbers 2 and 3 have been successfully achieved, while goal number 1 has been aborted.
-GPT-4 Judgement Predicted Grade: CORRECT Reasoning: The AI assistant’s response is comprehensive and accurate.
+Goal numbers 2 and 3 have been successfully achieved, while goal number 1 has been aborted.{mdframed} [hidealllines=true,backgroundcolor=gray!10,innerleftmargin=3pt,innerrightmargin=3pt,leftmargin=-3pt,rightmargin=-3pt] GPT-4 Judgement Predicted Grade: CORRECT Reasoning: The AI assistant’s response is comprehensive and accurate.
 It correctly states that the robot has successfully reached goals number 2 and 3 while goal number 1 has been aborted.
@@
 To evaluate our explainability proposal in the context of ROS 2 messages interpretation, we have defined a set of questions covering diverse aspects of the navigation process over the how, the what, the which, the when and the where dimensions.
-These questions provide insights into different stages and components of the three scenarios.
+These questions provide insights into the different stages and components of the three scenarios.
 To structure and improve the understanding and analysis of this set of questions, we have organized them into five different categories.
 These categories range from high-level overviews of navigation to detailed inquiries about trajectory planning, goal completion, and the functioning of the behavior tree.
-The categories also delve into the causes of unexpected behavior or errors in navigation, improving the system explainability whenever an unexpected action occurs.
-Through these questions, we aim to address key concerns in XAI, including understanding why decisions are made, establishing causality between actions, determining success or failure, identifying timing, and disguising the methods by which achievements are realized (Gunning and Aha, 2019).
+The categories also delve into the causes of unexpected behavior or errors in navigation, improving the explainability of the system whenever an unexpected action occurs.
+Through these questions, we aim to address key concerns in XAI, including understanding why decisions are made, establishing causality between actions, determining success or failure, identifying timing, and disguising the methods by which achievements are realized [79].
 The set of categories, their descriptions and the questions included in each one are described next.
-Navigation Process Overview
+Navigation Process Overview.
 These questions provide an overview of the navigation process captured in the Rosbag file.
@@
     How has the navigation task proceeded?
-Trajectory Planning and Replanning
+Trajectory Planning and Re-planning.
 These questions delve into the specifics of trajectory planning, a critical aspect of navigation.
@@
   3. 5.
-    Did the robot find any obstacle during the navigation?
-Goal Completion and Navigation Task Status
+    Did the robot find any obstacles during the navigation?
+Goal Completion and Navigation Task Status.
 These questions focus on the completion status of the navigation goals.
@@
     Have all objectives been successfully achieved or have any been cancelled or aborted?
-Specifics about Goals and Locations
+Specifics about Goals and Locations.
 These questions examine the specifics of goals and locations.
@@
   4. 13.
-    What was the linear velocity of the robot after receiving the goal number 1?
-Nav2 Behavior Tree and Node Status
+    What was the linear velocity of the robot after receiving goal number 1?
+Nav2 Behavior Tree and Node Status.
 These questions provide information about the underlying behavior tree and the state of its nodes in the navigation system.
@@
 Concerning memory consumption, all approaches within the three scenarios exhibit comparable values, differing by no more than 0.5 GB.
-Consequently, the inclusion of blockchain approaches into our black box recorder does not yield a remarkable impact on RAM comsuption in comparison to avoiding the use of anti-tampering techniques in this recording process.
+Consequently, the inclusion of blockchain approaches into our black box recorder does not yield a remarkable impact on RAM consumption in comparison to avoiding the use of anti-tampering techniques in this recording process.
 Regarding network traffic, approaches with higher process storage in the blockchain increase this metric, but do not exceed 0,20 MB/s, which is a very affordable quantity.
-Furthermore, approaches such as storing one hash every 100 messages nearly match the avoidance of storing any proof into the blockchain, indicating that including this approach do not have a remarkable impact in the system performance.
+Furthermore, approaches such as storing one hash every 100 messages nearly match the avoidance of storing any proof into the blockchain, indicating that including this approach does not have a remarkable impact on the system performance.
 Disk write operations do not form a bottleneck in system performance, with all approaches differing by no more than 0,40 MB/s, which is not a very significant impact.
-In addition, all the approaches that perform integrity proof storage on the blockchain achieve lower or similar results to the approach avoiding anti-tampering operations.
+In addition, all the approaches that perform integrity-proof storage on the blockchain achieve lower or similar results to the approach avoiding anti-tampering operations.
 Figure 9 shows the results described.
@@
 Figure 10 shows the results of the correctness of RAG evaluation, presenting both the overall results and assessments based on predefined categories.
-The results demonstrate remarkable consistency across the three scenarios, getting a significant number of correct answers, even though there has not been a prior tuning in the ROS 2 messages for autonomous agents navigation tasks.
+The results demonstrate remarkable consistency across the three scenarios, getting a significant number of correct answers, even though there has not been a prior tuning in the ROS 2 messages for autonomous agents’ navigation tasks.
 The slight percentage of incorrect answers remains within an acceptable threshold, given the specific context employed and the inclusion of general questions.
@@
 In this instance, the cancellation of a goal resulted in the execution and failures of Nav2 Behavior Tree nodes, not included in the other two scenarios and focused on navigation recovery.
-Therefore, the incorporation in a dataset of information related to the cancellation of a goal, results in LLM asnwers that usually diverge from the ground truth provided, as the model occasionally prioritized Behavior Trees information, other times focused on locations or timestamps.
+Therefore, the incorporation in a dataset of information related to the cancellation of a goal results in LLM answers that usually diverge from the ground truth provided, as the model occasionally prioritized Behavior Trees information, other times focused on locations or timestamps.
 This variability made it difficult to align with the ground truth provided, even if the obtained answers could be considered partially accurate.
-Introducing the prompt extension "Describe the sequence of events regarding navigation goals and include information about each of them." in question 1 significantly increased correctness in the first category, achieving results around and exceeding the 75% in all three scenarios.
+Introducing the prompt extension "Describe the sequence of events regarding navigation goals and include information about each of them." in question 1 significantly increased correctness in the first category, achieving results around and exceeding 75% in all three scenarios.
 This fact underscores the effectiveness of prompt engineering when tackling ambiguous or broad queries.
 The results also show consistently high correctness across all scenarios concerning questions related to trajectory planning and re-planning.
-The model is able to determine whether the modification of a predicted trajectory has been caused by the occurrence of an obstacle or other circumstance.
+The model can determine whether the modification of a predicted trajectory has been caused by the occurrence of an obstacle or other circumstance.
 Scenario 1 and Scenario 2 also exhibit relatively high correctness in questions related to goal completion and task status.
-However, the model encounters challenges in Scenario 3, occasionally failing to identify if a navigation that starts correctly concludes with an abortion or cancellation.
+However, the model encounters challenges in Scenario 3, occasionally failing to identify whether a navigation that starts correctly concludes with an abortion or cancellation.
 This issue may be attributed to misalignment during the retrieval process, specifically in the selection of the best chunk associated with a particular goal associated with numerous embeddings within a single collection from the vector database.
 Correctness in identifying specific targets and locations lacks some consistency, with results exceeding 50% in the first scenario, around 70% in the second and over 75% in the third.
-In particular, three scenarios encounter difficulties in identifying the initial position of the robot before starting the navigation, due to the model misaligns the queried position with the one assumed by the robot when starting the navigation process.
+In particular, three scenarios encounter difficulties in identifying the initial position of the robot before starting the navigation, as the model misaligns the queried position with the one assumed by the robot when starting the navigation process.
 This fact may also be attributed to deficiencies in the retrieval process.
-In Scenario 1, understanding the sequence of events or the order of goals results challenging for the model, leading to inaccuracies in specifying the second location or goal.
-Nevertheless, this challenge is not present in Scenario 3 and absent or reduced in Scenario 2, demonstrating how differences in data distribution can introduce varied responses even when using specific and unambiguous questions.
-Through questions included in the Nav2 Behavior Tree and node status category, we assessed the effectiveness of the model’s formal linguistic competence compared to its functional linguistic competence, a feature of present LLMs (Mahowald et al., 2023).
+In Scenario 1, understanding the sequence of events or the order of goals is a challenge for the model, leading to inaccuracies in specifying the second location or goal.
+Nevertheless, this challenge is not present in Scenario 3 and is absent or reduced in Scenario 2, demonstrating how differences in data distribution can introduce varied responses even when using specific and unambiguous questions.
+Through questions included in the Nav2 Behavior Tree and node status category, we assessed the effectiveness of the model’s formal linguistic competence compared to its functional linguistic competence, a feature of present LLMs [80].
 While the model excels in identifying Behavior Tree Nodes given its description, it faces challenges in pinpointing these components under specific conditions.
 As a result, the category correctness suffers a drop in Scenario 3, although overall results remain above 75%.
-In this scenario, the navigation recovery actions caused by the cancellation of the first goal introduced a challenge in the interpretation of the information by the LLM due to the execution of several Nav2 Behavior Tree nodes not present in the others scenarios.
+In this scenario, the navigation recovery actions caused by the cancellation of the first goal introduced a challenge in the interpretation of the information by the LLM due to the execution of several Nav2 Behavior Tree nodes not present in the other scenarios.
 Particularly, the mistakes stem from the model’s confusion in determining whether a previously failed Nav2 Behavior Tree node has recovered or finished its execution, as well as identifying all nodes that were executed during navigation.
@@
 6.4 Criteria-based Evaluation
-Figure 11 shows the remarkable performance of the Large Language Model’s (LLM) in the evaluated metrics, demonstrating consistency with slight variations in the scores.
+Figure 11 shows the remarkable performance of the selected LLM in the evaluated metrics, demonstrating consistency with slight variations in the scores.
 On average, the answers provided are accurate and in line with the reference, though there might be room for improvement.
@@
 The language model’s adaptability to challenges like obstacles and goal cancellations regarding the described categories, is shown in Figure 12.
-Figure 12: Large Language Model criteria-based evaluation by categories. Comparative visualizations for: (a) Scenario 1, (b) Scenario 2, and (c) Scenario 3.
-Concerning general questions about about events and the progression of the navigation task, we observe variations in scores across Scenarios 2 and 3.
+Figure 12: LLM criteria-based evaluation by categories. Comparative visualizations for: (a) Scenario 1, (b) Scenario 2, and (c) Scenario 3.
+Concerning general questions about events and the progression of the navigation task, we observe variations in scores across Scenarios 2 and 3.
 While coherence and conciseness are even improved, Scenario 1, involving direct navigation to three goals, outperforms in accuracy and helpfulness.
 This fact suggests that effective communication on general queries benefits from the absence of obstacles or goal cancellations, enhancing overall understanding.
-In the category ’Trajectory Planning and Replanning’, the questions are effectively addressed in all the three scenarios, showing scores consistently above 8 in every evaluated metric, even when obstacles are introduced or a goal is cancelled.
-These results underlines the system’s proficiency in handling trajectory-related questions despite potential new challenges.
+In the category ’Trajectory Planning and Re-planning’, the questions are effectively addressed in all three scenarios, showing scores consistently above 8 in every evaluated metric, even when obstacles are introduced or a goal is cancelled.
+These results underline the system’s proficiency in handling trajectory-related questions despite potential new challenges.
 Regarding the category ’Goal Completion and Navigation Task Status’, Scenarios 1 and 2 demonstrate the effectiveness of the system in conveying goal completion and navigation task status information.
 Nevertheless, the slightly lower scores for Scenario 3 indicate potential challenges in effectively communicating completion status when a goal is aborted.
-Concerning the ’Specifics about Goals and Locations’ category, the system consistently provide valuable goal and location information, even in the presence of obstacles or goal cancellations.
-This suggests a robust capacity to understand and localize specific details regardless of challenges introduced in each scenario.
+Concerning the ’Specifics about Goals and Locations’ category, the system consistently provides valuable goal and location information, even in the presence of obstacles or goal cancellations.
+This suggests a robust capacity to understand and localize specific details regardless of the challenges introduced in each scenario.
 The results regarding questions included in ’Nav2 Behavior Tree and Node Status’ demonstrate effectiveness in communication across all scenarios and metrics.
@@
 7 Conclusions
-This work presents an accountability and explicability architecture, implemented for ROS-based mobile robots.
+This work presents an accountability and explicability architecture, implemented for Robot Operating System (ROS)-based mobile robots.
 Our approach comprises a black box component for accountability that effectively addresses real-time recording challenges, ensuring message integrity and verifiability through blockchain technology, and an explainability component.
-This last component is based on the use of RAG and LLM through a private question-answering system designed to use the information recorded by the first component as context.
+This last component is based on the use of Retrieval Augmented Generation (RAG) and Large Language Models (LLMs) through a private question-answering system designed to use the information recorded by the first component as context.
 This solution demonstrates the feasibility of obtaining accurate, coherent, concise and helpful explanations for robot actions from raw data stored in Rosbag files, thereby enhancing human-robot interaction.
 Additionally, both main components have been integrated through the development of a third main functionality in charge of filtering non-relevant information and making messages closer to the final user, due to their specific nature can introduce challenges in their comprehension process.
-Furthermore, this last component releases software developers from making direct calls to the logging API, easing the development process and improving the scalability and adaptability of our solution for every navigation task performed by a ROS-based robot.
+Furthermore, this last component releases software developers from making direct calls to the logging Application Programming Interface (API), easing the development process and improving the scalability and adaptability of our solution for every navigation task performed by a ROS-based robot.
 The results achieved highlight the capability of our solution to guarantee the immutability of information generated by an agent without compromising its performance, a feature particularly important when trying to identify the causes behind an event.
@@
 Additionally, our RAG approach could improve in terms of retrieving the most relevant chunk of information from the context.
-This enhancement may involve exploring the use of separate collections within the vector database or employing a hybrid approach, such as integrating knowledge graphs (KG) with vector indexes.
+This enhancement may involve exploring the use of separate collections within the vector database or employing a hybrid approach, such as integrating Knowledge Graphs with vector indexes.
+Furthermore, by incorporating real-time explanation generation mechanisms into our solution, the responsiveness and adaptability of the system could be improved.
 These approaches can help to improve the accuracy, quality and overall consistency of the answers obtained across several scenarios, thereby enhancing the performance of our proposed system.
+Appendix A Supplementary Materials
+A.1 Prompt Template
+[…]
+Figure 13: Prompt template used to generate explanations with Llama 2 LLM.
+A.2 Explanations Obtained
+The 288 questions and answers obtained through the explanation generation and evaluation process are available on the aforementioned mentioned Github repository  [^4] https://github.com/laurafbec/immutable_explainable_BBR/blob/main/
+gpt4_eval_results/LLMs_QA_Results.xlsx
+.
+Additionally, a Python notebook has been developed to interact with the content of the previous sheet.
+This notebook is available on  [^5] https://github.com/laurafbec/immutable_explainable_BBR/blob/main/
+gpt4_eval_results/GPT4 Evaluation Results.ipynb
 […]
@@
 Data Availability
-Data and tools are available in a public repository.
+Data and tools are available in a public repository  [^6] https://github.com/laurafbec/immutable_explainable_BBR.
 Acknowledgments

### v3

@@
 Second, a component in charge of generating natural language explanations by harnessing the capabilities of Large Language Models (LLMs) over the data contained within the previously mentioned black box.
-The integration of blockchain and generative Artificial Intelligence (AI) solutions ensures the deployment of responsible, transparent, interpretable, and trustworthy agents.
 The study evaluates the performance of our solution in three different scenarios, each involving autonomous agent navigation functionalities.
@@
 Security improvements for autonomous agents in public spaces.
-Review of blockchain trends for immutable logs, EDRs, XARs, and LLMs.
+Review of blockchain for immutable logs, EDRs, XAI and blockchain, XARs, and LLMs.
 Accountability and Explainability architecture for autonomous agents.
@@
 1 Introduction
-The rising presence of mobile robots alongside humans has increased security and safety concerns.
-One of the main challenges in this context is the ability of robots to justify their actions and decisions, particularly in safety-critical scenarios [1].
-Therefore, it is essential to ensure that the events performed when developing an action or behavior are reliable and transparent [2].
-Equally important is the capacity to render these explanations comprehensible to non-expert users, offering them appropriate, effective, and natural responses to their queries [3, 4].
-The most substantial evidence in post-event analysis relates to the logging and auditing capabilities, which are essential to mitigate technical issues and provide forensic evidence to detect any data tampering or unauthorized leakage.
-Therefore, a trustworthy and secure log system is needed when generating a complete trace of a robot’s behavior [5].
-The system must ensure the integrity, confidentiality, and availability of the agent and the data.
-Additionally, it should possess the capability of incident detection, facilitating investigations, and ensuring compliance, ultimately improving the overall security of the robot.
-Regarding log systems vulnerabilities, covering tracks has become a common practice for attackers, erasing any traces that could lead to their identification or assisting in the analysis of their actions.
+The rising presence of mobile robots alongside humans has increased concerns about security and trust.
+One of the main challenges in this context is the ability of robots to justify their actions and decisions, particularly in safety-critical scenarios where their behavior may impact human well-being [1].
+Transparency and reliability in these justifications are essential [2], as such properties enable effective human-robot interaction and foster confidence, particularly for non-expert users who require clear and understandable explanations [3, 4].
+However, the complexity of robotic decision-making and the need for thorough post-event analysis make achieving these goals particularly difficult.
+To address these challenges, accountability must be ensured through a robust, tamper-proof logging system.
+The most substantial evidence for post-event analysis comes from logging and auditing capabilities, which are essential to mitigate technical issues and provide forensic evidence to detect data tampering or unauthorized breaches.
+These systems must guarantee the integrity, confidentiality, and availability of data while enabling incident detection, compliance, and forensic investigations [5].
+However, covering tracks in log systems has become a common practice for attackers, erasing any traces that could lead to their identification or assisting in the analysis of their actions.
 As a result, it is necessary to strengthen system properties against unauthorized modifications [6].
-This fact not only includes safeguarding the integrity of log data but also implementing measures to detect and counteract attempts at tampering.
-Accountability is equally required, enabling the detection and isolation of faulty behaviors and their origin.
-It must not only provide tamper resistance but also offer verifiability for checking the integrity of log entries against tampering.
-Moreover, it should guarantee search accessibility, enabling access to specific records whose integrity can be verified [7].
+This includes safeguarding the integrity of log data and implementing measures to detect and counteract attempts at tampering.
+Effective accountability solutions must ensure the detection and isolation of faulty behaviors and their origin.
+Moreover, they should guarantee search accessibility, enabling access to specific records whose integrity can be verified [7].
 Integrating accountability solutions into robotic systems often demands a trade-off with system performance.
-Specifically, features such as computational power and networking bandwidth can be negatively affected.
-The inclusion of anti-tampering techniques, such as full disk and communication encryption, may be restrictive due to resource limitations [8].
+Specifically, computational power and networking bandwidth may be adversely affected [8].
+The inclusion of anti-tampering techniques, such as full disk and communication encryption, can be restrictive due to resource limitations [9].
 Additionally, high-cost tamper-proof storage devices may prove economically unviable, especially when dealing with continuously generated log data.
-Therefore, it is necessary to establish an efficient and verifiable proof of integrity to guarantee the reproducibility and identification of the events that could arise during an autonomous agent action.
-For this evidence to be significant, it should belong to a persistent chain of custody, totally reliable and verifiable, main requirements for an auditable and transparent logging system [9].
-Blockchain solutions emerge as a mechanism to ensure data integrity, protection against unauthorized changes, and traceability.
-The consensus protocols established by all blockchain network members enable the detection and rejection of any destruction, manipulation, or non-permitted editing of data through illicit channels.
-The nature of this technology helps to avoid tampering risk, safeguarding data [10].
-Nevertheless, the potential evidence obtained poses significant challenges for analysis due to its diverse and varied nature, terminology, and the presence of substantial volumes of similar messages unrelated to anomalous events.
-These facts may result in limited usefulness and misalignment with the expectations of non-expert users [11].
-Regarding autonomous agents, the raw log messages produced by these systems are usually noisy and possess a semi-structured format.
+Therefore, it is necessary to establish an efficient and verifiable proof of integrity to guarantee the reproducibility and identification of the events that could arise during an autonomous agent action without compromising the robot’s performance.
+For this evidence to be significant, it should belong to a persistent chain of custody, fully reliable and verifiable, key requirements for an auditable and transparent logging system [10].
+To this end, blockchain solutions emerge as a mechanism to ensure data integrity, protection against unauthorized changes, and traceability.
+The consensus protocols established by all blockchain network members enable the detection and rejection of any destruction or manipulation of data through illicit channels.
+The nature of this technology mitigates tampering risk, safeguarding data [11].
+Nevertheless, while blockchain secures log data, the evidence obtained poses significant challenges for analysis due to its diverse and varied nature, terminology, and the presence of substantial volumes of similar messages unrelated to anomalous events.
+These factors may result in limited usefulness and misalignment with the expectations of non-expert users [12].
+The raw log messages produced by autonomous agents are usually noisy and semi-structured in format.
 In addition to this, it is necessary to deal with specific features such as large quantities of domain-specific language messages [13].
-In this particular context, explaining the reasons that make a robot carry out a specific and unexpected action is essential for diagnosing failures in goal achievement.
-Besides, making these explanations accessible to humans is equally important [13].
-This requirement has led to the development of eXplainable Autonomous Robots (XAR), which aims to provide mechanisms for presenting clear and understandable explanations for the actions and decisions of autonomous agents.
-XAR enhances the transparency and interpretability of these agents’ decision-making processes, often incorporating techniques like natural language explanations to improve clarity [14, 15].
-Large Language Models (LLMs) have gained significant attention due to their exceptional performance in diverse natural language processing tasks, including translation, coding or text summarization [16, 17, 18].
-One approach to improve the results of LLMs for a specific domain is through the use of retrieval-based methods.
-By retrieving relevant information from a large corpus of text, these models can provide explanations grounded in existing knowledge [19].
-The integration of Artificial Intelligence (AI) and blockchain has received significant attention in recent years, with researchers exploring various ways of combining these two areas to provide efficiency, security, reliability and transparency to different domains [20, 21].
-Our contribution demonstrates how bringing these technologies together can improve the behavioral analysis of autonomous agents, enhancing their interactions with humans and thus increasing their acceptance.
-This work presents an accountability and explainability architecture implemented for ROS-based mobile robots.
+Addressing this issue requires methods to organize and filter log data, making it accessible and meaningful.
+In this context, explaining the reasons behind a robot’s specific and unexpected actions is essential for diagnosing failures and comprehending deviations from intended goals.
+Equally important is ensuring these explanations are understandable to users [14].
+Explainable Artificial Intelligence (XAI) is a research area dedicated to enhancing the transparency and interpretability of Artificial Intelligence (AI) systems to ensure their responsible and ethical deployment.
+As autonomous robots grow in complexity, understanding and predicting their behavior has become increasingly challenging for non-expert users.
+This requirement has led to the development of eXplainable Autonomous Robots (XAR), which aims to provide mechanisms for presenting clear and understandable explanations for the actions and decisions of autonomous agents [15].
+XAR enhances the transparency and interpretability of these agents’ decision-making processes, often incorporating techniques like natural language explanations to improve clarity and user comprehension [16, 17].
+To further improve the effectiveness of XAR systems, Large Language Models (LLMs) excel in diverse natural language processing tasks, including translation, coding, or text summarization [18, 19, 20].
+Our approach integrates LLMs with blockchain-secured log data.
+By using retrieval-based techniques, this solution generates accurate and context-aware natural language explanations based on verifiable information [21].
+This integration ensures that explanations are clear, coherent, and grounded in a traceable and reliable context, effectively addressing user comprehension and data integrity concerns.
+This work presents an accountability and explainability architecture implemented for Robot Operating System (ROS)-based mobile robots.
 This solution includes two main components.
@@
 The second component generates natural language explanations from data in the black box, including the use of interactive natural language conversations as a means to provide comprehensive explainability for the recorded events.
-Both components are connected through an additional component responsible for processing raw data collected by the first one, eliminating non-essential information and building connections and relations between data.
+Both components are connected through an additional module responsible for processing raw data collected by the first one, eliminating non-essential information, and building connections and relations between data.
 These functionalities release software developers from the need to make calls to the logging Application Programming Interface (API) to build explanations for the code generated and the behaviors developed by the autonomous agent.
-Additionally, we have conducted a thorough evaluation of our solution, including its impact on system performance and the correctness, accuracy, helpfulness and coherence of the generated natural language explanations.
+To validate our approach, we conducted a comprehensive evaluation, assessing the architecture’s impact on system performance and analyzing the quality of the generated explanations.
+The evaluation considered key metrics, including correctness, accuracy, helpfulness, coherence, and user satisfaction.
+Our results demonstrate that the architecture effectively balances accountability and explainability without compromising the robot’s performance, providing a scalable and reliable framework for real-world applications.
 The remainder of this paper is organized as follows: Section 2 provides an overview of previous works and findings related to the study, focusing on the enhancement of immutable data recorders and explainability solutions for robots.
@@
 In this section, we explore key works related to various domains relevant to our approach.
-It includes progress in understanding how blockchain ensures the immutability of logs, research on the application of event data recorders in robotics, and significant contributions XAR field.
-Lastly, we examine the evolution of LLMs to provide natural language explanations.
+This includes advances in how blockchain ensures the immutability of logs, research on applying Event Data Recorders (EDRs) in robotics, significant contributions to integrating XAI with blockchain, and XAR.
+Finally, we examine the evolution of LLMs for generating natural language explanations and their advancements within robotics.
 2.1 Blockchain for Immutable Logs
 Blockchain is a decentralized solution that eliminates the necessity for trust in a central authority through its immutable and distributed ledger, composed of timestamped transaction blocks.
-These blocks are linked through hashing and storage of previous block content, ensuring the integrity of transactions.
+These blocks are linked through hashing and the storage of previous block content, ensuring the integrity of transactions.
 Any attempt to alter a transaction in one block requires the modification of all subsequent blocks, resulting in a computationally expensive process.
@@
 Nevertheless, closed networks may pose security risks, as authorized users could recalculate subsequent blocks in the event of a block modification [24].
-While permissionless blockchains have traditionally related to scalability and throughput challenges, approaches like Ethereum have transitioned to Proof of Stake (PoS), a more energy-efficient, secure, accessible and scalable consensus protocol [25, 26].
+While permissionless blockchains have traditionally related to scalability and throughput challenges, approaches like Ethereum have transitioned to Proof of Stake (PoS), a more energy-efficient, secure, accessible, and scalable consensus protocol [25, 26].
 Blockchain technology ensures log integrity and auditability, offering unique advantages over alternatives such as distributed storage systems or InterPlanetary File Systems (IPFSs), which may require central authorities for storage and verification [27, 28].
@@
 Nevertheless, inherent features of permissionless blockchain such as transparency and trust, have also boosted the development of secure logging systems [34].
-2.2 Robotic Event Data Recorders
-Event Data Recorders (EDRs), often referred to as black boxes, have traditionally played a crucial role in capturing and analyzing events within robotics systems.
-These recorders support the collection and storage of event data, helping to comprehend system behavior, detecting anomalies, and ensuring safety [35, 36].
-Therefore, these solutions are needed in forensic investigations, ethical assessments and transparency regarding autonomous agents [37].
-The integration of black box modules to enhance robots’ introspection capabilities contributes to improving system transparency by attributing responsibilities [38].
-These solutions typically include functionalities such as hashing, signing, and encryption [39].
-However, to ensure the integrity and authenticity of recorded events, anti-tamper features must also be in place [40].
-Using blockchain technology in event data recorders can guarantee the integrity of a transparent and auditable record of events.
-These features enhance accountability and trust in the recorded data, as it becomes difficult for any party to manipulate or tamper with the recorded events without detection [41].
-Additionally, blockchain involves decentralized and distributed storage of event data, eliminating the necessity for central authorities or intermediaries and reducing the risk of data loss.
+Our architecture uses the immutability of blockchain to store evidence of the integrity of recorded data, thus providing real-time guarantees against manipulation and ensuring that the explanations generated by our system are always based on non-manipulated data.
+2.2 Blockchain for Event Data Recorders in Robotics
+EDRs, or black boxes, are essential for capturing and analyzing events in robotic systems.
+These recorders help to comprehend system behavior, detect anomalies, and ensure safety [35, 36].
+Therefore, these solutions are needed in forensic investigations and ethical assessments regarding autonomous agents [37].
+The integration of black box modules to enhance robots’ introspection capabilities contributes to improving system transparency by attributing responsibilities [38, 9].
+These solutions typically include functionalities such as hashing, signing, and encryption [39], while ensuring event integrity through anti-tamper mechanisms [40].
+Using blockchain technology in EDRs ensures the integrity of a transparent and auditable record of events.
+These features enhance accountability and trust in the recorded data by preventing tampering [41].
+Decentralized and distributed storage further eliminates reliance on central authorities and mitigates data loss risks.
 Further research explores the application of blockchain technology to improve robotic capabilities, particularly event data recording and management.
-The Black Block Recorder (BBR) system proposed in [42], employs permissioned blockchain technology for immutable black box logging in robotic applications.
-The BBR system combines Rosbag2 and Distributed Ledger Technologies (DLTs) to ensure data integrity and immutability.
-The potential applications of blockchain in robot swarms, including secure communication, data logging, and consensus agreement are discussed in [43].
-In [44], a time-segmented consortium blockchain for robotic event registration is proposed, organizing and storing event records efficiently to allow for retrieval and analysis of specific events within specific time frames.
-Authors in [45] present a tool to bridge Robot Operating System (ROS) and the Ethereum blockchain, evaluating encryption efficiency and stability, transactions response time and packet loss rate.
-Nevertheless, these works lack capabilities to save integrity proofs in the blockchain at selected intervals, to enhance granularity, or to include service definitions or SCs to improve the verifiability process.
+The Black Block Recorder (BBR) system uses permissioned blockchain for immutable logging, combining Rosbag2 and Distributed Ledger Technologies (DLTs) to ensure data integrity [42].
+Blockchain applications in robot swarms, including secure communication, data logging, and consensus agreements, are discussed in [43].
+In [44], a time-segmented consortium blockchain for robotic event registration is proposed, enabling efficient event organization and retrieval within specific time frames.
+Authors in [45] present a tool to bridge ROS and the Ethereum blockchain, evaluating encryption efficiency and stability, transaction response time, and packet loss rate.
+Nevertheless, these works lack the capability to save integrity proofs in the blockchain at selected intervals, enhance granularity, or include service definitions or SCs to improve the verifiability process.
+2.3 Explainable Artificial Intelligence and Blockchain
+XAI aims to align AI decisions with human cognitive expectations by employing techniques such as contextual relationships, contrastive explanations, and methods like counterfactual reasoning and interpretable models that enhance interpretability and transparency [46, 47].
+Insights from social sciences show people tend to interpret explanations more positively when they reflect human-like reasoning, emphasizing intentions, goals, and contextual factors rather than purely technical or statistical descriptions [48].
+By combining XAI with blockchain transparent record-keeping, explanations can be retraced and validated, ensuring both interpretability and accountability in AI-driven systems [49, 50].
+The benefits of this integration are remarkable in various domains.
+In healthcare, XAI and blockchain have been combined for secure, transparent patient data management [51], logical disease prediction, and immutable medical diagnostics records [52].
+Similarly, in the legal domain applications, blockchain and XAI support tamper-proof decision auditing, using IPFS and Ethereum for cryptographic proofs that enable data verification and accountability [53, 54, 28].
+Privacy-preserving systems with homomorphic encryption and semantic ontologies further ensure data protection while maintaining explainability [55].
+Regarding autonomous agents like unmanned aerial vehicles or autonomous vehicles, blockchain and XAI ensure secure and transparent decision-making [56].
+This integration enhances collaboration, even in complex and time-critical scenarios [57].
+Decentralized edge models for collaborative AI, leveraging blockchain for secure communication and real-time data sharing, further optimize resources and enhance reliability in these environments [58].
+Our work introduces Rosbag2 files as an alternative to IPFS or distributed file systems for storing contextual data in autonomous robotic systems.
+This approach supports reproducibility and enables a more precise analysis of robotic behaviors [42].
+Rosbag files provide a standardized format for recording large volumes of robotic data, facilitating comprehensive interaction capture for improved analysis.
+To address the indeterminism of LLM algorithms, we propose using a permissionless blockchain to store integrity proofs of Rosbag messages, ensuring immutability, simplifying verification, and enhancing data reliability for explanation generation.
 2.4 Explainable Autonomous Robots
-Explainable Artificial Intelligence (XAI) is a research area focused on improving the transparency and interpretability of AI systems, thereby promoting their responsible deployment.
-The need for explainability in autonomous robots arises from the increasing complexity and autonomy of these systems, making it challenging for users to comprehend their behavior.
-Explainable Autonomous Robots (XARs) are designed to provide explanations for their actions and decisions to non-expert users.
-Such explanations enrich aspects like security and the trust level, preventing failures, errors, or conflicts in achieving objectives or undesired interactions [46].
-Despite the growing demand for explainability in autonomous agents, their use reveals a lack of practical and effective usability or interpretability in real environments [47].
-Therefore, it is a problem that requires special attention, being necessary to explain the performed decisions.
-This explanation should include alternatives, choices made, expectations of each option, decision criteria, and any unexpected events that may have led to the modification of the initially planned behavior [15, 48].
-As described in [49], an explanation regarding XAR is an "abductive reasoning that bridges the gap between the individual explaining, which in this case may be a robot, and the person to whom they are explaining".
-Self-explanatory capabilities in social robots have been provided through the development of architectures and frameworks [50, 51].
-Nevertheless, the essence of explainability regarding XAR lies in effective communication [49].
-Progress in natural language processing (NLP) technology through LLMs, based on the estimation of human statements and the relevance and position of the words contained in them, could improve the efficiency of human-robot interactions.
-This enhancement can create a sense that these systems truly comprehend and respond to individuals meaningfully, significantly benefiting their usefulness and applicability [1].
-2.4 Large Language Models
-LLMs have recently played a significant role in natural language processing, thanks to replacing recurrent neural networks with self-attention mechanisms [52].
-The design and implementation of the first Transformers models constitute the foundation of most LLM architectures.
-Trained on massive text data, LLMs include strong capabilities in natural language understanding and complex task solving [53].
-However, LLMs face two primary challenges.
-First, they tend to produce a significant amount of hallucinations due to their reliance on internal knowledge (weights) [54].
-Second, their large parameter sizes and the associated high update costs, make traditional pre-training and fine-tuning methods unfeasible.
-As a solution, Retrieval-Augmented Generation (RAG) methods emerge as a solution for LLMs to effectively interact with the external world [55].
-This strategy is particularly notable in question-answering systems, which represent the most widely adopted use cases for harnessing the explanatory and reasoning capabilities of LLMs.
-In this domain, RAG-based approaches are emerging as the standard for developing solutions that seamlessly integrate the advantages of a knowledge base (utilizing a vector store) and generative models.
-LLMs display emergent abilities, such as In-Context Learning (ICL), where LLMs make predictions based on contexts augmented with a few examples, helping the model to learn from analogy [56].
-This paradigm simplifies the inclusion of human knowledge into LLMs by modifying the prompt templates [57].
-ICL does not require altering the model’s parameters or training the model on a specific dataset; instead, it relies on guiding the LLM through the prompting interface to align its results with human values and preferences.
-Nevertheless, evaluating LLMs and the impact of previously described approaches poses several challenges due to their complexity and the nature of their outputs.
-These difficulties can be attributed to several factors, including the lack of suitable evaluation metrics, biases in generated text, and the cost of human assessment [58].
-Nevertheless, recently, employing LLMs as evaluators for LLMs-based solutions has been proven to be nearly as effective as human judges [59].
-Previous works have demonstrated the potential of using LLMs to generate natural language descriptions of logging systems [60] or robots’ behaviors [61], indicating room for improvements to deliver more accurate and complete interpretations.
-Nevertheless, to our knowledge, no previous research has investigated the capabilities of LLMs in generating natural language explanations within the context of an accountability approach for record interpretation.
-2.5 Artificial Intelligence and Blockchain
-XAI and blockchain technology serve as complementary tools to enhance transparency, accountability, trust and acceptance of AI systems.
-In recent years, there has been a growing recognition of the importance of understanding and interpreting the decisions made by AI systems, particularly within critical domains such as healthcare, finance, and autonomous vehicles [62, 63, 64].
-XAI aims to address this need by providing human-understandable explanations for AI predictions and decisions, enabling users to trust and effectively utilize AI systems.
-However, achieving transparency and interpretability in AI models poses significant challenges, especially when dealing with complex algorithms and large datasets.
-Through blockchain, a decentralized and immutable ledger can record and verify the decision-making processes of AI models in a transparent and auditable manner.
-By integrating blockchain, XAI initiatives can store explanations for AI predictions on a distributed ledger, allowing stakeholders to validate and verify the reasoning behind AI decisions.
-A method for auditing legal decisions using blockchain technology alongside AI tools is proposed in [21].
-Using IPFS and blockchain technology, this approach offers a secure and tamper-proof record to store details, judgements, and reasoning.
-The use of Generative AI over this data helps to build trust in the system, by analyzing and finding relevant information.
-Data privacy concerns are also addressed, helping to build a more transparent, accountable, and efficient legal system.
-A framework for auditing XAI decisions using blockchain technology is proposed in [28].
-This framework includes features for both data storage and verification, with data stored in IPFS and the Ethereum blockchain preserving a unique integrity proof (hash) of the data.
-This setup allows anyone to verify the integrity of the data, thereby providing transparency and accountability in XAI decision-making processes.
-However, these approaches do not directly benefit the autonomous agents’ domain.
-Instead of relying on IPFS for storing contextual information for explanations, the use of Rosbag2 files, a binary file format, ensures the reproducibility of robots’ behaviors [42].
-Rosbags, capable of recording thousands of messages per second, offer a comprehensive and standardized format for capturing robotic data, guaranteeing the recording of each interaction and facilitating accurate analysis.
-Furthermore, given the inherent indeterminism of LLM algorithms, ensuring the integrity of the messages contained in the Rosbag file through a permissionless blockchain solution offers public transparency and accountability.
-By storing a chain of integrity proofs of these messages on the blockchain, their immutability is guaranteed, and their verification is simplified in every analysis of the Rosbag file, thereby improving the reliability of the data utilized in the explanation generation process.
+XARs are designed to provide explanations for their actions and decisions to non-expert users.
+Such explanations enhance security and trust, helping to prevent failures, errors, conflicts, and undesired interactions while achieving objectives [59, 60].
+However, despite the growing demand for explainability in autonomous agents, their use reveals a lack of practical usability and effective interpretability in real environments [61].
+Effective explanations in XARs should detail alternatives, the choices made, expectations for each option, decision criteria, and any unexpected events that may have led to changes in the planned behavior [17, 62].
+These explanations often rely on abductive reasoning, which bridges the gap between the robot’s internal decision-making process and the user’s understanding [63].
+To support this, frameworks have been developed that enable robots to articulate their internal states and decisions, enhancing collaboration and improving communication with humans [64].
+Adaptive algorithms and needs-based motivational systems allow robots to provide real-time, context-aware explanations tailored to diverse audiences [65, 66].
+Behavior Trees (BTs) hierarchically decompose tasks into goals and subgoals, making complex behaviors easier to understand and adaptable to real-time changes [67].
+However, achieving a balance between simplicity and adequacy is essential when making complex information easy to understand.
+By focusing on the most relevant details and filtering non-essential information, explanations foster trust and collaboration by aligning with human cognitive capacities [68].
+For this purpose, our architecture incorporates a log curation module that processes raw robotic data, filtering it to retain only meaningful, high-relevance information, thus delivering clearer and more concise explanations directly tied to events.
+In addition to context-aware explanations, the core of explainability in XAR lies in effective communication [63].
+Progress in Natural Language Processing (NLP) and LLMs can improve the efficiency of human-robot interactions.
+This enhancement can create a sense that these systems truly comprehend and respond to individuals meaningfully, significantly improving their usefulness and applicability [1].
+2.5 Large Language Models in Robotics
+LLMs, enabled by self-attention mechanisms, have advanced NLP, improving task-solving capabilities and enabling more effective human-robot interaction [69, 70, 71].
+However, LLMs face challenges such as hallucinations, which arise due to reliance on internal knowledge, and the high update costs associated with large parameter sizes, making traditional pre-training and fine-tuning methods impractical.
+Retrieval-Augmented Generation (RAG) methods address these issues by enabling LLMs to interact effectively with external sources [72, 73].
+This approach is particularly advantageous in robotics, where real-time and accurate contextual understanding is essential for complex decision-making and task execution.
+Explainability and transparency have become critical aspects of deploying LLMs in autonomous agents.
+These models enable robots to translate high-level commands into actionable plans, improving performance in complex environments [74, 75, 76].
+When integrated with reward decomposition frameworks, LLMs can generate intuitive explanations, linking actions to object-specific properties, thereby reducing ambiguity [77].
+LLMs can also generate natural language descriptions of robot behaviors and logs, highlighting opportunities for improvement in accuracy and completeness [78, 79].
+Existing solutions, however, often rely on static models or predefined data, limiting their ability to handle dynamic environments.
+Our approach addresses these limitations by integrating blockchain-supported accountability with curated logging of relevant robotic data.
+This framework provides robust context for LLM-generated explanations, minimizing hallucinations and ensuring alignment with the robot’s actions and interactions within its environment.
+Evaluating LLMs in robotics presents additional challenges, including the absence of suitable metrics, biases in generated text, and the high cost of human evaluations [80].
+While human evaluations remain the gold standard for capturing subjective insights [48], they are resource-intensive and lack scalability [81].
+LLM-based evaluations offer a scalable alternative, achieving over 80% alignment with human judgments while delivering explainable outputs [82].
+A hybrid framework combining human assessments for nuanced insights and LLM evaluations for efficiency and repeatability provides a comprehensive and scalable solution for evaluating LLM outputs.
 3 System Architecture
@@
 Calling an SC function intended for blockchain writing involves the execution of a transaction that changes the state of the storage data on the blockchain.
-In our approach, this transaction is signed with the account’s private key via the Elliptic Curve Digital Signature Algorithm (ECDSA), effectively identifying the signer.
+In our approach, this transaction is signed with the account’s private key via the Elliptic Curve Digital Signature Algorithm, effectively identifying the signer.
 The defined SC also determines its ownership, implementing access control to ensure that only the contract owner can add or update message proofs.
@@
 Effective communication is essential for achieving explainability in autonomous robots.
-Our approach is focused on improving this feature through the use of Natural Language Processing (NLP) technology, enabling non-expert users to interact with agents in a meaningful way.
+Our approach is focused on improving this feature through the use of NLP technology, enabling non-expert users to interact with agents in a meaningful way.
 Our solution involves the use of RAG to retrieve data from outside the language model (non-parametric) and augment the prompts by adding the information recorded from our accountability component in context.
@@
 4.1 Black Box Recorder
-ROS is the most commonly used framework in robotics, consisting of a set of software libraries and tools for building robot applications.
-However, challenges in meeting the requirements of real-time embedded systems, and the need to improve its multi-platform capabilities led to a major upgrade to ROS 2, whose main change lies in using the Data Distribution Service (DDS) [65].
-Nodes constitute the fundamental building block in both ROS and ROS 2 applications.
-They consist of individual processes, each performing specific tasks and communicating with each other via messages.
-Data can be moved between nodes using different methods, including topics, services, and actions.
-Topics are used for message passing, while services provide synchronous request-reply interactions.
-Actions, on the other hand, are needed for long-running tasks that require feedback.
-This work focuses on the navigation functionality.
-To this end, we have used Nav2, a collection of tools for ROS 2, which is considered the successor of the ROS Navigation Stack.
-Nav2 provides the robot with the ability to perform complex navigation tasks.
-Its main functionalities include planning, control, localization, visualization and other additional features.
-Furthermore, it uses behavior trees to create customized navigation behavior via orchestrating many independent modular servers, which implies high flexibility in the navigation task and the specification of complex robot behaviors [66].
-Different independent task servers, intended to compute a path, control effort, or recovery, are communicated with the behavior tree over a ROS 2 interface, such as an action server or service.
-The use of behavior trees to generate explanations for robot behavior has been previously successfully evaluated, allowing seeking causal information to answer questions [67].
-ROS 2 nodes publish data that other nodes consume through ROS 2 topics.
-Such data include perception information from sensors or commands to actuators from decision-making nodes.
+ROS is the most widely used framework in robotics, offering a collection of software libraries and tools for developing robot applications.
+Nevertheless, limitations in fulfilling the requirements of real-time embedded systems and the necessity to enhance multi-platform support led to a major upgrade to ROS 2, whose main change lies in the adoption of the Data Distribution Service (DDS) [83].
+Nodes constitute the fundamental components in both ROS and ROS 2 applications.
+These nodes are individual processes that handle specific tasks and communicate with each other through messages.
+Data transfer between nodes can be achieved using various methods, such as topics, services, and actions.
+Topics facilitate message passing, services enable synchronous request-reply interactions, and actions are utilized for long-running tasks that need feedback.
+This work focuses on navigation capabilities.
+For this purpose, we utilized Nav2, a suite of tools for ROS 2 that is regarded as the successor to the ROS Navigation Stack.
+Nav2 equips the robot with the capacity to execute complex navigation tasks.
+Nav2’s core functionalities include planning, control, localization, visualization, and additional features.
+Furthermore, it employs BTs to create customized navigation behavior by orchestrating multiple independent modular servers, offering high flexibility in navigation tasks and the specification of complex robot behaviors [84].
+Different independent task servers, intended to compute a path, control effort, or recovery, are communicated with the BT over a ROS 2 interface, such as an action server or service.
+The use of BTs to generate explanations for robot behavior has been previously successfully evaluated, allowing seeking causal information to answer questions [85].
+Within ROS 2, nodes communicate by publishing data on topics, which can be consumed by other nodes.
+The data itself can range from sensor readings that provide the robot’s perception of its environment to control commands sent to actuators from decision-making nodes.
 Navigation reproducibility of autonomous robots in ROS 2 involves storing essential data, such as their poses, maps, odometry, velocity, and laser scan readings.
-To this aim, we have used Rosbag files, a binary file format designed to capture and store data generated during the execution of a ROS system.
+To this aim, we have used Rosbag files, designed to capture and store data generated during the execution of a ROS system.
 These files can reproduce the exact conditions and inputs from a previous run, making them particularly valuable for understanding system behavior.
@@
 Ethereum’s open-source nature ensures decentralization and accessibility, allowing participation without needing approval, a feature crucial for maintaining the integrity of recorded data.
-Additionally, Ethereum’s use of Proof of Stake (PoS) as a consensus protocol enhances energy efficiency and throughput compared to traditional protocols such as Proof of Work (PoW).
+Additionally, Ethereum’s use of PoS as a consensus protocol enhances energy efficiency and throughput compared to traditional protocols such as Proof of Work.
 In addition, Ethereum’s scalability, security and economic efficiency make it a very secure option for ensuring data immutability.
@@
 By analyzing messages from the recorded Rosbag topics, this service can identify key circumstances in the navigation process, such as the appearance of an obstacle or a change in the planned trajectory.
-The status of the navigation process is checked through the study of the ROS 2 action server NavigateToPose, used for commanding a robot to move to a specified pose.
-By default, Nav2 guarantees to find the shortest path under all conditions.
-The /plan topic includes the set of poses computed to reach a specific goal pose.
-If this initial path changes due to an external circumstance, the Euclidean distance between all the poses included in the pre-computed path will increase.
-This fact, along with the values obtained through the /scan topic, allows us to determine if the appearance of an obstacle has caused the route modification.
-These reasonings have been successfully checked in previous authors’ work [68] with the aim of building explanations for autonomous agents based on the use of ROS 2 topics.
+The ROS 2 NavigateToPose action server commands the robot to navigate to a specific goal.
+While Nav2 usually prioritizes the shortest path, external factors may require adjustments.
+The /plan topic provides the sequence of poses planned to reach the goal.
+If unexpected circumstances cause a deviation from this path, the Euclidean distance between consecutive poses in the precomputed plan will increase.
+This change, together with sensor data from the /scan topic, can indicate potential obstacles causing re-planning.
+These reasonings have been successfully checked in previous authors’ work [86] to build explanations for autonomous agents based on the use of ROS 2 topics.
 Additionally, the present approach uses the information provided by Nav2 behavior trees due to their potential in designing, visualizing, and understanding the behavior of autonomous robots.
@@
 5:     $logMsg\leftarrow message.data$
-6:     // Check a change in the planned path.
-7:     if $msgTopic$ == ’/plan’  then
-8:         $distance\leftarrow$ getEuclideanDistance($message.poses$)
-9:         if $distance>previousDistance$ DISTANCE_THRESHOLD then
-10:              $logMsg\leftarrow$ “Planned path has changed when trying to achieve goal pose number $nGoal$"
-11:              $previousDistance\leftarrow distance$
-12:              if min($last\_scan\_msg.ranges$) $<$ OBSTACLE_DISTANCE_THRESHOLD then
-13:                  $logMsg\mathrel{+}=$“ because there was an obstacle."
-14:              end if
-15:         end if
+6:     if $msgTopic$ == ’/plan’  then
+7:         $logMsg\leftarrow$ Log significant path changes and obstacles.
+8:     else if $msgTopic$ == ’/navigate_to_pose/_action/status’  then
+9:         $logMsg\leftarrow$ Track navigation status and log outcomes.
+10:     else if $msgTopic$ == ’/behavior_tree_log’  then
+11:         $logMsg\leftarrow$ Log behavior tree node status.
+12:     else if $msgTopic$ == ’/amcl_pose’  then
+13:         $logMsg\leftarrow$ Record robot position and orientation.
+14:     else if $msgTopic$ == ’/cmd_vel’  then
+15:         $logMsg\leftarrow$ Log linear and angular velocities.
 16:     end if
-17:     // Check the navigation status.
-18:     if $msgTopic$ == ’/navigate_to_pose/_action/status’  then
-19:         $logMsg\leftarrow$ “No navigation is running."
-20:         $statusList\leftarrow$ /navigate_to_pose/_action/status
-21:         for $status$ in $statusList$ do
-22:              $currentGoal\leftarrow statusList.id$
-23:              if $currentGoal$ != $previousGoal$ and $currentGoal$ not in $finishedGoals$ then
-24:                  $logMsg\leftarrow$ “Navigation to a new goal has started."
-25:              else
-26:                  $logMsg\leftarrow$ “Navigation to the goal "
-27:                  if $status$ = EXECUTING then
-28:                       $logMsg\mathrel{+}=$ “is in progress.”
-29:                  else if $status$ = SUCCEEDED then
-30:                       $logMsg\mathrel{+}=$ “has succeeded."
-31:                  else if $status$ = CANCELED then
-32:                       $logMsg\mathrel{+}=$ “was cancelled."
-33:                  else if $status$ = ABORTED then
-34:                       $logMsg\mathrel{+}=$ “has been aborted."
-35:                  end if
-36:                  if $status$ = SUCCEEDED or $status$ = CANCELED or $status$ = ABORTED  then
-37:                       $finishedGoals\cup=currentGoal$
-38:                  end if
-39:              end if
-40:              // Check the executed Behavior Tree nodes and their status.
-41:              if $msgTopic$ == ’/behavior_tree_log’  then
-42:                  $logMsg\leftarrow$“Node " $message.NodeName$ “performing "$btNodeDesc$“is in status " $message.nodeStatus$
-43:              end if
-44:              // Gets robot’s position and orientation.
-45:              if $msgTopic$ == ’/amcl_pose’  then
-46:                  $logMsg\leftarrow$ “Position: " $message.x,message.y$ “Orientation: " $message.z,message.w$
-47:              end if
-48:              if $msgTopic$ == ’/cmd_vel’  then
-49:                  $logMsg\leftarrow$ “Linear velocity: " $message.x,message.y$ “Angular velocity: " $message.z$
-50:              end if
-51:         end for
-52:     end if
-53:     $logMsg\leftarrow message.timestamp$ + $logMsg$
-54:     $curationLog\leftarrow curationLog$ + $logMsg$
-55:end for
-56:return $curationLog$
+17:     Append time-stamped $logMsg$ to $curationLog$
+18:end for
+19:return $curationLog$
 {mdframed} [hidealllines=true,backgroundcolor=gray!10,innerleftmargin=3pt,innerrightmargin=3pt,leftmargin=-3pt,rightmargin=-3pt] Rosbag Raw Messages action_msgs.msg.GoalStatusArray(status_list=action_msgs.msg.GoalStatus(goal_info= action_msgs.msg.GoalInfo( goal_id=unique_identifier_msgs.msg.UUID(uuid=array([190, 81, 48, 208, 248, 147, 78, 82, 188, 116, 202, 152, 203, 225, 149, 158], dtype=uint8)), stamp=builtin_interfaces.msg.Time(sec=27, nanosec=933000000)), status=2)]) nav2_msgs.msg.BehaviorTreeStatusChange(timestamp=builtin_interfaces.msg.Time(sec=1701354686, nanosec=130104968), node_name=’ComputePathToPose’, previous_status=’IDLE’, current_status=’RUNNING’){mdframed} [hidealllines=true,backgroundcolor=gray!10,innerleftmargin=3pt,innerrightmargin=3pt,leftmargin=-3pt,rightmargin=-3pt] Processed Data 1701354686 Navigation to the goal number 1 has started.
@@
 4.4 Software and Hardware Artifacts
-Our solution has been deployed in a Linux server running Ubuntu 22.04 as the Operating System (OS). The server is equipped with a 6-core 2.6 GHz CPU, 64 GB of memory, and a 4 GB GPU.
+Our solution has been deployed in a Linux server running Ubuntu 22.04 as the operating system. The server is equipped with a 6-core 2.6 GHz CPU, 64 GB of memory, and a 4 GB GPU.
 The chosen ROS 2 version is Humble.
@@
 The experimentation was executed entirely in Gazebo, an open-source solution for 3D robotics simulations.
-To provide a realistic scenario, we used a hospital simulation environment made available by Amazon Web Services (AWS) [^2] https://github.com/aws-robotics/aws-robomaker-hospital-world.
+To provide a realistic scenario, we used a hospital simulation environment made available by Amazon Web Services [^2] https://github.com/aws-robotics/aws-robomaker-hospital-world.
 Additionally, we selected RB-1, an autonomous mobile robot based on the ROS 2 platform produced by Robotnik [87].
@@
 As a text embedding model we have chosen Instructor-xl [91], a high-performance model designed to generate text embeddings tailored to tasks such as classification, clustering, text evaluation or retrieval, a fact particularly important in the processes performed for any RAG solution.
-Table 2: Robag topics rate and average rate per message
+Table 2: Rosbag topics rate and average rate per message.
 Topic Avg. Rate (Hz) Avg. Size Msg. (B)
@@
 This work aimed to assess the effectiveness of these technologies in delivering natural language explanations to individuals who are not experts.
-The experiments conducted within this proposal focus on navigation tasks included in the Navigation Functionality of the European Robotics League Consumer Service Robots Challenge (ERL Consumer) [74].
+The experiments conducted within this proposal focus on navigation tasks included in the Navigation Functionality of the European Robotics League Consumer Service Robots Challenge [92].
 This functionality evaluates the capabilities of a robot to navigate accurately and autonomously within a given scenario, containing furniture, objects, and people.
@@
 This approach, together with those operating at fixed intervals, has been compared with the generation of a Rosbag file that covers identical topics but lacks anti-tampering properties.
-Each experiment has been repeated six times, totalling 108 runs, to derive insights into the characteristics of the black box-like device.
+Each experiment has been repeated six times, totaling 108 runs, to derive insights into the characteristics of the black box-like device.
 5.1.1 Metrics
@@
 These variables are highly significant for system-overhead analysis given that elevated load conditions can potentially influence the robot’s behavior and event-recording capabilities.
-In particular, we have obtained a full report about system load, CPU, RAM, network traffic and disk write operations updated every second, facilitating the tracking of overall system activity and an easy comparison with preceding values.
+In particular, we have obtained a full report about system load, CPU, RAM, network traffic, and disk write operations updated every second, facilitating the tracking of overall system activity and an easy comparison with preceding values.
 To describe the overall distribution and characteristics of these results, we calculated their average as a statistical measure.
@@
 5.2 Explainability Evaluation
-Despite the remarkable performance of LLMs across diverse language tasks, evaluating these models presents challenges, especially in ensuring factual consistency.
-The extensive capabilities of LLMs and the limitations of current benchmarks in capturing human preferences have given rise to emerging trends in LLM evaluations, such as employing strong LLMs as judges [59].
+Despite the remarkable performance of LLMs in various language tasks, evaluating these models presents challenges, especially in ensuring factual consistency.
+The extensive capabilities of LLMs and the limitations of current benchmarks in capturing human preferences have led to emerging trends in LLM evaluations, such as the use of strong LLMs as judges [82].
 Traditional similarity-based metrics, such as BLEU and ROUGE, have proven to misalign with human evaluation.
 While human evaluation remains the primary standard for assessing human preferences, it is exceptionally time-consuming and expensive.
-To automate the evaluation process, several works have adopted the ’LLM-as-a-judge’ approach, which focuses on using strong LLMs like GPT-4.
+Furthermore, this approach significantly impacts the scalability of evaluations, making it necessary to use more effective and unbiased methods [81, 93].
+To automate the evaluation process, several works have adopted the ‘LLM-as-a-judge’ approach, which focuses on using strong LLMs like GPT-4.
 These models are often trained with RLHF, resulting in strong human alignment [94].
-Previous studies have evidenced the consistent agreement of these models with human grading in over 80% of judgments [76, 77, 78].
-Given these characteristics, we opted for GPT-4 to assess the answers generated by our explainability component.
+Previous studies have evidenced the consistent agreement of these models with human grading [95, 96, 97, 98].
+Moreover, research shows that LLMs can replicate most of conclusions from human evaluations, particularly in tasks involving prediction accuracy and explanation helpfulness, even across varied settings and domains [81, 99, 100].
+However, while LLMs excel in scalability and efficiency, their strengths can be enhanced by integrating the nuanced understanding and subjective judgments that humans provide, particularly in complex or domain-specific contexts [101].
+For example, human evaluators can detect subtle inconsistencies or interpret ambiguous explanations in tasks that require deeper contextual understanding, which an automated method might overlook [48, 102].
+To address these limitations, we propose a hybrid evaluation method that combines the scalability of LLM-based evaluations with the depth of human assessments.
+In this approach, a strong LLM such as GPT-4 is used to evaluate a broad set of model outputs, ensuring cost-effective and consistent initial assessments.
+Human evaluators then validate a representative subset of these outputs to uncover overlooked biases or misalignments.
+This approach integrates the efficiency and scalability of automated methods with the reliability of human judgment, particularly for nuanced cases.
+By adopting this hybrid methodology, we balance efficiency and accuracy, ensuring that the evaluation process captures both objective metrics and subjective preferences effectively.
+This dual-layered evaluation enables comprehensive assessments while addressing the inherent limitations of relying exclusively on either LLMs or humans.
+5.2.1 LLM-as-a-judge Evaluation
+The need for a robust LLM aligned with human preferences led us to select GPT-4 to evaluate the answers produced by our explainability component.
 Through LLM-as-a-judge, we can obtain not only scores but also explanations, making their outputs interpretable.
@@
 Furthermore, the LLM must not only provide correct information but also deliver it in a meaningful and valuable way to the user.
-Therefore, to assess the performance of the selected LLM on specific attributes, our approach also includes a criteria-based evaluation covering accuracy, conciseness, coherence, and helpfulness.
+Therefore, to assess the performance of the selected LLM on specific attributes, this approach also includes a criteria-based evaluation covering accuracy, conciseness, coherence, and helpfulness.
 Selected criteria metrics are described in Table 3.
@@
 A complete example of the answers and evaluations obtained for a selected question in the three selected scenarios is included in Figure 6.
-Table 3: Criteria-based evaluation metrics
-Criteria metric Description
+Table 3: Criteria-based evaluation metrics in LLM-as-a-judge evaluation.
+Metric Description
 Accuracy Accuracy evaluates the extent to which the response conforms to the reference, ranging from being completely unrelated to being completely accurate and in line with the reference.
@@
 Accuracy: [[10]] Conciseness: [[7]] Coherence: [[10]] Helpfulness: [[10]]", score: 1.0
-Figure 6: Examples of correct answers, RAG, and criteria-based evaluations obtained across the three scenarios for the question ’How many goals have been reached by the robot?’
+Figure 6: Examples of correct answers, RAG, and criteria-based evaluations obtained across the three scenarios for the question ’How many goals have been reached by the robot?’ with LLM-as-a-judge evaluation.
+5.2.2 Human Evaluation
+To validate the explanation results generated by the GPT-4 model, we conducted an online study with 17 participants with diverse demographics.
+The age distribution ranged from 18 to 55 years, with the majority (47.06%) in the 26-35 age group, followed by 35.29% in the 18-25 age group, 11.76% in the 36-45 age group, and 5.88% in the 46-55 age group.
+In terms of gender, 62.5% were male and 37.5% were female.
+Participants’ highest level of education varied, with 37.5% holding a PhD, 37.5% having a Master’s Degree, and 25% possessing a Bachelor’s Degree.
+Regarding expertise in explainability for autonomous agents, 37.5% were beginners, 31.25% had intermediate expertise, 12.5% were advanced, and 18.75% had no experience.
+The evaluation process involved three distinct questionnaires, one corresponding to each scenario.
+Participants assessed 16 explanations per scenario, resulting in a total of 48 explanations evaluated by each individual.
+Each explanation was rated using five questions derived from the metrics outlined in Table 4, based on the Explanation Satisfaction Scale proposed by Hoffman et al.
+in [102].
+This scale represents a widely recognized framework for measuring human responses to XAI systems.
+To ensure the evaluation aligned with human preferences, we used a 5-point Likert scale.
+Answers in this approach are distributed in 1 (I disagree strongly), 2 (I disagree somewhat), 3 (I’m neutral about it), 4 (I agree somewhat), and 5 (I agree strongly).
+In addition to the explanation evaluations and optional demographic information, the questionnaire provided general instructions describing the robot’s task and included a map for each scenario.
+The questionnaires were distributed via a mailing list, and participants were informed about the anonymity and research purposes of the study.
+Consent was obtained upon submission of their answers.
+Participation was voluntary, which made the process cost-effective but also introduced potential drawbacks, such as longer response times and biases that may arise from factors like age, education level, or familiarity with autonomous system explainability approaches.
+Statistical analysis is essential in human evaluations to ensure rigor and reliability, particularly when interpreting subjective judgments and addressing variability among participants.
+Likert-scale data, which is ordinal, requires careful statistical treatment to identify significant trends and differences while avoiding misinterpretation of the results.
+By applying appropriate statistical methods, we strengthen the credibility of our conclusions and provide a robust evaluation of the LLM’s results across scenarios.
+To ensure the reliability and robustness of the evaluation, a series of statistical tests were applied.
+Reliability analysis of the scale criteria was conducted using Cronbach’s Alpha, a robust metric for assessing internal consistency and ensuring that the scale reliably measures the intended constructs, with a commonly accepted threshold for reliability being $\alpha>0.7$ [103, 104].
+To further explore the relationships among variables in the dataset, correlation analysis was performed using Pearson correlation matrices, which provide insights into the strength and direction of linear relationships between variables [105].
+For inferential analysis, a normality check was conducted using the Shapiro-Wilk test to determine whether the data met the assumptions required for parametric tests [106].
+Since the data violated these assumptions, the non-parametric Friedman test was employed [107].
+The Friedman test is particularly suited for analyzing repeated measures on the same participants, as it does not require data to follow a normal distribution [108].
+This test assessed whether significant differences existed in participant responses across the three scenarios, offering a reliable framework for understanding the impact of different conditions on the perceived quality and effectiveness of explanations.
+Finally, post-hoc analysis using the Nemenyi test was performed to examine participant behavior in specific explanations where significant differences were observed, offering a deeper understanding of the underlying patterns in their responses.
+The statistical analyses were developed through Python scripts (version 3.10.12), utilizing libraries such as pingouin (version 0.5.5) for Cronbach’s Alpha, scipy.stats (version 1.14.1) for the Shapiro-Wilk and Friedman tests, and pandas (version 2.3.3) for data manipulation.
+This approach ensured accuracy and reproducibility in the execution of the tests.
+Table 4: Questions and metrics used for human evaluation.
+Metric Question
+Understandability From the explanation, I understand how the task proceeded.
+Satisfaction The explanation is satisfactory.
+Informativeness The explanation has sufficient detail.
+Completeness The explanation seems complete.
+Usefulness The explanation is useful to my goals
 5.2.3 Categories and Evaluation Questions
-To evaluate our explainability proposal in the context of ROS 2 messages interpretation, we have defined a set of questions covering diverse aspects of the navigation process over the how, the what, the which, the when and the where dimensions.
+To evaluate our explainability proposal in the context of ROS 2 messages interpretation, we defined a set of questions covering diverse aspects of the navigation process over the how, the what, the which, the when and the where dimensions.
 These questions provide insights into the different stages and components of the three scenarios.
-To structure and improve the understanding and analysis of this set of questions, we have organized them into five different categories.
-These categories range from high-level overviews of navigation to detailed inquiries about trajectory planning, goal completion, and the functioning of the behavior tree.
+To structure and improve the understanding and analysis of this set of questions, we organized them into five different categories.
+These categories range from high-level overviews of navigation to detailed inquiries about trajectory planning, goal completion, and the functioning of the BT.
 The categories also delve into the causes of unexpected behavior or errors in navigation, improving the explainability of the system whenever an unexpected action occurs.
@@
 These findings may encourage the adoption of our approach to ensure integrity in messages from autonomous agents.
-6.3 Correctness Evaluation
+6.3 Correctness Evaluation Using LLM-as-a-judge
 Figure 10 shows the results of the correctness of RAG evaluation, presenting both the overall results and assessments based on predefined categories.
-The results demonstrate remarkable consistency across the three scenarios, getting a significant number of correct answers, even though there has not been a prior tuning in the ROS 2 messages for autonomous agents’ navigation tasks.
+The results demonstrate consistency across the three scenarios, getting a significant number of correct answers, even though there has not been a prior tuning in the ROS 2 messages for autonomous agents’ navigation tasks.
 The slight percentage of incorrect answers remains within an acceptable threshold, given the specific context employed and the inclusion of general questions.
 Figure 10: (a) Overall correctness in the three scenarios. (b) Correctness by category in the three scenarios.
-Regarding individual categories, the correctness in the category ’Navigation Process Overview’ is relatively consistent across all scenarios, indicating a good understanding of events and progress related to navigation.
+Regarding individual categories, the correctness in the category ‘Navigation Process Overview’ is relatively consistent across all scenarios, indicating a good understanding of events and progress related to navigation.
 However, the broad nature of the question "What has happened in this ROS 2 log regarding navigation?" initially yielded a correctness below 65% in the third scenario.
 In this instance, the cancellation of a goal resulted in the execution and failures of Nav2 Behavior Tree nodes, not included in the other two scenarios and focused on navigation recovery.
-Therefore, the incorporation in a dataset of information related to the cancellation of a goal results in LLM answers that usually diverge from the ground truth provided, as the model occasionally prioritized Behavior Trees information, other times focused on locations or timestamps.
+Therefore, the incorporation in a dataset of information related to the cancellation of a goal results in LLM answers that usually diverge from the ground truth provided, as the model occasionally prioritized BTs information, other times focused on locations or timestamps.
 This variability made it difficult to align with the ground truth provided, even if the obtained answers could be considered partially accurate.
@@
 Nevertheless, this challenge is not present in Scenario 3 and is absent or reduced in Scenario 2, demonstrating how differences in data distribution can introduce varied responses even when using specific and unambiguous questions.
-Through questions included in the Nav2 Behavior Tree and node status category, we assessed the effectiveness of the model’s formal linguistic competence compared to its functional linguistic competence, a feature of present LLMs [80].
-While the model excels in identifying Behavior Tree Nodes given its description, it faces challenges in pinpointing these components under specific conditions.
-As a result, the category correctness suffers a drop in Scenario 3, although overall results remain above 75%.
-In this scenario, the navigation recovery actions caused by the cancellation of the first goal introduced a challenge in the interpretation of the information by the LLM due to the execution of several Nav2 Behavior Tree nodes not present in the other scenarios.
-Particularly, the mistakes stem from the model’s confusion in determining whether a previously failed Nav2 Behavior Tree node has recovered or finished its execution, as well as identifying all nodes that were executed during navigation.
-This can be attributed to the fact that the complexities of the recovery actions and the dynamic nature of Behavior Tree executions during navigation recovery introduce complexities that the model may not fully generalize from the training data, underscoring the need for additional fine-tuning in such scenario-specific contexts in case the obtained correctness needs to be improved.
-6.4 Criteria-based Evaluation
-Figure 11 shows the remarkable performance of the selected LLM in the evaluated metrics, demonstrating consistency with slight variations in the scores.
+Through questions included in the ‘Nav2 Behavior Tree and Node Status’ category, we assessed the effectiveness of the model’s formal linguistic competence compared to its functional linguistic competence, a feature of present LLMs [110].
+While the model excels in identifying BT nodes given its description, it faces challenges in pinpointing these components under specific conditions.
+As a result, the criteria correctness suffers a drop in Scenario 3, although overall results remain above 75%.
+In this scenario, the navigation recovery actions caused by the cancellation of the first goal introduced a challenge in the interpretation of the information by the LLM due to the execution of several Nav2 BT nodes not present in the other scenarios.
+Particularly, the mistakes stem from the model’s confusion in determining whether a previously failed Nav2 BT node has recovered or finished its execution, as well as identifying all nodes that were executed during navigation.
+This can be attributed to the fact that the complexities of the recovery actions and the dynamic nature of BT executions during navigation recovery introduce complexities that the model may not fully generalize from the training data, underscoring the need for additional fine-tuning in such scenario-specific contexts in case the obtained correctness needs to be improved.
+6.4 Criteria-based Evaluation Using LLM-as-a-judge
+Figure 11 shows the performance of the selected LLM in the evaluated metrics, demonstrating consistency with slight variations in the scores.
 On average, the answers provided are accurate and in line with the reference, though there might be room for improvement.
 While responses generally provide useful information, there are instances where additional depth could enhance their utility.
-In addition, most answers usually include a remarkable balance between conciseness and effectiveness, showing a proficient level of brevity, as well as clear and logical organization.
+In addition, most answers usually include a good balance between conciseness and effectiveness, with clear and logical organization.
 The overall results show no notable differences in these metrics across the three scenarios.
@@
 Concerning general questions about events and the progression of the navigation task, we observe variations in scores across Scenarios 2 and 3.
-While coherence and conciseness are even improved, Scenario 1, involving direct navigation to three goals, outperforms in accuracy and helpfulness.
+While coherence and conciseness improve, Scenario 1, which involves direct navigation to three goals, scores higher in accuracy and helpfulness.
 This fact suggests that effective communication on general queries benefits from the absence of obstacles or goal cancellations, enhancing overall understanding.
-In the category ’Trajectory Planning and Re-planning’, the questions are effectively addressed in all three scenarios, showing scores consistently above 8 in every evaluated metric, even when obstacles are introduced or a goal is cancelled.
-These results underline the system’s proficiency in handling trajectory-related questions despite potential new challenges.
-Regarding the category ’Goal Completion and Navigation Task Status’, Scenarios 1 and 2 demonstrate the effectiveness of the system in conveying goal completion and navigation task status information.
-Nevertheless, the slightly lower scores for Scenario 3 indicate potential challenges in effectively communicating completion status when a goal is aborted.
-Concerning the ’Specifics about Goals and Locations’ category, the system consistently provides valuable goal and location information, even in the presence of obstacles or goal cancellations.
-This suggests a robust capacity to understand and localize specific details regardless of the challenges introduced in each scenario.
-The results regarding questions included in ’Nav2 Behavior Tree and Node Status’ demonstrate effectiveness in communication across all scenarios and metrics.
-This emphasizes the potential for Behavior Trees to complement language models in explainability, offering insight into the decision-making processes during navigation tasks.
-In summary, although results can be improved in accuracy regarding general questions, the language model exhibits adaptability to different navigation scenarios, facing obstacles and changes in navigation tasks.
-The consistency in conveying information related to Nav2 Behavior Trees suggests a robust understanding of underlying navigation processes.
+In the category ‘Trajectory Planning and Re-planning’, questions are effectively addressed across all three scenarios, showing scores above 8 in every metric, even when obstacles or goal cancellations are introduced.
+These results highlight the system’s capability to manage trajectory-related queries effectively despite challenges.
+Regarding the category ‘Goal Completion and Navigation Task Status’, Scenarios 1 and 2 effectively convey information about goal completion and navigation task progress.
+Nevertheless, slightly lower scores in Scenario 3 indicate challenges in effectively communicating completion status when a goal is aborted.
+In the ‘Specifics about Goals and Locations’ category, the system provides valuable goal and location information, even in the presence of obstacles or goal cancellations.
+This suggests a strong ability to understand and localize specific details regardless of the challenges introduced in each scenario.
+The results regarding questions included in ‘Nav2 Behavior Tree and Node Status’ category demonstrate effectiveness in communication across all scenarios and metrics.
+This emphasizes the potential for BTs to complement language models in explainability, offering insight into the decision-making processes during navigation tasks.
+In summary, although results can be improved in accuracy for general questions, the language model shows adaptability across different navigation scenarios, including those with obstacles and task changes.
+The consistency in conveying information related to Nav2 BTs suggests a solid understanding of underlying navigation processes.
 The model also performs well when providing information about trajectory planning, goal completion, and specifics about goals and locations, regardless of the specific challenges introduced in each scenario.
-Its robustness in effective communication, despite technical vocabulary, details and the introduced challenges, is demonstrated through the achievement of responses that are not only generally accurate but also easily comprehensible to users without expertise, which enhances the model’s practical utility.
+Nevertheless, while evaluation using LLMs excels in scalability and efficient resource usage, it is still recommended to test a subset of questions with human evaluators.
+This ensures a deeper understanding of how well the approach aligns with human expectations and effectively communicates, particularly in scenarios involving complex or nuanced details.
+6.5 Human Evaluation of LLM-generated Explanations
+This section presents an analysis of the results from the human evaluation conducted with 17 participants, as stated in Section 5.2.2.
+6.5.1 Distribution Analysis
+Figure 13 depicts the distribution of Likert scale responses for the five evaluation criteria across the three scenarios, showing the frequency of each response value.
+The highest ratings predominate across all scenarios and criteria, suggesting that participants generally found the explanations satisfactory.
+However, Scenario 2 and Scenario 3 exhibit greater variability, with an increase in mid-range scores for certain criteria, such as Usefulness.
+These trends are particularly noticeable in broad questions, such as those in Category 1 regarding the navigation process overview, or when the model’s answers do not align with the ground truth.
+In such cases, participant opinions diverge significantly, even in the presence of factual inaccuracies, indicating a potential bias in how the utility is perceived when discrepancies with the ground truth arise.
+Figure 13: Distribution of Likert scale responses across scenarios and evaluation criteria.
+6.5.2 Reliability Analysis
+The reliability analysis resulted in an overall Cronbach’s Alpha value of 0.9792, indicating good internal consistency and confirming that the evaluation criteria reliably capture the intended dimensions of the robot’s explanations.
+Additionally, we calculated Cronbach’s Alpha if an item is deleted for each criterion.
+This metric helps assess the impact of removing an individual item on the scale’s overall reliability.
+As summarized in Table 5, the "Alpha if Deleted" values are consistently close to the overall Cronbach’s Alpha.
+Specifically, these values range from 0.9699 (Usefulness) to 0.9798 (Informativeness), with all values slightly lower than the overall result.
+This suggests that removing any single item would only have a minimal impact on the reliability of the scale, indicating that each criterion contributes positively to the measurement of overall quality.
+Further, we examined the item-total correlations, which quantify the strength of the relationship between each criterion and the total score.
+These correlations range from 0.9286 for the Informativeness criteria to 0.9660 for Usefulness.
+The high item-total correlations support the relevance of each item in assessing the key dimensions of the robot’s explanations.
+Table 5: Cronbach’s Alpha and item-wise "Alpha if Deleted" in human evaluation.
+Criteria Cronbach’s Alpha if Deleted Item-Total Correlation
+Completeness 0.9724 0.9492
+Informativeness 0.9798 0.9286
+Satisfaction 0.9752 0.9567
+Understandability 0.9724 0.9502
+Usefulness 0.9699 0.9660
+Finally, we assessed the inter-item correlations, which provide insight into how strongly the criteria correlate with one another.
+The inter-item correlation matrix, presented in Table 6, shows values ranging from 0.878 (between Informativeness and Satisfaction) to 0.981 (between Satisfaction and Usefulness).
+These high correlations indicate that the criteria are closely related but not redundant, as each captures a unique and complementary aspect of the evaluation process.
+Table 6: Inter-Item correlation matrix in human evaluation.
+Completeness Informativeness Satisfaction Understandability Usefulness
+Completeness 1.000 0.976 0.905 0.911 0.911
+Informativeness 0.976 1.000 0.878 0.887 0.881
+Satisfaction 0.905 0.878 1.000 0.936 0.981
+Understandability 0.911 0.887 0.936 1.000 0.947
+Usefulness 0.911 0.881 0.981 0.947 1.000
+6.5.3 Correlation Analysis
+Building on the findings from the reliability analysis, we further explored the relationships among the variables in our dataset through Pearson correlation matrices.
+While Cronbach’s Alpha results demonstrated the internal consistency of the evaluation criteria, the correlation analysis extends this by examining how these criteria relate to each other across different scenarios and explanation categories.
+Figure 14 presents the Pearson correlation matrix for participants’ responses across the three scenarios.
+It reveals strong associations, indicating a high level of consistency in participants’ assessments.
+This consistency suggests that the explanations provided were clear and easily comparable, allowing participants to reliably evaluate them regardless of the specific context of the robot’s task.
+Figures 15 and 16 depict the relationships among evaluation criteria and question categories.
+Regarding the criteria, Completeness and Informativeness exhibit a strong positive correlation, as do Understandability and Usefulness.
+This indicates that participants are more likely to perceive explanations as useful when they are also clear.
+In contrast, the correlation between Informativeness and Usefulness is weaker (0.55), highlighting a divergence in participants’ evaluations.
+This divergence may stem from biases previously discussed, where participants find it challenging to judge the usefulness of an explanation that is not fully accurate.
+Regarding the question categories, the ‘Navigation Process Overview’ and ‘Specifics About Goals and Locations’ categories exhibit a strong correlation (0.99), indicating that participants tended to provide similar ratings for explanations in these areas.
+This suggests that these types of explanations are perceived similarly.
+In contrast, the ‘Trajectory Planning and Re-planning’ category shows weak or negative correlations with other categories.
+This discrepancy arises from differences in evaluating broad, open-ended explanations versus more specific ones.
+Figure 14: Correlation matrix of responses by scenario.
+Figure 15: Correlation matrix of responses by criteria.
+Figure 16: Correlation matrix of responses by category.
+6.5.4 Inferential Statistics
+Statistical analysis often requires understanding the underlying data distribution to ensure the appropriate methods are applied.
+Since parametric tests assume normality, violations of this assumption can compromise the validity of the results.
+To assess normality, we applied the Shapiro-Wilk test, a robust method for small to moderate-sized samples, across various groupings, including scenarios, categories, questions, and criteria.
+The test results consistently revealed significant deviations from normality, with $p$-values well below 0.05, leading to the rejection of the null hypothesis of normality for all groups.
+These findings confirm that the dataset does not satisfy the assumptions required for parametric tests, such as ANOVA or t-tests.
+As a result, non-parametric methods were adopted in subsequent analyses to ensure the robustness and validity of statistical inferences.
+The Friedman test was used due to its ability to assess differences in related groups when the data does not meet the assumptions of normality required for parametric tests.
+The results of the Friedman test across the three scenarios reveal significant differences in participant ratings for several explanations, indicating varying levels of agreement across conditions.
+In Scenario 1, significant differences were observed for explanations to Questions 1 and 2, suggesting diverse responses.
+In contrast, most other questions in this scenario showed no significant differences ($p\geq 0.05$).
+Explanations 1 and 2 in all three scenarios addressed broad questions, whose answers diverged from the ground truth, as discussed in the GPT-4 evaluation in 6.3.
+Broad questions often result in responses that prioritize varied information, such as BTs, locations, or timestamps, making it difficult to align the answers with the ground truth.
+This fact justifies the significant variability in participant ratings for these explanations.
+In Scenario 2, significant differences were observed for Explanations 1, 7, 8, and 15, with $p$-values below the 0.05 threshold, while other questions did not show significant variations.
+Specifically, the explanations for Questions 7 and 8 were misaligned with the ground truth, leading to significant variability in participant assessments.
+In these cases, participants’ ratings lacked a clear trend across the evaluated criteria.
+A similar pattern was observed in Explanation 15 of Scenario 2, where the model provided irrelevant information that did not directly address the user’s inquiry.
+This lack of relevance and alignment likely contributed to the observed variability in participant ratings.
+Finally, Scenario 3 showed significant differences for Explanations 1, 2, 7, and 16, indicating that the conditions for these explanations led to notable variations in participant ratings.
+Explanation 7, does not address the answer to the corresponding question, including no relevant information.
+Explanation 16 in Scenario 3 caused variability in participants’ answers due to its incomplete nature, highlighting the importance of content clarity in generating consistent and meaningful explanations.
+These findings suggest that while most explanations remain consistent, certain ones resonate differently depending on the scenario, nature of the question, and LLMs indeterminism.
+Figure 17: Friedman test results.
+The post-hoc Nemenyi test was conducted for explanations that showed significant differences in participant ratings based on the Friedman test results.
+This allowed us to identify specific participant pairs with notable divergences in their responses across the three scenarios.
+The Nemenyi test revealed that most participants were consistent in their evaluations, but certain pairs exhibited significant variability.
+For example, in Scenario 1, the evaluation for Explanations 1 and 2 highlighted differences involving Participant 8 and several others.
+In Scenario 2, significant differences emerged for Explanation 1 among Participants 3, 6, 7, and 13.
+Similarly, in Scenario 3, variability was observed for Explanations 1 and 16, where selected participants differed significantly from others.
+These discrepancies are likely due to natural human biases, differences in interpretation, and subjective judgments that are inherent in human responses.
+Overall, while the explanations were generally clear and consistent for most participants, the observed variability underscores the role of individual perspectives and biases in shaping evaluations, reflecting the inherent complexity of human judgment.
+6.6 Similarity Between Human and LLM Evaluations
+As described in 5.2, previous studies have demonstrated that LLMs can effectively approximate human judgment in various evaluation tasks, including those involving subjective interpretation.
+In this work, we conducted a focused validation of LLM-generated evaluations by comparing a subset of 48 explanations from one of six evaluation runs performed by the LLM against human assessments.
+This subset included 16 explanations per scenario, allowing us to validate the LLM’s judgment within a practical scope.
+Figures 18 and 19 depict the mean responses from 17 human participants and GPT-4, respectively, acting as a judge in this subset.
+The comparison highlights notable similarities in scoring patterns across scenarios and explanations, suggesting a high degree of alignment between human and LLM evaluations.
+Both evaluations consistently identify high-quality explanations with scores clustered around 4.5–5, indicating shared agreement on the quality of certain explanations.
+Similarly, lower-rated explanations identified by humans are also flagged by the LLM, revealing a mutual recognition of weaker content.
+However, while the overall trends are similar, there are some differences in the exact ratings and outlier values between the two evaluations.
+These differences could be attributed to nuances in how the LLM interprets the explanations compared to human evaluators, who might take into account subtle contextual or subjective factors that an LLM might miss.
+Despite these slight variations, the visual correspondence between the human and LLM heatmaps indicates that the LLM provides a reliable approximation of human judgment, making it a promising tool for scaling evaluation processes without sacrificing much in terms of accuracy.
+This similarity could be valuable in demonstrating that LLMs are suitable alternatives for tasks like explanation evaluation, especially in large-scale studies where human evaluation may be impractical.
+Figure 18: Heatmap of mean responses from human evaluators Across Scenarios and Explanations.
+Figure 19: Heatmap of mean responses from the LLM-as-Judge evaluation across scenarios and explanations.
 7 Conclusions
@@
 This last component is based on the use of Retrieval Augmented Generation (RAG) and Large Language Models (LLMs) through a private question-answering system designed to use the information recorded by the first component as context.
-This solution demonstrates the feasibility of obtaining accurate, coherent, concise and helpful explanations for robot actions from raw data stored in Rosbag files, thereby enhancing human-robot interaction.
+This solution demonstrates the feasibility of obtaining accurate, coherent, concise and helpful explanations for robot actions from raw data stored in Rosbag files, providing satisfactory results for users and thereby enhancing human-robot interaction.
 Additionally, both main components have been integrated through the development of a third main functionality in charge of filtering non-relevant information and making messages closer to the final user, due to their specific nature can introduce challenges in their comprehension process.
@@
 This fact strengthens trust and confidence in the use of autonomous agents across diverse applications and contexts.
-While our approach has shown remarkable results in navigation functionalities regarding several scenarios and circumstances, future work should involve adapting our solution to other applications by simply adapting the recording and the processing components to the relevant topics of the area or domain of use.
+While our approach has shown promising results in navigation functionalities regarding several scenarios and circumstances, future work should involve adapting our solution to other applications by simply adapting the recording and the processing components to the relevant topics of the area or domain of use.
 Furthermore, the consistency and robustness of responses generated by our explainability approach could benefit from refinement across different scenarios, especially when categorizing questions under specific subjects.
-Therefore, future work should explore the impact of In-Context Learning (ICL) techniques on prompt generation, in addition to fine-tuning the model for the specific context of autonomous agent navigation tasks.
+Therefore, future work should explore the impact of In-Context Learning techniques on prompt generation, in addition to fine-tuning the model for the specific context of autonomous agent navigation tasks.
 Additionally, our RAG approach could improve in terms of retrieving the most relevant chunk of information from the context.
@@
 A.2 Explanations Obtained
-The 288 questions and answers obtained through the explanation generation and evaluation process are available on the aforementioned mentioned Github repository  [^4] https://github.com/laurafbec/immutable_explainable_BBR/blob/main/
+The 288 questions and answers obtained through the explanation generation and evaluation process are available in the aforementioned mentioned Github repository  [^4] https://github.com/laurafbec/immutable_explainable_BBR/blob/main/
 gpt4_eval_results/LLMs_QA_Results.xlsx
@@
 Additionally, a Python notebook has been developed to interact with the content of the previous sheet.
-This notebook is available on  [^5] https://github.com/laurafbec/immutable_explainable_BBR/blob/main/
+This notebook at the following link  [^5] https://github.com/laurafbec/immutable_explainable_BBR/blob/main/
 gpt4_eval_results/GPT4 Evaluation Results.ipynb
@@
 Acknowledgments
-This research has been partially funded by the Recovery, Transformation, and Resilience Plan, financed by the European Union (Next Generation) thanks to the TESCAC project (Traceability and Explainability in Autonomous Systems for improved Cybersecurity) granted by INCIBE to the University of León; and by grant PID2021-126592OB-C21 funded by MCIN/AEI/10.13039/501100011033 EDMAR (Explainable Decision Making in Autonomous Robots) project, PID2021-126592OB-C21 funded by
-MCIN/AEI/10.13039/501100011033 and by ERDF ”A way of making Europe”.
+This publication is part of the TESCAC project, financed “by European Union NextGeneration-EU, the Recovery Plan, Transformation and Resilience, through INCIBE”.
+In addition, this work has been partially funded by the project EDMAR, PID2021-126592OB-C21, funded by MCIN/AEI/10.13039/501100011033 and by ERDF A way of making Europe.

### v4

@@
 The deployment of autonomous agents in environments involving human interaction has increasingly raised security concerns.
-Consequently, understanding the circumstances behind an event becomes critical, requiring the development of capabilities to justify their behaviors to non-expert users.
-Such explanations are essential in enhancing trustworthiness and safety.
-Additionally, they contribute to improving communication, bridging the gap between the agent and the user, thereby improving the effectiveness of their interactions.
-This work presents an accountability and explainability architecture implemented for mobile Robot Operating System (ROS)-based robots.
-The proposed solution consists of two main components.
-First, a black box-like element to provide accountability, featuring anti-tampering properties achieved through blockchain technology.
-Second, a component in charge of generating natural language explanations by harnessing the capabilities of Large Language Models (LLMs) over the data contained within the previously mentioned black box.
-The study evaluates the performance of our solution in three different scenarios, each involving autonomous agent navigation functionalities.
-This evaluation includes a thorough examination of accountability and explainability metrics, demonstrating the effectiveness of our approach in using accountable data from robot actions to obtain coherent, accurate and understandable explanations, even when facing challenges inherent in the use of autonomous agents in real-world scenarios.
-keywords:
+Consequently, understanding the circumstances behind an event becomes critical, requiring the development of capabilities to justify their behavior to non-expert users.
+Such explanations are essential to fostering trust and ensuring safety.
+Additionally, they contribute to improving communication by bridging the gap between the agent and the user, thereby enhancing the effectiveness of their interactions.
+This paper presents an accountability and explainability architecture implemented for mobile Robot Operating System-based robots.
+The proposed solution comprises two main components.
+First, a black box-like element used to provide accountability, featuring anti-tampering properties achieved through blockchain technology.
+Second, a component in charge of generating natural language explanations by harnessing the capabilities of Large Language Models over the data contained within the previously mentioned black box.
+This study evaluates the performance of our solution in three different scenarios, each involving autonomous agent navigation functionalities.
+This evaluation includes a thorough examination of accountability and explainability metrics, demonstrating the effectiveness of our approach in using accountable data from robot actions to obtain coherent, accurate and understandable explanations, even when facing the challenges inherent in the use of autonomous agents in real-world scenarios.
+Index Terms:
+Accountability, explainability, autonomous agents, robotics, blockchain, human–robot interaction, large language models
 […]
-{highlights}
-Security improvements for autonomous agents in public spaces.
-Review of blockchain for immutable logs, EDRs, XAI and blockchain, XARs, and LLMs.
-Accountability and Explainability architecture for autonomous agents.
-Comprehensive assessment across diverse navigation scenarios and metrics.
-Future research directions for accountable and explainable autonomous agents.
-1 Introduction
-The rising presence of mobile robots alongside humans has increased concerns about security and trust.
-One of the main challenges in this context is the ability of robots to justify their actions and decisions, particularly in safety-critical scenarios where their behavior may impact human well-being [1].
-Transparency and reliability in these justifications are essential [2], as such properties enable effective human-robot interaction and foster confidence, particularly for non-expert users who require clear and understandable explanations [3, 4].
-However, the complexity of robotic decision-making and the need for thorough post-event analysis make achieving these goals particularly difficult.
-To address these challenges, accountability must be ensured through a robust, tamper-proof logging system.
-The most substantial evidence for post-event analysis comes from logging and auditing capabilities, which are essential to mitigate technical issues and provide forensic evidence to detect data tampering or unauthorized breaches.
-These systems must guarantee the integrity, confidentiality, and availability of data while enabling incident detection, compliance, and forensic investigations [5].
-However, covering tracks in log systems has become a common practice for attackers, erasing any traces that could lead to their identification or assisting in the analysis of their actions.
-As a result, it is necessary to strengthen system properties against unauthorized modifications [6].
-This includes safeguarding the integrity of log data and implementing measures to detect and counteract attempts at tampering.
+=-21pt
+I Introduction
+The increasing integration of mobile robots into human environments has raised significant concerns regarding their transparency and trustworthiness.
+A key challenge in this context is ensuring that robots can provide meaningful justifications for their actions and decisions, particularly in safety-critical scenarios where their behavior may impact human well-being [1].
+Transparent and reliable explanations are essential [2], as they not only support effective human-robot interactions but also help build trust, particularly among non-expert users who depend on clear, accessible interpretations of robotic behavior [3, 4].
+However, the complexity of robotic decision-making processes and the need for thorough post-event analysis make it particularly difficult to achieve these goals.
+To address these challenges, accountability must be ensured using a robust tamper-proof logging system.
+The most effective evidence for post-event analysis comes from logging and auditing capabilities, which are essential for mitigating technical issues and providing forensic evidence to detect data tampering or unauthorized breaches.
+These systems must guarantee the integrity, confidentiality, and availability of data while enabling incident detection, compliance, and forensic investigation [5].
+However, covering tracks in log systems has become common practice for attackers, erasing any traces that could lead to their identification or assisting in the analysis of their actions.
+Therefore, it is necessary to strengthen the system properties to prevent unauthorized modifications [6].
+This includes safeguarding the integrity of log data and implementing measures to detect and counteract tampering attempts.
 Effective accountability solutions must ensure the detection and isolation of faulty behaviors and their origin.
@@
 Integrating accountability solutions into robotic systems often demands a trade-off with system performance.
-Specifically, computational power and networking bandwidth may be adversely affected [8].
+Specifically, the computational power and networking bandwidth may be adversely affected [8].
 The inclusion of anti-tampering techniques, such as full disk and communication encryption, can be restrictive due to resource limitations [9].
-Additionally, high-cost tamper-proof storage devices may prove economically unviable, especially when dealing with continuously generated log data.
+Additionally, high-cost tamper-proof storage devices may prove economically unviable, particularly when dealing with continuously generated log data.
 Therefore, it is necessary to establish an efficient and verifiable proof of integrity to guarantee the reproducibility and identification of the events that could arise during an autonomous agent action without compromising the robot’s performance.
 For this evidence to be significant, it should belong to a persistent chain of custody, fully reliable and verifiable, key requirements for an auditable and transparent logging system [10].
-To this end, blockchain solutions emerge as a mechanism to ensure data integrity, protection against unauthorized changes, and traceability.
-The consensus protocols established by all blockchain network members enable the detection and rejection of any destruction or manipulation of data through illicit channels.
-The nature of this technology mitigates tampering risk, safeguarding data [11].
-Nevertheless, while blockchain secures log data, the evidence obtained poses significant challenges for analysis due to its diverse and varied nature, terminology, and the presence of substantial volumes of similar messages unrelated to anomalous events.
+To this end, blockchain solutions have emerged as a mechanism to ensure data integrity, protection against unauthorized changes, and traceability.
+Consensus protocols established by all blockchain network members enable the detection and rejection of any destruction or manipulation of data through illicit channels.
+The nature of this technology mitigates tampering risk and safeguards data [11].
+Nevertheless, while blockchain secures log data, the evidence obtained poses significant challenges for analysis, due to its diverse and varied nature, terminology, and the presence of substantial volumes of similar messages unrelated to anomalous events.
 These factors may result in limited usefulness and misalignment with the expectations of non-expert users [12].
-The raw log messages produced by autonomous agents are usually noisy and semi-structured in format.
-In addition to this, it is necessary to deal with specific features such as large quantities of domain-specific language messages [13].
+The raw log messages produced by autonomous agents are typically noisy and semi-structured.
+In addition, it is necessary to deal with specific features such as large quantities of domain-specific language messages [13].
 Addressing this issue requires methods to organize and filter log data, making it accessible and meaningful.
-In this context, explaining the reasons behind a robot’s specific and unexpected actions is essential for diagnosing failures and comprehending deviations from intended goals.
-Equally important is ensuring these explanations are understandable to users [14].
+In this context, explaining the reasons behind a robot’s specific and unexpected actions is essential for diagnosing failures and comprehending deviations from the intended goals.
+Equally important is ensuring that these explanations are understandable to users.
 Explainable Artificial Intelligence (XAI) is a research area dedicated to enhancing the transparency and interpretability of Artificial Intelligence (AI) systems to ensure their responsible and ethical deployment.
@@
 This requirement has led to the development of eXplainable Autonomous Robots (XAR), which aims to provide mechanisms for presenting clear and understandable explanations for the actions and decisions of autonomous agents [14].
-XAR enhances the transparency and interpretability of these agents’ decision-making processes, often incorporating techniques like natural language explanations to improve clarity and user comprehension [16, 17].
-To further improve the effectiveness of XAR systems, Large Language Models (LLMs) excel in diverse natural language processing tasks, including translation, coding, or text summarization [18, 19, 20].
+XAR enhances the transparency and interpretability of these agents’ decision-making processes, often incorporating techniques such as natural language explanations to improve clarity and user comprehension [15, 16].
+To further improve the effectiveness of XAR systems, Large Language Models (LLMs) are used in diverse natural language processing tasks, including command translation and task planning [17].
 Our approach integrates LLMs with blockchain-secured log data.
-By using retrieval-based techniques, this solution generates accurate and context-aware natural language explanations based on verifiable information [21].
+Using retrieval-based techniques, this solution generates accurate and context-aware natural language explanations based on verifiable information [18].
 This integration ensures that explanations are clear, coherent, and grounded in a traceable and reliable context, effectively addressing user comprehension and data integrity concerns.
-This work presents an accountability and explainability architecture implemented for Robot Operating System (ROS)-based mobile robots.
-This solution includes two main components.
-First, a black box-like component to provide accountability that includes anti-tampering features by distributing integrity proofs from the events recorded by the black box component.
-The second component generates natural language explanations from data in the black box, including the use of interactive natural language conversations as a means to provide comprehensive explainability for the recorded events.
-Both components are connected through an additional module responsible for processing raw data collected by the first one, eliminating non-essential information, and building connections and relations between data.
-These functionalities release software developers from the need to make calls to the logging Application Programming Interface (API) to build explanations for the code generated and the behaviors developed by the autonomous agent.
-To validate our approach, we conducted a comprehensive evaluation, assessing the architecture’s impact on system performance and analyzing the quality of the generated explanations.
-The evaluation considered key metrics, including correctness, accuracy, helpfulness, coherence, and user satisfaction.
-Our results demonstrate that the architecture effectively balances accountability and explainability without compromising the robot’s performance, providing a scalable and reliable framework for real-world applications.
-The remainder of this paper is organized as follows: Section 2 provides an overview of previous works and findings related to the study, focusing on the enhancement of immutable data recorders and explainability solutions for robots.
-Section 3 delves into the system architecture, offering a detailed description of the designed approaches.
-Section 4 details the technical elements of the designed system attending to the proposed architecture.
-Section 5 discusses the experimental process, the scenarios and the variables employed for assessment.
-Section 6 outlines the obtained results, particularly addressing accountability and explainability metrics.
-Finally, Section 7 summarises the obtained conclusions.
-2 Related Works
+Although recent studies have advanced accountability and explainability in autonomous robotic systems, many existing approaches still face important limitations.
+Several rely on fixed templates or static rule-based methods that lack flexibility to adapt to dynamic and unpredictable environments.
+Consequently, the explanations produced may be generic, rigid, or poorly aligned with the actual context of the robot’s behavior.
+Moreover, explainability techniques are often developed separately from accountability mechanisms, leading to explanations not grounded in verifiable data.
+This disconnect limits their reliability and effectiveness, particularly in safety-critical settings where trust and traceability are essential.
+In addition, current evaluation methods for these explanations face significant challenges, including the absence of standardized and domain-specific metrics, biases in the generated text, and the high cost of human evaluations [19].
+Even though human evaluations remain the gold standard for capturing subjective insights [20], they are resource-intensive and lack scalability [21].
+Alternatively, LLM-based evaluation methods offer a scalable alternative, demonstrating over 80% alignment with human judgments while offering interpretable and explainable assessments [22, 23].
+I-A Contributions
+This work presents an accountability and explainability architecture designed for Robot Operating System (ROS)-based mobile robots.
+This solution integrates two main components.
+The first is a black box-like module that provides accountability through anti-tampering features by distributing integrity proofs from the events recorded by the black box.
+The second component generates natural language explanations based on black box data, including interactive natural language conversations as a means of providing comprehensive explainability for the recorded events.
+Both components are connected through an additional module responsible for processing the raw data collected by the first module, eliminating non-essential information, and building connections and relations between data.
+These functionalities release software developers from the need to make calls to the logging Application Programming Interface (API) to build explanations to document the code and behaviors of the autonomous agent.
+To validate our approach, we conducted a comprehensive evaluation to examine both the impact of the architecture on system performance and the quality of the generated explanations.
+In order to capture detailed insights while maintaining efficiency and reproducibility, we developed a hybrid evaluation methodology that combines human judgments with LLM-based assessments.
+The evaluation focused on key criteria such as correctness, accuracy, helpfulness, coherence, and user satisfaction.
+This integrated framework provides a robust, scalable, and reliable means of assessing LLM-generated explanations in robotic systems.
+Our findings demonstrate that the proposed architecture successfully balances accountability and explainability without compromising robotic performance, thereby offering a practical and scalable solution for deployment in real-world applications.
+Therefore, the primary contributions of this work are summarized as follows:
+  1. 1.
+    A unified architecture that combines blockchain-based tamper-proof logging with LLMs to generate grounded natural language explanations of robot behavior, enabling accountability and transparency.
+  2. 2.
+    Real-time integrity and verifiability of robotic event data through a blockchain-based black box component, with minimal impact on robot performance, even at high data rates, addressing a critical requirement for practical deployment.
+  3. 3.
+    A log curation and processing module that filters noisy, semi-structured robotic logs and extracts meaningful context, significantly improving the relevance, structure, and coherence of the generated explanations.
+  4. 4.
+    An empirical analysis of the trade-offs between system performance and logging granularity, offering practical design insights for developing robust, interpretable, and efficient autonomous agents
+  5. 5.
+    A hybrid and comprehensive evaluation methodology combining LLM-as-judge and human-based assessments, designed to leverage the strengths of both approaches in validating the quality, accuracy, and user satisfaction of the generated explanations for complex robotic behaviors such as goal cancellations, re-planning, and obstacle avoidance.
+I-B Structure of the Paper
+The remainder of this paper is organized as follows: Section II provides an overview of previous studies and findings related to the study, focusing on the enhancement of immutable data recorders and explainability solutions for robots.
+Section III delves into the system architecture, offering a detailed description of the designed approaches.
+Section IV describes the technical elements of the system designed for the proposed architecture.
+Section V presents the evaluation process, including the scenarios and assessment variables.
+Section VI outlines the results obtained, particularly addressing accountability and explainability metrics.
+Finally, Section VII summarizes the obtained conclusions.
+II Related Works
 In this section, we explore key works related to various domains relevant to our approach.
 This includes advances in how blockchain ensures the immutability of logs, research on applying Event Data Recorders (EDRs) in robotics, significant contributions to integrating XAI with blockchain, and XAR.
-Finally, we examine the evolution of LLMs for generating natural language explanations and their advancements within robotics.
-2.1 Blockchain for Immutable Logs
+Finally, we examine the evolution of LLMs to generate natural language explanations and their advancements in robotics.
+II-A Blockchain for Immutable Logs
 Blockchain is a decentralized solution that eliminates the necessity for trust in a central authority through its immutable and distributed ledger, composed of timestamped transaction blocks.
-These blocks are linked through hashing and the storage of previous block content, ensuring the integrity of transactions.
-Any attempt to alter a transaction in one block requires the modification of all subsequent blocks, resulting in a computationally expensive process.
-As a result, data stored in the blockchain is immutable [22, 23].
-There are two primary types of blockchains based on access controls: permissionless blockchains allow open access and public transaction visibility, while permissioned blockchains limit participation and transaction visibility.
-Nevertheless, closed networks may pose security risks, as authorized users could recalculate subsequent blocks in the event of a block modification [24].
-While permissionless blockchains have traditionally related to scalability and throughput challenges, approaches like Ethereum have transitioned to Proof of Stake (PoS), a more energy-efficient, secure, accessible, and scalable consensus protocol [25, 26].
-Blockchain technology ensures log integrity and auditability, offering unique advantages over alternatives such as distributed storage systems or InterPlanetary File Systems (IPFSs), which may require central authorities for storage and verification [27, 28].
+These blocks are linked through the hashing and storage of the previous block content, ensuring the integrity of transactions.
+Any attempt to alter a transaction in one block requires modification of all subsequent blocks, resulting in a computationally expensive process.
+As a result, the data stored in the blockchain is immutable [24, 25].
+There are two primary types of blockchains based on access controls: permissionless blockchains allow open access and public transaction visibility, whereas permissioned blockchains limit participation and transaction visibility.
+Nevertheless, closed networks may pose security risks, as authorized users could recalculate subsequent blocks in the event of block modification [26].
+While permissionless blockchains have traditionally been related to scalability and throughput challenges, approaches such as Ethereum have transitioned to Proof of Stake (PoS), a more energy-efficient, secure, accessible, and scalable consensus protocol [27, 28].
+Blockchain technology ensures log integrity and auditability, offering unique advantages over alternatives such as distributed storage systems or InterPlanetary File Systems (IPFSs), which may require central authorities for storage and verification [29].
 Blockchain offers consensus and immutability, coupled with the capability of integrating logic mechanisms such as Smart Contracts (SCs).
-SCs contain functions that either write values onto the blockchain, thus altering its state, or read data from it [29].
-These features make blockchain the preferred choice for applications where data integrity and trust are required [30].
+SCs contain functions that either write values onto the blockchain, thus altering its state, or reading data from it [30].
+These features make blockchain the preferred choice for applications in which data integrity and trust are required [31].
 Different secure logging systems rely on permissioned blockchain frameworks for log immutability [32], introducing accountability even in embedded systems [7].
-To get a balance between throughput and cost-effectiveness, some works combine both categories [32], developing approaches that include a permissioned blockchain to ensure per-entry immutability and public non-repudiation solution [33].
-Nevertheless, inherent features of permissionless blockchain such as transparency and trust, have also boosted the development of secure logging systems [34].
-Our architecture uses the immutability of blockchain to store evidence of the integrity of recorded data, thus providing real-time guarantees against manipulation and ensuring that the explanations generated by our system are always based on non-manipulated data.
-2.2 Blockchain for Event Data Recorders in Robotics
+To achieve a balance between throughput and cost-effectiveness, some works combine both categories [33], developing approaches that include a permissioned blockchain to ensure per-entry immutability and a public non-repudiation solution [34].
+Nevertheless, inherent features of permissionless blockchains such as transparency and trust, have boosted the development of secure logging systems [35].
+While previous work often records cryptographic hashes on the blockchain to ensure log integrity and employs SCs for data management, our approach goes further by using these verified records as contextual input for generating natural language explanations of autonomous robot behavior.
+Additionally, our system includes features that enable public accountability and prevent denial of recorded actions.
+This design provides strong guarantees against data manipulation and ensures that the explanations produced by our system are always based on trustworthy, tamper-proof information.
+By bridging immutable records with interpretable explanations, our method enhances transparency and clarity in communicating robot actions.
+II-B Blockchain for Event Data Recorders in Robotics
 EDRs, or black boxes, are essential for capturing and analyzing events in robotic systems.
-These recorders help to comprehend system behavior, detect anomalies, and ensure safety [35, 36].
-Therefore, these solutions are needed in forensic investigations and ethical assessments regarding autonomous agents [37].
+These recorders help comprehend system behavior, detect anomalies, and ensure safety [36, 37].
+Therefore, these solutions are required for forensic investigations and ethical assessments of autonomous agents [38].
 The integration of black box modules to enhance robots’ introspection capabilities contributes to improving system transparency by attributing responsibilities [39, 9].
-These solutions typically include functionalities such as hashing, signing, and encryption [39], while ensuring event integrity through anti-tamper mechanisms [40].
-Using blockchain technology in EDRs ensures the integrity of a transparent and auditable record of events.
+These solutions typically include hashing, signing, and encryption [40], while ensuring event integrity through anti-tamper mechanisms [41].
+The blockchain technology in EDRs ensures the integrity of a transparent and auditable record of events.
 These features enhance accountability and trust in the recorded data by preventing tampering [42].
@@
 In [45], a time-segmented consortium blockchain for robotic event registration is proposed, enabling efficient event organization and retrieval within specific time frames.
-Authors in [45] present a tool to bridge ROS and the Ethereum blockchain, evaluating encryption efficiency and stability, transaction response time, and packet loss rate.
-Nevertheless, these works lack the capability to save integrity proofs in the blockchain at selected intervals, enhance granularity, or include service definitions or SCs to improve the verifiability process.
-2.3 Explainable Artificial Intelligence and Blockchain
-XAI aims to align AI decisions with human cognitive expectations by employing techniques such as contextual relationships, contrastive explanations, and methods like counterfactual reasoning and interpretable models that enhance interpretability and transparency [46, 47].
-Insights from social sciences show people tend to interpret explanations more positively when they reflect human-like reasoning, emphasizing intentions, goals, and contextual factors rather than purely technical or statistical descriptions [48].
-By combining XAI with blockchain transparent record-keeping, explanations can be retraced and validated, ensuring both interpretability and accountability in AI-driven systems [49, 50].
-The benefits of this integration are remarkable in various domains.
-In healthcare, XAI and blockchain have been combined for secure, transparent patient data management [51], logical disease prediction, and immutable medical diagnostics records [52].
-Similarly, in the legal domain applications, blockchain and XAI support tamper-proof decision auditing, using IPFS and Ethereum for cryptographic proofs that enable data verification and accountability [53, 54, 28].
-Privacy-preserving systems with homomorphic encryption and semantic ontologies further ensure data protection while maintaining explainability [55].
-Regarding autonomous agents like unmanned aerial vehicles or autonomous vehicles, blockchain and XAI ensure secure and transparent decision-making [56].
+In [46], the authors present a tool to bridge ROS and Ethereum blockchain, evaluating encryption efficiency and stability, transaction response time, and packet loss rate.
+Existing blockchain-based logging solutions in robotics usually focus on low-level data storage and network performance metrics, and often lack mechanisms to record integrity proofs of structured and filtered messages at configurable intervals.
+Furthermore, these solutions generally miss granular recording policies and the use of SCs to improve data verifiability.
+In contrast, our approach guarantees tamper-proof evidence by periodically anchoring integrity proofs of curated log data on the blockchain, while also generating natural language explanations grounded in verifiable and filtered information.
+This integration significantly improves both the accountability of robotic systems and the comprehensibility of the explanations provided to users in dynamic contexts.
+II-C Explainable Artificial Intelligence and Blockchain
+XAI aims to align AI decisions with human cognitive expectations by employing techniques such as contextual relationships, contrastive explanations, counterfactual reasoning, and interpretable models that enhance interpretability and transparency [47, 48].
+Insights from social sciences show that people tend to interpret explanations more positively when they reflect human-like reasoning, emphasizing intentions, goals, and contextual factors rather than purely technical or statistical descriptions [20].
+By combining XAI with transparent blockchain record-keeping, explanations can be retraced and validated, ensuring both interpretability and accountability in AI-driven systems [49, 50].
+The benefits of this integration are remarkable in several domains.
+In healthcare, XAI and blockchain have been combined for secure, transparent patient data management, logical disease prediction, and immutable medical diagnostic records [51].
+Similarly, in legal domain applications, blockchain and XAI support tamper-proof decision auditing, using IPFS and Ethereum for cryptographic proofs that enable data verification and accountability [52, 29].
+Privacy-preserving systems with homomorphic encryption and semantic ontologies further ensure data protection, while maintaining explainability [53].
+Regarding autonomous agents, such as unmanned aerial vehicles or autonomous vehicles, blockchain and XAI ensure secure and transparent decision-making [54].
 This integration enhances collaboration, even in complex and time-critical scenarios [55].
-Decentralized edge models for collaborative AI, leveraging blockchain for secure communication and real-time data sharing, further optimize resources and enhance reliability in these environments [58].
-Our work introduces Rosbag2 files as an alternative to IPFS or distributed file systems for storing contextual data in autonomous robotic systems.
+Decentralized edge models for collaborative AI leverage blockchain for secure communication and real-time data sharing, further optimizing resources, and enhancing reliability in these environments [56].
+Despite previous advances, providing explanations grounded in robot event data remains challenging due to the complexity, volume, and heterogeneity of the generated information.
+To address this, our proposal introduces Rosbag2 files as an alternative to IPFS or distributed file systems for storing contextual data in autonomous robotic systems.
 This approach supports reproducibility and enables a more precise analysis of robotic behaviors [43].
 Rosbag files provide a standardized format for recording large volumes of robotic data, facilitating comprehensive interaction capture for improved analysis.
-To address the indeterminism of LLM algorithms, we propose using a permissionless blockchain to store integrity proofs of Rosbag messages, ensuring immutability, simplifying verification, and enhancing data reliability for explanation generation.
-2.4 Explainable Autonomous Robots
-XARs are designed to provide explanations for their actions and decisions to non-expert users.
+To mitigate the inherent indeterminism of LLM algorithms, we propose using a permissionless blockchain to store integrity proofs of Rosbag messages, ensuring data immutability, simplifying verification, and increasing reliability in explanation generation.
+II-D Explainable Autonomous Robots
+XARs are designed to provide explanations of their actions and decisions to non-expert users.
 Such explanations enhance security and trust, helping to prevent failures, errors, conflicts, and undesired interactions while achieving objectives [57, 58].
-However, despite the growing demand for explainability in autonomous agents, their use reveals a lack of practical usability and effective interpretability in real environments [61].
-Effective explanations in XARs should detail alternatives, the choices made, expectations for each option, decision criteria, and any unexpected events that may have led to changes in the planned behavior [17, 62].
+However, despite the growing demand for explainability in autonomous agents, their use lacks practical usability and interpretability in real environments [59].
+Effective explanations in XARs should detail the alternatives, choices made, expectations for each option, decision criteria, and any unexpected events that may have led to changes in the planned behavior [16, 60].
 These explanations often rely on abductive reasoning, which bridges the gap between the robot’s internal decision-making process and the user’s understanding [61].
-To support this, frameworks have been developed that enable robots to articulate their internal states and decisions, enhancing collaboration and improving communication with humans [64].
+To support this, frameworks have been developed that enable robots to articulate their internal states and decisions, enhance collaboration and improve communication with humans [62].
 Adaptive algorithms and needs-based motivational systems allow robots to provide real-time, context-aware explanations tailored to diverse audiences [63, 64].
 Behavior Trees (BTs) hierarchically decompose tasks into goals and subgoals, making complex behaviors easier to understand and adaptable to real-time changes [65].
-However, achieving a balance between simplicity and adequacy is essential when making complex information easy to understand.
-By focusing on the most relevant details and filtering non-essential information, explanations foster trust and collaboration by aligning with human cognitive capacities [68].
-For this purpose, our architecture incorporates a log curation module that processes raw robotic data, filtering it to retain only meaningful, high-relevance information, thus delivering clearer and more concise explanations directly tied to events.
-In addition to context-aware explanations, the core of explainability in XAR lies in effective communication [63].
-Progress in Natural Language Processing (NLP) and LLMs can improve the efficiency of human-robot interactions.
-This enhancement can create a sense that these systems truly comprehend and respond to individuals meaningfully, significantly improving their usefulness and applicability [1].
-2.5 Large Language Models in Robotics
-LLMs, enabled by self-attention mechanisms, have advanced NLP, improving task-solving capabilities and enabling more effective human-robot interaction [69, 70, 71].
+Nevertheless, achieving a balance between simplicity and adequacy remains a key challenge when conveying complex or technical information in an understandable way.
+Explanations that are overly detailed can overwhelm users, while those that are too simplified may omit critical context necessary for constructing accurate and meaningful interpretations [66].
+Our architecture addresses these limitations by incorporating a log curation module that automatically processes and filters raw robotic data, retaining only the most relevant information for explanation generation.
+By emphasizing essential details and excluding non-essential information, our system produces explanations that align with human cognitive capacities, thereby fostering trust and facilitating effective collaboration.
+In addition to providing context-aware explanations, effective explainability in XAR fundamentally relies on clear communication [61].
+While prior research has acknowledged the potential of Natural Language Processing (NLP) and LLMs to enhance human-robot communication [61, 1], their integration with robot-specific data streams remains limited.
+This limitation arises because LLMs primarily depend on probabilistic patterns rather than direct access to data from the robot, which can lead to responses that are linguistically plausible yet factually inaccurate.
+Our approach addresses this challenge by grounding explanations in curated robotic data and leveraging the natural language capabilities of LLMs, thereby producing clear and accessible explanations that enhance transparency and foster trust between users and autonomous systems, including non-expert audiences.
+II-E Large Language Models in Robotics
+LLMs, enabled by self-attention mechanisms, have advanced NLP, improving task-solving capabilities and enabling more effective human-robot interaction [67, 68].
 However, LLMs face challenges such as hallucinations, which arise due to reliance on internal knowledge, and the high update costs associated with large parameter sizes, making traditional pre-training and fine-tuning methods impractical.
@@
 This approach is particularly advantageous in robotics, where real-time and accurate contextual understanding is essential for complex decision-making and task execution.
-Explainability and transparency have become critical aspects of deploying LLMs in autonomous agents.
-These models enable robots to translate high-level commands into actionable plans, improving performance in complex environments [74, 75, 76].
+Explainability and transparency have become critical for deploying LLMs in autonomous agents.
+These models enable robots to translate high-level commands into actionable plans, improving their performance in complex environments [71, 72, 73].
 When integrated with reward decomposition frameworks, LLMs can generate intuitive explanations, linking actions to object-specific properties, thereby reducing ambiguity [74].
 LLMs can also generate natural language descriptions of robot behaviors and logs, highlighting opportunities for improvement in accuracy and completeness [75, 76].
-Existing solutions, however, often rely on static models or predefined data, limiting their ability to handle dynamic environments.
-Our approach addresses these limitations by integrating blockchain-supported accountability with curated logging of relevant robotic data.
-This framework provides robust context for LLM-generated explanations, minimizing hallucinations and ensuring alignment with the robot’s actions and interactions within its environment.
-Evaluating LLMs in robotics presents additional challenges, including the absence of suitable metrics, biases in generated text, and the high cost of human evaluations [80].
-While human evaluations remain the gold standard for capturing subjective insights [48], they are resource-intensive and lack scalability [81].
-LLM-based evaluations offer a scalable alternative, achieving over 80% alignment with human judgments while delivering explainable outputs [82].
-A hybrid framework combining human assessments for nuanced insights and LLM evaluations for efficiency and repeatability provides a comprehensive and scalable solution for evaluating LLM outputs.
-3 System Architecture
+However, existing solutions often rely on static models or predefined data, which limit their ability to handle dynamic environments.
+Furthermore, these solutions rarely provide mechanisms for ground explanations in verifiable, tamper-proof evidence of robot behaviors.
+Our approach addresses these limitations by integrating blockchain-supported accountability with the curated logging of relevant robotic data.
+This framework provides a robust context for LLM-generated explanations, minimizes hallucinations, and ensures alignment with the robot’s actions and interactions within its environment.
+III System Architecture
 Our approach comprises two main components: a tamper-proof black box-like device, and a module responsible for generating natural language explanations for the data stored by the former.
-The first component offers services for building, storing, and verifying integrity proofs through SC specifications.
-The explainability component provides natural language explanations by implementing RAG, thereby enhancing the obtained results using the data recorded by the first component as an external source to enrich the context in the answer formulation process.
+The first component offers services for building, storing, and verifying integrity proofs through the SC specifications.
+The explainability component provides natural language explanations by implementing RAG, thereby enhancing the results obtained using the data recorded by the first component as an external source to enrich the context in the answer formulation process.
 Both components are integrated through the existence of two main asynchronous tasks focused on cleaning and processing the raw data generated by the black box-like device to increase the effectiveness of the explainability engine.
-Figure 1 depicts a generic architecture of our proposal, highlighting in blue color the functionalities related to accountability, and in green those related to explainability.
-Below we provide a deeper description of the approaches designed and followed in order to develop and integrate the above-mentioned components.
+Figure 1 depicts the generic architecture of our proposal, highlighting the functionalities related to accountability in blue, and those related to explainability in green.
+Below, we provide a deeper description of the approaches designed and followed to develop and integrate the above-mentioned components.
 Figure 1: System architecture.
-3.1 Black Box Recorder
-Our first component focuses on recording accountable information that captures essential messages generated during the navigation task of an autonomous agent.
-The selected data will guarantee the reproducibility of the mentioned task, including details like the robot’s poses, the employed map, odometry data, linear and angular velocity, and laser scan readings.
-The immutability of the previously described information is achieved through blockchain.
-Blockchain technology usually faces limitations regarding the size of the data that can be written into a single transaction.
-Efficiently storing autonomous agent messages becomes challenging due to these constraints.
-Our proposed solution involves selectively storing the hash of messages at intervals, providing in addition the chance of using different storage frequencies depending on the topic and its relevance.
+III-A Black Box Recorder
+The first component focuses on recording accountable information that captures the essential messages generated during the navigation task of an autonomous agent.
+The selected data will guarantee the reproducibility of the mentioned task, including details such as the robot’s poses, employed map, odometry data, linear and angular velocity, and laser scan readings.
+The immutability of the previously described information is achieved using blockchain.
+Blockchain technology usually faces limitations in terms of the size of the data that can be written into a single transaction.
+Efficient storage of autonomous agent messages is challenging owing to these constraints.
+Our proposed solution involves selectively storing the hashes of messages at intervals, providing the chance to use different storage frequencies depending on the topic and its relevance.
 This fact not only enhances the granularity of stored message proofs but also provides a more adaptable solution.
-Our anti-tampering approach for this accountability proposal comprises two key components: a chain of integrity proofs coupled with the SC specification.
-This section delves into their design and justifies how both align with the constraints of autonomous agents.
-Key features such as consistency are guaranteed by the blockchain’s consensus protocol, where the content of each block is hashed and stored in the next one.
-Additionally, availability is also assured through the distributed and decentralized nature of the blockchain, ensuring accessibility across the entire network.
-3.1.1 Chain of Integrity Proofs
+Our anti-tampering approach to this accountability proposal comprises two key components: a chain of integrity proofs coupled with an SC specification.
+This section delves into their design and explains how both align with the constraints of the autonomous agents.
+Key features such as consistency are guaranteed by blockchain’s consensus protocol, in which the content of each block is hashed and stored in the next one.
+In addition, availability is assured by the distributed and decentralized nature of the blockchain, ensuring accessibility across the entire network.
+III-A1 Chain of Integrity Proofs
 To preserve the integrity of the messages without compromising system performance, our solution entails building a chain where each message hash is linked with the previous hash value.
-This approach ensures that any alteration of the selected messages will result in a different final hash, easing tampering detection.
-The verification process is also simplified since only the last hash in the chain needs to be recomputed and compared with the last stored value.
-A single modification of a hash will disrupt the entire chain from that point onward, simplifying the identification of tampering.
+This approach ensures that any alteration of the selected messages results in a different final hash, easing tampering detection.
+The verification process is also simplified because only the last hash in the chain must be recomputed and compared with the last stored value.
+A single hash modification disrupts the entire chain from that point onward, simplifying the identification of tampering.
 We define a chain entry ($CE$) with a counter $i\in 1..m$ to be linked with the previous one by concatenating the prior digest $h_{i-1}$ with the current log message $Msg_{i}$.
-hash() denotes the cryptographic hash function, with SHA-256 being the chosen one, commonly used for data integrity verification and digital signatures.
+hash() denotes the cryptographic hash function, with SHA-256 being the chosen one, which is commonly used for data integrity verification and digital signatures.
 The initial hash value $h_{0}$ of the chain is calculated by applying the selected hash function to a set of $l=32$ random bytes ($RB$), to generate a 256-bit value.
-Introducing initial entropy into the chain of integrity proofs enhances privacy, ensuring that integrity proofs for similar initial log messages from different records are not identical.
-Data privacy is further prioritized by storing in the blockchain only the hashes, rather than the contents of the messages.
+Introducing initial entropy into the chain of integrity proofs enhances privacy, ensuring that the integrity proofs for similar initial log messages from different records are not identical.
+Data privacy is further prioritized by storing only the hashes in the blockchain, rather than the content of the messages.
 $CE_{i}$ $=h_{i}$ (1)
@@
 $h_{i}$ $=\text{hash}(h_{i-1}||Msg_{i})$
-By chaining every hash with the previous one, we optimize the validation of stored messages because the last linked digest should correspond to the final proof saved in the blockchain.
+By chaining every hash with the previous one, we optimize the validation of the stored messages because the last linked digest should correspond to the final proof saved in the blockchain.
 Additionally, a service has been developed to generate the chain of hashes using the recorded data, ensuring partial validation of the integrity proofs without modifying the stored messages or adding non-essential information to the transactions saved in the blockchain.
 The number of proofs included in a blockchain transaction depends on the transaction cost limit set by the blockchain framework.
-Therefore, being $j\in 1..m$ a transaction counter, each newly arriving chain entry is added in a transaction payload $P_{j}$ until the limit is reached.
-Being sign() a shorthand for a signature function and $PrivKey$ the private key of the account, transaction $T_{j}$ is transmitted from the service in charge of recording the accountable data to the blockchain.
+Therefore, being $j\in 1..m$ a transaction counter, each newly arriving chain entry is added to a transaction payload $P_{j}$ until the limit is reached.
+Being sign() a shorthand for a signature function and $PrivKey$ the private key of the account, transaction $T_{j}$ is transmitted from the service in charge of recording accountable data to the blockchain.
 $T_{j}$ $=(P_{j},S_{j})$ (2)
@@
 A service interruption or a disruption in the behavior of the autonomous agent will lead to the aggregation of the remaining hash values into a single transaction, even if there is capacity for additional values.
-This practice guarantees the immutability of the last messages sent by a robot upon completing a task or achieving a goal.
-3.1.2 Smart Contract
-Calling an SC function intended for blockchain writing involves the execution of a transaction that changes the state of the storage data on the blockchain.
+This practice guarantees the immutability of the last messages sent by the robot upon completing a task or achieving a goal.
+III-A2 Smart Contract
+Calling an SC function intended for blockchain writing involves the execution of a transaction that changes the state of storage data on the blockchain.
 In our approach, this transaction is signed with the account’s private key via the Elliptic Curve Digital Signature Algorithm, effectively identifying the signer.
@@
 Moreover, our SC includes a function to verify the existence of a message proof on the blockchain.
-This fact allows external parties to independently confirm whether a specific message hash has been stored in the contract and when it was stored.
-This functionality contributes to the system’s transparency, enabling users or auditors to check the validity of message proofs without relying solely on the contract owner, thereby promoting public accountability.
-3.2 Explainability Component
+This allows external parties to independently confirm whether a specific message hash has been stored in the contract and when it is stored.
+This functionality contributes to the transparency of the system, enabling users or auditors to check the validity of message proofs without relying solely on the contract owner, thereby promoting public accountability.
+III-B Explainability Component
 Effective communication is essential for achieving explainability in autonomous robots.
-Our approach is focused on improving this feature through the use of NLP technology, enabling non-expert users to interact with agents in a meaningful way.
+Our approach focuses on improving this feature through the use of NLP technology, enabling non-expert users to interact with agents in a meaningful way.
 Our solution involves the use of RAG to retrieve data from outside the language model (non-parametric) and augment the prompts by adding the information recorded from our accountability component in context.
-Before this integration, the information undergoes processing by a dedicated module responsible for acquiring and processing the raw data generated by the autonomous agent, serving as contextual input in the question-answering system.
-Given the volume of messages, their varied nature, their length, the presence of non-meaningful records generated by the autonomous agent, and the need to establish specific relationships between messages to enhance the context of the engine, the log curation component extracts relevant information.
-This functionality serves to enhance the effectiveness of explanations provided to non-expert users.
-A more comprehensive description of this component is provided in Section 4.
-After processing the raw data stored by our accountability component, the ingestion process performs the chunking task to break down extensive text segments into more manageable units.
-To ensure comparability during relevancy searches at inference time, both this data and the user query are converted into numerical representations through embedding language models.
-Next, based on the embedding of the user query, its relevant text is identified in the document collection by a similarity search in the embedding space.
-The user-provided prompt is then extended with relevant information obtained in the search and added to the context.
+Before this integration, the information is processed by a dedicated module responsible for acquiring and processing the raw data generated by the autonomous agent, which serves as contextual input in the question-answering system.
+Given the volume of messages, their varied nature, their length, the presence of non-meaningful records generated by the autonomous agent, and the need to establish specific relationships between messages to enhance the context of the engine, the log curation component extracts the relevant information.
+This functionality enhances the effectiveness of the explanations provided to non-expert users.
+A more comprehensive description of this component is provided in Section IV.
+After processing the raw data stored by our accountability component, the ingestion process performs a chunking task to break down extensive text segments into more manageable units.
+To ensure comparability during relevancy searches at inference time, both the data and user query are converted into numerical representations by embedding language models.
+Next, based on the embedding of the user query, the relevant text is identified in the document collection using a similarity search in the embedding space.
+The user-provided prompt is then extended with the relevant information obtained in the search and added to the context.
 This enhanced prompt is sent to the LLM.
-As the context contains pertinent external data along with the original prompt, the model output is relevant and accurate.
-The addition of new accountable data to our explainability component requires asynchronous updates of the accountable data and its embedding representation.
-These updates ensure that the model remains sensitive to changes, addressing limitations introduced by offline training.
-4 Implementation
-The solution presented in Section 3 can be implemented using different technologies. This section details the specific solutions chosen for developing our proposal.
-4.1 Black Box Recorder
+Because the context contains pertinent external data along with the original prompt, the model output is relevant and accurate.
+The addition of new accountable data to our explainability component requires asynchronous updates of accountable data and its embedding representation.
+These updates ensure that the model is sensitive to changes, thereby addressing the limitations introduced by offline training.
+IV Implementation
+The solution presented in Section III can be implemented using various technologies. This section details the specific solutions used to develop our proposal.
+IV-A Black Box Recorder
 ROS is the most widely used framework in robotics, offering a collection of software libraries and tools for developing robot applications.
 Nevertheless, limitations in fulfilling the requirements of real-time embedded systems and the necessity to enhance multi-platform support led to a major upgrade to ROS 2, whose main change lies in the adoption of the Data Distribution Service (DDS) [77].
-Nodes constitute the fundamental components in both ROS and ROS 2 applications.
-These nodes are individual processes that handle specific tasks and communicate with each other through messages.
+Nodes constitute fundamental components of both ROS and ROS 2 applications.
+These nodes are individual processes that handle specific tasks and communicate with one another through messages.
 Data transfer between nodes can be achieved using various methods, such as topics, services, and actions.
-Topics facilitate message passing, services enable synchronous request-reply interactions, and actions are utilized for long-running tasks that need feedback.
-This work focuses on navigation capabilities.
-For this purpose, we utilized Nav2, a suite of tools for ROS 2 that is regarded as the successor to the ROS Navigation Stack.
+Topics facilitate message passing, services enable synchronous request-reply interactions, and actions are utilized for long-running tasks that require feedback.
+This study focuses on navigation capabilities.
+For this purpose, we utilized Nav2, a suite of tools for ROS 2 that is regarded as the successor of the ROS Navigation Stack.
 Nav2 equips the robot with the capacity to execute complex navigation tasks.
 Nav2’s core functionalities include planning, control, localization, visualization, and additional features.
-Furthermore, it employs BTs to create customized navigation behavior by orchestrating multiple independent modular servers, offering high flexibility in navigation tasks and the specification of complex robot behaviors [84].
-Different independent task servers, intended to compute a path, control effort, or recovery, are communicated with the BT over a ROS 2 interface, such as an action server or service.
-The use of BTs to generate explanations for robot behavior has been previously successfully evaluated, allowing seeking causal information to answer questions [85].
-Within ROS 2, nodes communicate by publishing data on topics, which can be consumed by other nodes.
-The data itself can range from sensor readings that provide the robot’s perception of its environment to control commands sent to actuators from decision-making nodes.
-Navigation reproducibility of autonomous robots in ROS 2 involves storing essential data, such as their poses, maps, odometry, velocity, and laser scan readings.
-To this aim, we have used Rosbag files, designed to capture and store data generated during the execution of a ROS system.
-These files can reproduce the exact conditions and inputs from a previous run, making them particularly valuable for understanding system behavior.
-The set of topics included in the Rosbag files generated are detailed in Table 1.
-Table 1: Topics name included in the Rosbag file their description.
+Furthermore, it employs BTs to create customized navigation behavior by orchestrating multiple independent modular servers, offering high flexibility in navigation tasks and specifications of complex robot behaviors [78].
+Different independent task servers, intended to compute a path, control effort, or recovery, communicate with the BT over a ROS 2 interface, such as an action server or service.
+The use of BTs to generate explanations for robot behavior has previously been successfully evaluated, allowing causal information to answer questions [79].
+In ROS 2, nodes communicate by publishing data on topics that can be consumed by other nodes.
+The data can range from sensor readings that provide the robot’s perception of its environment to control commands sent to actuators from decision-making nodes.
+The navigation reproducibility of autonomous robots in ROS 2 involves storing essential data such as their poses, maps, odometry, velocity, and laser scan readings.
+To this end, we used Rosbag files, designed to capture and store data generated during the execution of a ROS system.
+These files can reproduce the exact conditions and inputs from a previous run, making them particularly valuable for understanding the system behavior.
+The set of topics included in the generated Rosbag files generated are detailed in Table I.
+The selected topics capture the core aspects of the robot’s operation, including localization, environment perception, navigation status, decision-making logic, and motion control.
+This selection ensures that the recorded data provides a coherent and sufficient basis for interpreting and reproducing the robot’s navigation behavior.
+TABLE I: Topics name included in the Rosbag file their description.
 Topic name 1 Information provided
@@
 tf Coordinate frame transformations in the robotic system.
-To ensure the immutability of generated Rosbag files, Ethereum has been selected as the blockchain framework.
+Ethereum has been selected as the blockchain framework to ensure the immutability of generated Rosbag files.
 Ethereum, an open-source public blockchain, introduces the cryptocurrency Ether and self-executing programs (SCs) for code execution on the blockchain.
-Ethereum’s open-source nature ensures decentralization and accessibility, allowing participation without needing approval, a feature crucial for maintaining the integrity of recorded data.
-Additionally, Ethereum’s use of PoS as a consensus protocol enhances energy efficiency and throughput compared to traditional protocols such as Proof of Work.
-In addition, Ethereum’s scalability, security and economic efficiency make it a very secure option for ensuring data immutability.
-The high-level Algorithms 1 and 2 depict the main steps in recording and proof verification services.
+Ethereum’s open-source nature ensures decentralization and accessibility, allowing participation without requiring approval, which is crucial for maintaining the integrity of the recorded data.
+Additionally, Ethereum’s use of PoS as a consensus protocol enhances energy efficiency and throughput compared to traditional protocols such as the Proof of Work.
+Furthermore, Ethereum’s scalability, security and economic efficiency make it a secure option for ensuring data immutability.
+High-level Algorithms 1 and 2 depict the main steps in the recording and proof verification services.
 Algorithm 1 Accountable Information Recording
-1:topicsSet = [$selectedTopics$], $serviceCallValue$, $contractAddress$, $contractAbi$, BLOCKS_LIMIT, $privateKey$
-2:$rosbag$
-3:Boolean $recording\leftarrow$ false
-4:String $initialNonce\leftarrow$ sha256($randomBytes$)
-5:// Recording process attending to a service call.
-6:$recording\leftarrow serviceCallValue$
-7:if $recording$ = $true$ then
-8:     String $previousHash\leftarrow$$initialNonce$
-9:     for $message$ in $topicsSet$ do
-10:         if $messageNumber$ $topicRate$ == 0  then
-11:              $chainedHash\leftarrow$ sha256($message$,$previousHash$)
-12:              $previousHash\leftarrow$ $chainedHash$
-13:         end if
-14:         // Building chain of hashes.
-15:         if $payloadSize$ < BLOCKS_LIMIT then
-16:              $payload\leftarrow$ $payload\cup chainedHash$
-17:         else
-18:              // Load Solitidy contract.
-19:              loadBcContract($contractAddress$, $contractAbi$)
-20:              // Saving chain of hashes into blockchain.
-21:              $transaction\leftarrow$buildTransaction($payload$)
-22:              $signedTransaction\leftarrow$ signTransaction($transaction$, $privateKey$)
-23:              sendTransaction($signed\_transaction$)
-24:              for $hash$ in $payload$ do
-[…]
-26:              end for
-27:         end if
-28:     end for
-29:end if
-30:return $rosbag$
+0:  $topicsSet$ = [$selectedTopics$], $serviceCallValue$, $contractAddress$, $contractAbi$, BLOCKS_LIMIT, $privateKey$
+0:  $rosbag$
+1:  $recording\leftarrow$ false
+2:  $initialNonce\leftarrow$ sha256($randomBytes$)
+3:  $recording\leftarrow serviceCallValue$
+4:  $proof\leftarrow\{\}$
+5:  if $recording=true$ then
+6:     $previousHash\leftarrow initialNonce$
+7:     for $message$ in $topicsSet$ do
+8:        if $messageNumber$ mod $topicRate$ = 0 then
+9:           $chainedHash\leftarrow$ sha256($message$, $previousHash$)
+10:           $previousHash\leftarrow chainedHash$
+11:        end if
+12:        if $payloadSize$ BLOCKS_LIMIT then
+13:           $payload\leftarrow payload\cup chainedHash$
+14:        else
+15:           loadBcContract($contractAddress$, $contractAbi$)
+16:           $transaction\leftarrow$ buildTransaction($payload$)
+17:           $signedTx\leftarrow$ signTransaction($transaction$, $privateKey$)
+18:           sendTransaction($signedTx$)
+19:           for $hash$ in $payload$ do
+20:              saveBC($proof[hash]$ = BLOCK_NUMBER)
+21:           end for
+22:        end if
+23:     end for
+24:  end if
+25:  return  rosbag
 Algorithm 2 Proof Verification
-1:$hashValue$
-2:$BlockNumber$
-3:// Load Solitidy contract.
-4:loadBcContract($contractAddress$, $contractAbi$)
-5:$transaction\leftarrow$buildTransaction($hashValue$)
-6:$signedTransaction\leftarrow$ signTransaction($transaction$, $privateKey$)
-7:sendTransaction($signed\_transaction$)
-8:$blockNumber\leftarrow$readBC($hashValue$)
-9:if $blockNumber$<>0 then
-10:     $msg\leftarrow$ “The hash value is stored in $blockNumber$."
-11:else
-12:     $msg\leftarrow$ “The hash value is not stored."
-13:end if
-14:return $msg$
-Our approach addresses challenges such as the drop in robot performance during the Rosbag recording process and costs associated with storing blockchain data.
-We accomplish this by saving chained hashes from messages at specific intervals, ensuring the integrity of data while preserving system performance.
+0:  $hashValue$, $contractAddress$, $contractAbi$, $privateKey$
+0:  $msg$
+1:  loadBcContract($contractAddress$, $contractAbi$)
+2:  $transaction\leftarrow$ buildTransaction($hashValue$)
+3:  $signedTx\leftarrow$ signTransaction($transaction$, $privateKey$)
+4:  sendTransaction($signedTx$)
+5:  $blockNumber\leftarrow$ readBC($hashValue$)
+6:  if $blockNumber\neq 0$ then
+7:     $msg\leftarrow$ “The hash value is stored in block ” $+blockNumber$
+8:  else
+9:     $msg\leftarrow$ “The hash value is not stored.”
+10:  end if
+11:  return  $msg$
+Our approach addresses challenges such as the drop in robot performance during the Rosbag recording process and the costs associated with storing blockchain data.
+We accomplish this by saving chained hashes from messages at specific intervals, ensuring the integrity of the data while preserving system performance.
 These intervals define how often an integrity proof is stored and can be established based on the relevance of messages from a topic, regarding their required immutability features.
-Furthermore, message proofs are stored in real-time on the Ethereum blockchain, eliminating the need to modify the content of the Rosbag file.
-Detection of any alteration to the messages within this file is facilitated by a service that generates the hash chain from the file at intervals set for each topic during the recording process.
-To verify the existence of any of these hashes in the blockchain, the blockchain logic has been extended by developing an SC using the Solidity language.
+Furthermore, message proofs are stored in real-time in the Ethereum blockchain, eliminating the need to modify the content of the Rosbag file.
+The detection of any alteration to the messages within this file is facilitated by a service that generates a hash chain from the file at intervals set for each topic during the recording process.
+Blockchain logic has been extended by developing a SC using Solidity language to verify the existence of any of these hashes in the blockchain.
 Solidity is an open-source programming language that facilitates the creation of decentralized applications, contracts, protocols, and agreements.
-This contract provides functionalities for storing and verifying proof information for hashes stored on the Ethereum blockchain by associating each hash value with the block number where it is stored.
-Verification of transaction ownership is also conducted during the storage of these proofs.
-Therefore, this approach combines immutability, decentralization, traceability, and cost efficiency to address challenges related to the real-time recording and storage of Rosbag files securely and transparently.
-4.2 Log Curation Component
-To filter relevant information from the Rosbag file messages, a raw data processing node has been implemented.
-Given the size, variety, and intricate nature of messages contained in the Rosbag file, the inclusion of this raw information might be challenging for non-expert users.
-Therefore, it is necessary to include a component that processes this data, eliminating non-essential content in the natural language explanation generation process.
-By analyzing messages from the recorded Rosbag topics, this service can identify key circumstances in the navigation process, such as the appearance of an obstacle or a change in the planned trajectory.
-The ROS 2 NavigateToPose action server commands the robot to navigate to a specific goal.
-While Nav2 usually prioritizes the shortest path, external factors may require adjustments.
+This contract provides functionalities for storing and verifying proof information for hashes stored on the Ethereum blockchain by associating each hash value with the block number in which it is stored.
+The verification of transaction ownership is also conducted during the storage of these proofs.
+Therefore, this approach combines immutability, decentralization, traceability, and cost efficiency to securely and transparently address challenges related to the real-time recording and storage of Rosbag files.
+IV-B Log Curation Component
+To filter the relevant information from the Rosbag file messages, a raw data-processing node has been implemented.
+Given the size, variety, and intricate nature of the messages contained in the Rosbag file, the inclusion of this raw information might be challenging for non-expert users.
+To address this, a dedicated component has been developed to refine the data by discarding non-essential content before generating natural language explanations.
+Furthermore, by analyzing messages from the recorded Rosbag topics, this service can identify key circumstances in the navigation process, such as the appearance of an obstacle or a change in the planned trajectory.
+The ROS 2 NavigateToPose action server commands the robot to navigate toward a specific goal.
+Although Nav2 usually prioritizes the shortest path, external factors may require adjustments.
 The /plan topic provides the sequence of poses planned to reach the goal.
 If unexpected circumstances cause a deviation from this path, the Euclidean distance between consecutive poses in the precomputed plan will increase.
-This change, together with sensor data from the /scan topic, can indicate potential obstacles causing re-planning.
-These reasonings have been successfully checked in previous authors’ work [86] to build explanations for autonomous agents based on the use of ROS 2 topics.
-Additionally, the present approach uses the information provided by Nav2 behavior trees due to their potential in designing, visualizing, and understanding the behavior of autonomous robots.
-Other information such as the present position of the robot and its velocity is also processed through this component.
-A general overview of its behavior is described through the Algorithm 3.
-As a result, through this component, ROS 2 developers are relieved from having to call the logging API to generate explanations.
+This change, together with sensor data from the /scan topic, can indicate potential obstacles that cause re-planning.
+These reasonings have been successfully verified in previous authors’ work [80] to build explanations for autonomous agents based on the use of ROS 2 topics.
+Additionally, the present approach uses the information provided by Nav2 behavior trees because their potential for designing, visualizing, and understanding the behavior of autonomous robots.
+Other information, such as the present position of the robot and its velocity is also processed through this component.
+A general overview of its behavior is presented in Algorithm 3.
+As a result, through this component, ROS 2 developers are relieved from calling the logging API to generate explanations.
 This functionality streamlines the development process, allowing developers to focus on core functionalities without the need for managing log calls.
 This approach reduces debugging efforts and boosts efficient resource usage, scalability, and adaptability in the development lifecycle, contributing to a more productive natural language explanation generation process.
-An example of how raw ROS messages are transformed through this component is included in Figure 2.
+An example of the transformation of ROS messages through this component is shown in Figure 2.
 Algorithm 3 Log Curation Process
-1:$rosbag$, $btNodeDesc$
-2:$curationLog$
-3:for $message$ in $rosbag$ do
-4:     $msgTopic\leftarrow$ getMsgTopic($message$)
-5:     $logMsg\leftarrow message.data$
-6:     if $msgTopic$ == ’/plan’  then
-7:         $logMsg\leftarrow$ Log significant path changes and obstacles.
-8:     else if $msgTopic$ == ’/navigate_to_pose/_action/status’  then
-9:         $logMsg\leftarrow$ Track navigation status and log outcomes.
-10:     else if $msgTopic$ == ’/behavior_tree_log’  then
-11:         $logMsg\leftarrow$ Log behavior tree node status.
-12:     else if $msgTopic$ == ’/amcl_pose’  then
-13:         $logMsg\leftarrow$ Record robot position and orientation.
-14:     else if $msgTopic$ == ’/cmd_vel’  then
-15:         $logMsg\leftarrow$ Log linear and angular velocities.
+0:  $rosbag$, $btNodeDesc$
+0:  $curationLog$
+1:  for $message$ in $rosbag$ do
+2:     $msgTopic\leftarrow\text{getMsgTopic}(message)$
+3:     $logMsg\leftarrow message.data$
+[…]
+5:        $logMsg\leftarrow$ Track navigation status
+[…]
+7:        $logMsg\leftarrow$ Record log outcomes
+[…]
+9:        $logMsg\leftarrow$ Log path changes and obstacles
+[…]
+11:        $logMsg\leftarrow$ Log behavior tree node status
+[…]
+13:        $logMsg\leftarrow$ Record robot position and orientation
+[…]
+15:        $logMsg\leftarrow$ Log linear and angular velocities
 16:     end if
@@
 19:  return  $curationLog$
-{mdframed} [hidealllines=true,backgroundcolor=gray!10,innerleftmargin=3pt,innerrightmargin=3pt,leftmargin=-3pt,rightmargin=-3pt] Rosbag Raw Messages action_msgs.msg.GoalStatusArray(status_list=action_msgs.msg.GoalStatus(goal_info= action_msgs.msg.GoalInfo( goal_id=unique_identifier_msgs.msg.UUID(uuid=array([190, 81, 48, 208, 248, 147, 78, 82, 188, 116, 202, 152, 203, 225, 149, 158], dtype=uint8)), stamp=builtin_interfaces.msg.Time(sec=27, nanosec=933000000)), status=2)]) nav2_msgs.msg.BehaviorTreeStatusChange(timestamp=builtin_interfaces.msg.Time(sec=1701354686, nanosec=130104968), node_name=’ComputePathToPose’, previous_status=’IDLE’, current_status=’RUNNING’){mdframed} [hidealllines=true,backgroundcolor=gray!10,innerleftmargin=3pt,innerrightmargin=3pt,leftmargin=-3pt,rightmargin=-3pt] Processed Data 1701354686 Navigation to the goal number 1 has started.
-1701354686 Navigation to the goal number 1 is in progress.
-1701354686 Nav2 Behavior Tree node NavigateRecovery that recovers from unexpected situations, is running.
 Figure 2: Example of raw Rosbag messages and information obtained from them after the Log Curation process.
-4.3 Explainability Component
-The output of the previously described node serves as an external source in our implementation of the RAG approach.
-This development is based on the use and extension of localGPT  [^1] https://github.com/PromtEngineer/localGPT, an open-source question-answering system designed to interact with documents conversationally while safeguarding the privacy of the underlying information.
+IV-C Explainability Component
+The output of the previously described node serves as an external source in the implementation of the RAG approach.
+This development is based on the use and extension of localGPT  [^1] https://github.com/PromtEngineer/localGPT, an open-source question-answering system designed to interact conversationally with documents while safeguarding the privacy of the underlying information.
 Running locally, it ensures data privacy and secure interactions with the interpreted and recorded data, facilitating the generation of natural language explanations.
@@
 Therefore, this implementation combines the strengths of robotics frameworks, blockchain technology, and natural language processing to provide a comprehensive solution for real-time recording, secure storage, and user-friendly interpretation of robotic system data.
-It addresses challenges related to performance, storage costs, and user comprehension, making it an effective approach when understanding the behavior of autonomous robots.
-4.4 Software and Hardware Artifacts
-Our solution has been deployed in a Linux server running Ubuntu 22.04 as the operating system. The server is equipped with a 6-core 2.6 GHz CPU, 64 GB of memory, and a 4 GB GPU.
+It addresses challenges related to performance, storage costs, and user comprehension, making it an effective approach for understanding the behavior of autonomous robots.
+IV-D Software and Hardware Artifacts
+Our solution has been deployed on a Linux server running Ubuntu 22.04 as the operating system. The server is equipped with a 6-core 2.6 GHz CPU, 64 GB of memory, and a 4 GB GPU.
 The chosen ROS 2 version is Humble.
-Details regarding the topics incorporated into the Rosbag file, along with their average rates, are provided in Table 2.
-These topics play a crucial role in ensuring the reproducibility of the navigation process, facilitating data analysis and debugging purposes.
-The experimentation was executed entirely in Gazebo, an open-source solution for 3D robotics simulations.
-To provide a realistic scenario, we used a hospital simulation environment made available by Amazon Web Services [^2] https://github.com/aws-robotics/aws-robomaker-hospital-world.
+Details regarding the topics incorporated into the Rosbag file, along with their average rates, are provided in Table II.
+These topics play a crucial role in ensuring the reproducibility of the navigation process, facilitating data analysis and debugging.
+The experimentation was executed entirely in Gazebo, an open-source solution for 3D robotic simulations.
+To provide a realistic scenario, we used a hospital simulation environment made available by the Amazon Web Services [^2] https://github.com/aws-robotics/aws-robomaker-hospital-world.
 Additionally, we selected RB-1, an autonomous mobile robot based on the ROS 2 platform produced by Robotnik [81].
 This combination of tools and platforms enabled us to assess the impact and effectiveness of our approach.
-The interaction with Ethereum has been developed through the web3.py library (version 6.6.1). Ganache (version 2.7.1) has been used as the Ethereum network simulator.
+Interaction with Ethereum has been developed through the web3.py library (version 6.6.1). Ganache (version 2.7.1) has been used as the Ethereum network simulator.
 RAG solution holds Langchain (version 0.325), and ChromaDB (version 0.4.6).
@@
 This LLM integrates Reinforcement Learning from Human Feedback (RLHF) and shows versatility in handling general-purpose tasks, ranging from answering questions, text generation, and language translation to more demanding tasks such as reasoning or coding.
-Preliminary tests of Llama 2, compared to other LLMs such as Mistral [89] or Zephyr [90], led us to choose this model for evaluating our data, due to the promising results obtained in these assessments conducted during the last quarter of 2023, when this work was in its development phase.
-As a text embedding model we have chosen Instructor-xl [91], a high-performance model designed to generate text embeddings tailored to tasks such as classification, clustering, text evaluation or retrieval, a fact particularly important in the processes performed for any RAG solution.
-Table 2: Rosbag topics rate and average rate per message.
+Preliminary tests of Llama 2, compared to other LLMs such as Mistral [83] or Zephyr [84], led us to choose this model for evaluating our data due to its promising results and suitability for the available hardware during the development phase of this work.
+As a text embedding model we have chosen Instructor-xl [85], a high-performance model designed to generate text embeddings tailored to tasks such as classification, clustering, text evaluation or retrieval, which is particularly important in the processes performed for any RAG solution.
+TABLE II: Rosbag topics rate and average rate per message.
 Topic Avg. Rate (Hz) Avg. Size Msg. (B)
@@
 tf 387.278 195.413
-5 Evaluation
-To evaluate our proposal, we conducted an experimental study comparing the performance of an accountability system built in accordance with Section 4.
-This study covered several approaches for storing integrity proofs across various scenarios.
-Additionally, we explored and adapted the capabilities of RAG solutions and LLMs in question-answering systems, using as a source the data gathered by our black box-like component, through a comprehensive set of questions.
-This work aimed to assess the effectiveness of these technologies in delivering natural language explanations to individuals who are not experts.
+V Evaluation
+To evaluate our proposal, we conducted an experimental study to compare the performance of an accountability system built in accordance with Section IV.
+This study covered several approaches to storing integrity proofs across various scenarios.
+Additionally, we explored and adapted the capabilities of RAG solutions and LLMs in question-answering systems, using the data gathered by our black box-like component as a source, through a comprehensive set of questions.
+This study aimed to assess the effectiveness of these technologies in delivering natural language explanations to individuals who are not experts.
 The experiments conducted within this proposal focus on navigation tasks included in the Navigation Functionality of the European Robotics League Consumer Service Robots Challenge [86].
-This functionality evaluates the capabilities of a robot to navigate accurately and autonomously within a given scenario, containing furniture, objects, and people.
-Details like the initial position and the number of waypoints the robot needs to visit are provided beforehand.
-Furthermore, the coordinates of the waypoints are communicated sequentially to the robot during runtime.
+This functionality evaluates the capability of a robot to navigate accurately and autonomously within a given scenario, including furniture, objects, and people.
+Details, such as the initial position and the number of waypoints the robot needs to visit, are provided beforehand.
+Furthermore, the coordinates of the waypoints are communicated sequentially to the robot during the runtime.
 Each waypoint is characterized by its $X$, $Y$, and $\Theta$ coordinates.
-Existing elements in the environment, such as furniture, doors, or walls, must be mapped.
-The obstacles, with varying shapes and sizes, remain unknown to the teams beforehand and may differ between runs.
+Existing elements in the environment, such as furniture, doors, and walls, must be mapped.
+Obstacles, with varying shapes and sizes, remain unknown to the teams beforehand and may differ between runs.
 Our experiments are structured around three distinct assessment scenarios in line with the navigation functionality previously described.
-In the initial scenario, the robot must navigate to three waypoints, encountering no obstacles along the way.
+In the initial scenario, the robot must navigate three waypoints, encountering no obstacles along the way.
 The second scenario presents the robot with the challenge of adjusting its previously planned path to reach the second goal.
-This adjustment becomes necessary due to the appearance of an obstacle obstructing the original planned route, requiring a deviation to achieve the goal.
-The third scenario introduces a new obstacle that blocks the only door needed to reach the first goal.
-Consequently, this obstacle prompts the cancellation of the first waypoint objective.
-The second and third goals are achieved with the same restrictions as detailed in the second scenario.
-The three scenarios are represented in Figure 4.
-The entire source code is accessible online on GitHub [^3] https://github.com/laurafbec/immutable_explainable_BBR.git, and is licensed under GPLv3.
+This adjustment becomes necessary because the appearance of an obstacle obstructing the original planned route, which requires a deviation to achieve the goal.
+The third scenario introduces a new obstacle that blocks the only door required to reach the first goal.
+Consequently, this obstacle leads to the cancellation of the first waypoint objective.
+The second and third goals are achieved with the same restrictions as those detailed in the second scenario.
+The three scenarios are illustrated in Figure 4.
+The entire source code is accessible online on GitHub [^3] https://github.com/laurafbec/immutable_explainable_BBR.git, and licensed under the GPLv3.
 Figure 4: Scenario maps with trajectories, goals, and obstacles. (a) Scenario 1: No obstacles scenario. (b) Scenario 2: One obstacle scenario. (c) Scenario 3: Two obstacles scenario.
-5.1 Black Box Recorder Evaluation
-In evaluating the accountability component, we conducted six experiments across three previously described scenarios.
-These experiments involved storing integrity proofs on the blockchain, with hash calculations occurring at intervals of one every 10, 25, 50, or 100 messages.
+V-A Black Box Recorder Evaluation
+To evaluate the accountability component, we conducted six experiments using the three previously described scenarios.
+These experiments involved storing integrity proofs on the blockchain, with hash calculations occurring at intervals of one for every 10, 25, 50, or 100 messages.
 Moreover, our solution introduces the flexibility to save proofs of evidence at different rates based on the topic, ensuring non-tampering for critical topics or those with lower publishing rates where message loss could be crucial.
@@
 Each experiment has been repeated six times, totaling 108 runs, to derive insights into the characteristics of the black box-like device.
-5.1.1 Metrics
+V-A1 Metrics
 We employed different metrics to evaluate the impact of the accountability system in the chosen scenarios.
-First, through the topics’ publication rate and the number of messages stored per topic in each Rosbag file, we analyzed the average rate of lost messages based on the employed blockchain storage approach.
+First, using the topics’ publication rate and the number of messages stored per topic in each Rosbag file, we analyzed the average rate of lost messages based on the blockchain storage approach.
 Next, regarding system performance, we studied variables such as CPU, memory and disk usage.
 Specifically, we examined the following variables: system load 1-, 5- and 15-minute average, CPU usage (%), memory usage (GB), sent network traffic (MB/s) and disk writing rate (MB/s).
-These variables are highly significant for system-overhead analysis given that elevated load conditions can potentially influence the robot’s behavior and event-recording capabilities.
-In particular, we have obtained a full report about system load, CPU, RAM, network traffic, and disk write operations updated every second, facilitating the tracking of overall system activity and an easy comparison with preceding values.
+These variables are highly significant for system-overhead analysis given that elevated load conditions can potentially influence a robot’s behavior and event-recording capabilities.
+In particular, we obtained a full report on system load, CPU, RAM, network traffic, and disk write operations updated every second, facilitating the tracking of overall system activity and an easy comparison with preceding values.
 To describe the overall distribution and characteristics of these results, we calculated their average as a statistical measure.
-This information is included in Section 6.
-5.2 Explainability Evaluation
-Despite the remarkable performance of LLMs in various language tasks, evaluating these models presents challenges, especially in ensuring factual consistency.
+This information is presented in Section VI.
+V-B Explainability Evaluation
+Despite the remarkable performance of LLMs in various language tasks, evaluating these models presents challenges, particularly in ensuring factual consistency.
 The extensive capabilities of LLMs and the limitations of current benchmarks in capturing human preferences have led to emerging trends in LLM evaluations, such as the use of strong LLMs as judges [22].
-Traditional similarity-based metrics, such as BLEU and ROUGE, have proven to misalign with human evaluation.
-While human evaluation remains the primary standard for assessing human preferences, it is exceptionally time-consuming and expensive.
-Furthermore, this approach significantly impacts the scalability of evaluations, making it necessary to use more effective and unbiased methods [81, 93].
-To automate the evaluation process, several works have adopted the ‘LLM-as-a-judge’ approach, which focuses on using strong LLMs like GPT-4.
+Traditional similarity-based metrics, such as BLEU and ROUGE, have been proven to be misaligned with human evaluations.
+Although human evaluation remains the primary standard for assessing human preferences, it is exceptionally time-consuming and expensive.
+Furthermore, this approach significantly affects the scalability of evaluations, making it necessary to use more effective and unbiased methods [21, 87].
+To automate the evaluation process, several works have adopted the ‘LLM-as-a-judge’ approach, which focuses on using strong LLMs, such as GPT-4.
 These models are often trained with RLHF, resulting in strong human alignment [88].
-Previous studies have evidenced the consistent agreement of these models with human grading [95, 96, 97, 98].
+Previous studies have shown the consistent agreement between these models and human grading [89, 90, 91].
 Moreover, research shows that LLMs can replicate most of conclusions from human evaluations, particularly in tasks involving prediction accuracy and explanation helpfulness, even across varied settings and domains [21, 92, 93].
 However, while LLMs excel in scalability and efficiency, their strengths can be enhanced by integrating the nuanced understanding and subjective judgments that humans provide, particularly in complex or domain-specific contexts [94].
-For example, human evaluators can detect subtle inconsistencies or interpret ambiguous explanations in tasks that require deeper contextual understanding, which an automated method might overlook [48, 102].
-To address these limitations, we propose a hybrid evaluation method that combines the scalability of LLM-based evaluations with the depth of human assessments.
-In this approach, a strong LLM such as GPT-4 is used to evaluate a broad set of model outputs, ensuring cost-effective and consistent initial assessments.
-Human evaluators then validate a representative subset of these outputs to uncover overlooked biases or misalignments.
-This approach integrates the efficiency and scalability of automated methods with the reliability of human judgment, particularly for nuanced cases.
+For example, human evaluators can detect subtle inconsistencies or interpret ambiguous explanations in tasks that require deeper contextual understanding, which an automated method may overlook [20, 95].
+To leverage the strengths of both approaches, we propose a hybrid evaluation method that combines the scalability of LLM-based evaluations with the depth of human assessment.
+In this approach, a strong LLM such as GPT-4 is used to evaluate a broad set of model outputs, ensuring cost-effective and consistent initial assessment.
+Human evaluators then validated a representative subset of these outputs to uncover overlooked biases or misalignments.
+This approach integrates the efficiency and scalability of automated methods with the reliability of human judgment, particularly in nuanced cases.
 By adopting this hybrid methodology, we balance efficiency and accuracy, ensuring that the evaluation process captures both objective metrics and subjective preferences effectively.
 This dual-layered evaluation enables comprehensive assessments while addressing the inherent limitations of relying exclusively on either LLMs or humans.
-5.2.1 LLM-as-a-judge Evaluation
-The need for a robust LLM aligned with human preferences led us to select GPT-4 to evaluate the answers produced by our explainability component.
+V-B1 LLM-as-a-judge Evaluation
+The need for a robust LLM aligned with human preferences led us to select the GPT-4 model to evaluate the answers produced by our explainability component.
 Through LLM-as-a-judge, we can obtain not only scores but also explanations, making their outputs interpretable.
-The whole process is represented in Figure 5.
+The entire process is illustrated in Figure 5.
 Figure 5: Explainability evaluation.
-LLMs exhibit a certain degree of indeterminism, defined as the model’s ability to offer different answers when asked the same question or prompt several times.
-This behavior derives from their training, in which the models learn to generate outputs by sampling from probabilistic distributions over possible word sequences, introducing randomness to their response.
-The effect of this indeterminism must be considered in any LLM evaluation.
-In addition, RAG for question-answering systems is one of the most difficult LLM use cases to evaluate.
+LLMs exhibit a certain degree of indeterminism, defined as a model’s ability to offer different answers when asked the same question or prompt several times.
+This behavior is derived from their training, in which the models learn to generate outputs by sampling from probabilistic distributions over possible word sequences, thereby introducing randomness to their responses.
+The effect of indeterminism must be considered in LLM evaluations.
+In addition, the use of RAG for question-answering systems is one of the most difficult LLM use cases for evaluation.
 A basic RAG setup includes two main sources of error: the retriever and the LLM.
-The evaluation of the retriever aims to determine the relevance of its document selection concerning a query.
+The evaluation of the retriever aims to determine the relevance of document selection for a query.
 To this end, we conducted an experimental study with 16 different questions on the data generated and processed during the navigation task in the three previously described scenarios.
-This set of questions has been evaluated six times in each scenario, giving a total of 288 questions and answers to perform the analysis.
-This approach ensures a thorough examination of the performance of the RAG system in various scenarios, showing potential strengths and weaknesses in our solution.
-Using the context provided by the Rosbag data and an evaluation question, our approach includes an evaluation component for the generated explanations that contrasts the prediction made by the evaluated LLM and, thanks to the retrieved selection, compares it to an annotated ground truth.
-This process yields a correctness metric that scores whether the retriever returned no relevant documents (0) or whether it returned a perfect selection of documents (1).
-The final score, an average of all the scores in the dataset, is used to judge the retriever’s overall performance.
-A comprehensive evaluation of an LLM should consider multiple metrics that include correctness as well as other aspects.
+This set of questions has been evaluated six times for each scenario, giving a total of 288 questions and answers to perform the analysis.
+This approach ensures a thorough examination of the performance of the RAG system in various scenarios, demonstrating the potential strengths and weaknesses of our solution.
+Using the context provided by the Rosbag data and a given evaluation question, our approach includes an evaluation component that assesses the generated explanations by comparing the prediction made by the LLM against an annotated ground truth.
+The comparison yields a correctness score of 1 if the prediction matches the ground truth, and 0 otherwise.
+This binary metric reflects the extent to which retrieved information supports accurate explanations.
+The final score, computed as the average across the dataset, measures the retriever’s overall effectiveness in providing context that leads to correct, grounded responses.
+A comprehensive evaluation of an LLM should consider multiple metrics including correctness and other aspects.
 A language model should produce text that is not only correct but also logically structured and easy to understand.
-Produced responses must be appropriately detailed without unnecessary verbosity.
-Furthermore, the LLM must not only provide correct information but also deliver it in a meaningful and valuable way to the user.
-Therefore, to assess the performance of the selected LLM on specific attributes, this approach also includes a criteria-based evaluation covering accuracy, conciseness, coherence, and helpfulness.
-Selected criteria metrics are described in Table 3.
-This structured assessment offers a means to assess the model’s effectiveness in generating responses, providing a broader understanding of the model’s capabilities and its alignment with the used scenarios.
-Employing a scorecard defined for each of these features, our solution assigns scores between 0 and 10 to the responses it generates.
-The use of this scorecard facilitates a detailed and quantitative analysis of the LLM’s output, ensuring a comprehensive understanding of both its strengths and areas for improvement.
-A complete example of the answers and evaluations obtained for a selected question in the three selected scenarios is included in Figure 6.
-Table 3: Criteria-based evaluation metrics in LLM-as-a-judge evaluation.
+The responses must be appropriately detailed without unnecessary verbosity.
+Furthermore, the LLM must not only provide correct information but also deliver it in a meaningful and valuable manner to the user.
+Therefore, to assess the performance of the selected LLM on specific attributes, this approach also includes a criteria-based evaluation of accuracy, conciseness, coherence, and helpfulness.
+The selected criteria metrics are described in Table III.
+This structured assessment offers a means to assess the model’s effectiveness in generating responses, providing a broader understanding of the model’s capabilities and its alignment with the scenarios used.
+By employing a scorecard defined for each of these features, our solution assigns scores between 0 and 10 to the responses it generates.
+The use of this scorecard facilitates a detailed and quantitative analysis of LLM output, ensuring a comprehensive understanding of both its strengths and areas for improvement.
+A complete example of the answers and evaluations obtained for a selected question in the three scenarios is shown in Figure 6.
+TABLE III: Criteria-based evaluation metrics in LLM-as-a-judge evaluation.
 Metric Description
-Accuracy Accuracy evaluates the extent to which the response conforms to the reference, ranging from being completely unrelated to being completely accurate and in line with the reference.
-Helpfulness Helpfulness evaluates the degree to which the response provides valuable information, ranging from providing no helpful information to being extremely helpful and offering comprehensive information.
-Conciseness Conciseness assesses the brevity of the response, ranging from being extremely verbose to highly concise with no unnecessary words.
-Coherence Coherence examines the organizational structure and logical flow of the response, from extremely inconsistent to highly coherent, with a clear and logical organization.
-{mdframed} [hidealllines=true,backgroundcolor=blue!20,innerleftmargin=3pt,innerrightmargin=3pt,leftmargin=-3pt,rightmargin=-3pt, skipabove=-3pt,skipbelow=-10pt] Question: How many goals have been reached by the robot?
-{mdframed}[hidealllines=true,backgroundcolor=blue!10,innerleftmargin=3pt,innerrightmargin=3pt,leftmargin=-3pt,rightmargin=-3pt] Scenario 1 Answer: Based on the provided context, we know that the robot has reached 3 goals.
-The context mentions that the goal number 1, 2 and 3 have been reached successfully.
-Therefore, the answer to your question is 3.{mdframed} [hidealllines=true,backgroundcolor=gray!10,innerleftmargin=3pt,innerrightmargin=3pt,leftmargin=-3pt,rightmargin=-3pt] GPT-4 Judgement Predicted Grade: CORRECT Reasoning: The AI assistant’s response is accurate, concise, coherent, and helpful.
-It directly answers the user’s question by confirming that the robot has reached three goals.
-The assistant also provides clarification by referencing the context, which adds to the response’s accuracy and helpfulness.
-There are no unnecessary words or unclear statements, making the response concise and coherent.
-Accuracy: [[10]] Conciseness: [[10]] Coherence: [[10]] Helpfulness: [[10]]", score: 1.0{mdframed} [hidealllines=true,backgroundcolor=blue!10,innerleftmargin=3pt,innerrightmargin=3pt,leftmargin=-3pt,rightmargin=-3pt] Scenario 2 Answer: Based on the context provided, the robot has reached 3 goals.{mdframed} [hidealllines=true,backgroundcolor=gray!10,innerleftmargin=3pt,innerrightmargin=3pt,leftmargin=-3pt,rightmargin=-3pt] GPT-4 Judgement Predicted Grade: CORRECT Reasoning: The assistant’s response is completely accurate as it matches the information given in the ground truth.
-The response is concise, with no unnecessary words or information.
-It is highly coherent and has a clear structure.
-The response is extremely helpful as it gives the exact information asked for by the user.
-Therefore, the ratings are as follows: Accuracy: [[10]] Conciseness: [[10]] Coherence: [[10]] Helpfulness: [[10]]", score: 1.0{mdframed} [hidealllines=true,backgroundcolor=blue!10,innerleftmargin=3pt,innerrightmargin=3pt,leftmargin=-3pt,rightmargin=-3pt] Scenario 3 Answer: Based on the provided context,the robot has successfully reached 2 out of 3 goals.
-The successful goals are: * Goal number 2: Navigation to the goal number 2 has succeeded.
-Position: -9.14068582195674, -25.977117025927537.
-Orientation: -0.6580310307304338,0.7529908117605705.
-* Goal number 3: Navigation to the goal number 3 has succeeded.
-Position: -1.8811226231679194, -28.773672979732826.
-Orientation: -0.5946646510401259,0.803973850820613.
-The aborted goal is: * Goal number 1: Navigation to the goal number 1 has been aborted.
-Position: -5.062160931312302, -8.330878461767831.
-Orientation: -0.8760794639388502,0.4821667479872674.
-Therefore, the answer to the user’s question is: The robot has successfully reached 2 out of 3 goals.
-Goal numbers 2 and 3 have been successfully achieved, while goal number 1 has been aborted.{mdframed} [hidealllines=true,backgroundcolor=gray!10,innerleftmargin=3pt,innerrightmargin=3pt,leftmargin=-3pt,rightmargin=-3pt] GPT-4 Judgement Predicted Grade: CORRECT Reasoning: The AI assistant’s response is comprehensive and accurate.
-It correctly states that the robot has successfully reached goals number 2 and 3 while goal number 1 has been aborted.
-The assistant also provides additional details about the positions and orientations of the goals.
-While these extra details are not explicitly asked for in the user’s question, they may be useful for the user’s understanding.
-Therefore, the assistant’s response is also helpful.
-However, the response is not concise because it includes more information than necessary, leading to some redundancy in the explanation.
-The assistant’s response is coherent as it logically organizes the information about each goal.
-Accuracy: [[10]] Conciseness: [[7]] Coherence: [[10]] Helpfulness: [[10]]", score: 1.0
+Accuracy Measures how well the response aligns with the reference, ranging from completely inaccurate to fully accurate.
+Helpfulness Assesses the informational value of the response, from uninformative to highly insightful and comprehensive.
+Conciseness Evaluates the brevity and relevance of the response, from overly verbose to succinct and focused.
+Coherence Judges the logical structure and flow, from disorganized or inconsistent to well-structured and logically connected.
 Figure 6: Examples of correct answers, RAG, and criteria-based evaluations obtained across the three scenarios for the question ’How many goals have been reached by the robot?’ with LLM-as-a-judge evaluation.
-5.2.2 Human Evaluation
+V-B2 Human Evaluation
 To validate the explanation results generated by the GPT-4 model, we conducted an online study with 17 participants with diverse demographics.
@@
 In terms of gender, 62.5% were male and 37.5% were female.
-Participants’ highest level of education varied, with 37.5% holding a PhD, 37.5% having a Master’s Degree, and 25% possessing a Bachelor’s Degree.
-Regarding expertise in explainability for autonomous agents, 37.5% were beginners, 31.25% had intermediate expertise, 12.5% were advanced, and 18.75% had no experience.
-The evaluation process involved three distinct questionnaires, one corresponding to each scenario.
-Participants assessed 16 explanations per scenario, resulting in a total of 48 explanations evaluated by each individual.
-Each explanation was rated using five questions derived from the metrics outlined in Table 4, based on the Explanation Satisfaction Scale proposed by Hoffman et al.
-in [102].
-This scale represents a widely recognized framework for measuring human responses to XAI systems.
-To ensure the evaluation aligned with human preferences, we used a 5-point Likert scale.
+The participants’ highest level of education varied, with 37.5% holding a PhD, 37.5% having a master’s degree, and 25% possessing a bachelor’s degree.
+Regarding expertise in explainability of autonomous agents, 37.5% were beginners, 31.25% had intermediate expertise, 12.5% were advanced, and 18.75% had no experience.
+The evaluation process involved three questionnaires, with one corresponding to each scenario.
+The participants assessed 16 explanations per scenario, resulting in a total of 48 explanations evaluated by each individual.
+Each explanation was rated using five questions derived from the metrics outlined in Table IV, based on the Explanation Satisfaction Scale proposed by Hoffman et al.
+[95].
+This scale is a widely recognized framework for measuring human responses to XAI systems.
+To ensure that the evaluation aligned with human preferences, we used a 5-point Likert scale.
 Answers in this approach are distributed in 1 (I disagree strongly), 2 (I disagree somewhat), 3 (I’m neutral about it), 4 (I agree somewhat), and 5 (I agree strongly).
 In addition to the explanation evaluations and optional demographic information, the questionnaire provided general instructions describing the robot’s task and included a map for each scenario.
-The questionnaires were distributed via a mailing list, and participants were informed about the anonymity and research purposes of the study.
+The questionnaires were distributed via a mail list, and participants were informed about the anonymity and research purposes of the study.
 Consent was obtained upon submission of their answers.
-Participation was voluntary, which made the process cost-effective but also introduced potential drawbacks, such as longer response times and biases that may arise from factors like age, education level, or familiarity with autonomous system explainability approaches.
+Participation was voluntary, which made the process cost-effective but also introduced potential drawbacks, such as longer response times and biases that may arise from factors such as age, education level, or familiarity with autonomous system explainability approaches.
 Statistical analysis is essential in human evaluations to ensure rigor and reliability, particularly when interpreting subjective judgments and addressing variability among participants.
 Likert-scale data, which is ordinal, requires careful statistical treatment to identify significant trends and differences while avoiding misinterpretation of the results.
-By applying appropriate statistical methods, we strengthen the credibility of our conclusions and provide a robust evaluation of the LLM’s results across scenarios.
-To ensure the reliability and robustness of the evaluation, a series of statistical tests were applied.
-Reliability analysis of the scale criteria was conducted using Cronbach’s Alpha, a robust metric for assessing internal consistency and ensuring that the scale reliably measures the intended constructs, with a commonly accepted threshold for reliability being $\alpha>0.7$ [103, 104].
-To further explore the relationships among variables in the dataset, correlation analysis was performed using Pearson correlation matrices, which provide insights into the strength and direction of linear relationships between variables [105].
-For inferential analysis, a normality check was conducted using the Shapiro-Wilk test to determine whether the data met the assumptions required for parametric tests [106].
-Since the data violated these assumptions, the non-parametric Friedman test was employed [107].
-The Friedman test is particularly suited for analyzing repeated measures on the same participants, as it does not require data to follow a normal distribution [108].
+By applying appropriate statistical methods, we strengthen the credibility of our conclusions and provide a robust evaluation of LLM’s results across scenarios.
+A series of statistical tests was applied to ensure the reliability and robustness of the evaluation.
+Reliability analysis of the scale criteria was conducted using Cronbach’s Alpha, a robust metric for assessing internal consistency and ensuring that the scale reliably measures the intended constructs, with a commonly accepted threshold for reliability of $\alpha>0.7$ [96].
+To further explore the relationships among variables in the dataset, correlation analysis was performed using Pearson correlation matrices, which provide insights into the strength and direction of the linear relationships between variables [97].
+For inferential analysis, a normality check was conducted using the Shapiro-Wilk test to determine whether the data met the assumptions required for parametric tests.
+As the data violated these assumptions, the non-parametric Friedman test was employed [98].
+The Friedman test is particularly suited for analyzing repeated measures on the same participants, as it does not require data to follow a normal distribution.
 This test assessed whether significant differences existed in participant responses across the three scenarios, offering a reliable framework for understanding the impact of different conditions on the perceived quality and effectiveness of explanations.
-Finally, post-hoc analysis using the Nemenyi test was performed to examine participant behavior in specific explanations where significant differences were observed, offering a deeper understanding of the underlying patterns in their responses.
-The statistical analyses were developed through Python scripts (version 3.10.12), utilizing libraries such as pingouin (version 0.5.5) for Cronbach’s Alpha, scipy.stats (version 1.14.1) for the Shapiro-Wilk and Friedman tests, and pandas (version 2.3.3) for data manipulation.
-This approach ensured accuracy and reproducibility in the execution of the tests.
-Table 4: Questions and metrics used for human evaluation.
+Finally, a post-hoc analysis using the Nemenyi test was performed to examine participant behavior in specific explanations where significant differences were observed, offering a deeper understanding of the underlying patterns in their responses.
+Statistical analyses were performed through Python scripts (version 3.10.12), utilizing libraries such as pingouin (version 0.5.5) for Cronbach’s Alpha, scipy.stats (version 1.14.1) for the Shapiro-Wilk and Friedman tests, and pandas (version 2.3.3) for data manipulation.
+This approach ensured accuracy and reproducibility during the text execution.
+TABLE IV: Questions and metrics used for human evaluation.
 Metric Question
@@
 Completeness The explanation seems complete.
-Usefulness The explanation is useful to my goals
-5.2.3 Categories and Evaluation Questions
-To evaluate our explainability proposal in the context of ROS 2 messages interpretation, we defined a set of questions covering diverse aspects of the navigation process over the how, the what, the which, the when and the where dimensions.
+Usefulness The explanation is useful to my goals.
+V-B3 Categories and Evaluation Questions
+To evaluate our explainability proposal in the context of ROS 2 message interpretation, we defined a set of questions covering diverse aspects of the navigation process over the how, the what, the which, the when and the where dimensions.
 These questions provide insights into the different stages and components of the three scenarios.
-To structure and improve the understanding and analysis of this set of questions, we organized them into five different categories.
-These categories range from high-level overviews of navigation to detailed inquiries about trajectory planning, goal completion, and the functioning of the BT.
-The categories also delve into the causes of unexpected behavior or errors in navigation, improving the explainability of the system whenever an unexpected action occurs.
+To structure and improve the understanding and analysis of this set of questions, we organized them into five categories.
+These categories range from high-level overviews of navigation to detailed inquiries regarding trajectory planning, goal completion, and BT functioning.
+These categories also delve into the causes of unexpected behavior or errors in navigation, improving the explainability of the system whenever an unexpected action occurs.
 Through these questions, we aim to address key concerns in XAI, including understanding why decisions are made, establishing causality between actions, determining success or failure, identifying timing, and disguising the methods by which achievements are realized [99].
-The set of categories, their descriptions and the questions included in each one are described next.
+The set of categories, their descriptions and questions included in each one are described next.
 Navigation Process Overview.
-These questions provide an overview of the navigation process captured in the Rosbag file.
-Understanding what has happened and how the navigation task has proceeded provides a high-level context that sets the stage for more detailed questions in later categories.
+These questions provide an overview of the navigation process captured at the Rosbag file.
+Understanding what has occurred and how the navigation task has proceeded provides a high-level context that sets the stage for more detailed questions in later categories.
 This category comprises the following questions:
@@
 Trajectory Planning and Re-planning.
-These questions delve into the specifics of trajectory planning, a critical aspect of navigation.
+These questions delve into the specifics of trajectory planning, which is a critical aspect of navigation.
 They address whether the robot has re-planned its trajectory, the reasons behind this behavior, and whether it has encountered any obstacles.
-These details are necessary for understanding how the robot dynamically adjusts its path.
+These details are necessary to understand how a robot dynamically adjusts its path.
 The questions included in this category are:
@@
 Specifics about Goals and Locations.
-These questions examine the specifics of goals and locations.
+These questions examine the specifics of the goals and locations.
 Understanding the location of goals, the robot’s linear velocity during navigation, its initial position and orientation, and the velocity after receiving a goal contribute to a detailed understanding of the robot’s movements and performance.
-Questions included in this category are:
+The questions included in this category are:
   1. 10.
@@
   3. 12.
-    What is the initial position and orientation of the robot?
+    What are the initial position and orientation of the robot?
   4. 13.
@@
   3. 16.
-    Which Behavior Tree’s nodes were used during navigation?
-6 Experimental Results and Discussion
-This section summarizes the experiments’ results involving both the black box-like engine and the explainability component when using our solution in the previously depicted assessment scenarios.
-6.1 Message Loss Rates
-Figure 7 shows the percentage of lost messages in the Rosbag file for experiments detailed in Section 5, capturing the topics outlined in Table 2, attending to their publication rates and to the approaches selected to store integrity proofs in blockchain.
+    Which Behavior Tree nodes were used during navigation?
+VI Experimental Results and Discussion
+This section summarizes the results of the experiments involving both the black box-like engine and explainability component when using our solution in the previously depicted assessment scenarios.
+VI-A Message Loss Rates
+Figure 7 shows the percentage of lost messages in the Rosbag file for the experiments detailed in Section V, capturing the topics outlined in Table II, attending to their publication rates, and to the approaches selected to store integrity proofs in the blockchain.
 The results across the three scenarios demonstrate a consistent trend, showing uniform message loss patterns in the creation of immutable Rosbag files.
-Choosing to store a single message hash for every 100 messages results in a loss of less than 5% compared to avoiding anti-tampering measures in the Rosbag recording process, even for topics with rates exceeding 300 Hz.
-This approach stands out as the most efficient option for integrating anti-tampering techniques into our accountability solution.
-Alternative strategies, such as saving one proof every 50 messages or adjusting proof intervals based on topic rates, achieve losses below 10% for each topic.
-This is particularly noteworthy, ensuring the integrity of messages based on their significance in a given scenario or task.
-Introducing an integrity proof every 25 messages leads to an overall reduction in message recording, especially evident when topic rates surpass 100 Hz.
-Storing one hash every 10 messages could lead to losses exceeding 30% for topics with rates higher or around 100 Hz.
+Choosing to store a single message hash for every 100 messages results in a loss of less than 5% compared with avoiding anti-tampering measures in the Rosbag recording process, even for topics with rates exceeding 300 Hz.
+This approach is the most efficient option for integrating anti-tampering techniques into our accountability solution.
+Alternative strategies, such as saving one proof every 50 messages or adjusting proof intervals based on topic rates, achieve losses of less than 10% for each topic.
+This is particularly noteworthy, as it ensures the integrity of messages based on their significance in a given scenario or task.
+Introducing an integrity proof every 25 messages leads to an overall reduction in message recording, particularly when topic rates surpass 100 Hz.
+Storing one hash every 10 messages could lead to losses exceeding 30% for topics with rates higher than or around 100 Hz.
 Beyond the consistent linear trend of message loss with high topic rates, a slight deviation is observed in this approach when topics fall below 50 Hz.
@@
 Figure 7: Message loss rates in Rosbag Files under several anti-tampering techniques. Comparative visualizations for: (a) Scenario 1, (b) Scenario 2, and (c) Scenario 3.
-6.2 System Benchmarking
-The system load or, the running thread demand on the system as the average number of running and idle threads, is shown in Figure 8.
+VI-B System Benchmarking
+The system load or, running thread demand on the system as the average number of running and idle threads, is shown in Figure 8.
 The average of 1 minute is higher than the average of 5 or 15 minutes in all scenarios, showing a load increase.
 However, the values of this metric do not exceed the CPU count, which is 6 in our experimental setup, indicating no performance issues or resource saturation.
-The inclusion of runnable and uninterruptible tasks in this metric means that the load average may increase due to a disk I/O workload, not just CPU demand, giving a more general view of system overhead.
-Other more specific metrics are described next to clarify the possible existence of a bottleneck in our approach.
+The inclusion of runnable and uninterruptible tasks in this metric means that the load average may increase due to a disk I/O workload, not just CPU demand, giving a more general view of the system overhead.
+Other more specific metrics are described next to clarify the possible existence of a bottleneck in the proposed approach.
 Figure 8: Load average under several anti-tampering techniques. Comparative visualizations for: (a) Scenario 1, (b) Scenario 2, and (c) Scenario 3.
@@
 Regarding network traffic, approaches with higher process storage in the blockchain increase this metric, but do not exceed 0,20 MB/s, which is a very affordable quantity.
-Furthermore, approaches such as storing one hash every 100 messages nearly match the avoidance of storing any proof into the blockchain, indicating that including this approach does not have a remarkable impact on the system performance.
+Furthermore, approaches such as storing one hash every 100 messages nearly match the avoidance of storing any proof into the blockchain, indicating that including this approach does not have a remarkable impact on system performance.
 Disk write operations do not form a bottleneck in system performance, with all approaches differing by no more than 0,40 MB/s, which is not a very significant impact.
-In addition, all the approaches that perform integrity-proof storage on the blockchain achieve lower or similar results to the approach avoiding anti-tampering operations.
+In addition, all approaches that perform integrity-proof storage on the blockchain achieve lower or similar results to the approach that avoids anti-tampering operations.
 Figure 9 shows the results described.
@@
 These findings may encourage the adoption of our approach to ensure integrity in messages from autonomous agents.
-6.3 Correctness Evaluation Using LLM-as-a-judge
-Figure 10 shows the results of the correctness of RAG evaluation, presenting both the overall results and assessments based on predefined categories.
-The results demonstrate consistency across the three scenarios, getting a significant number of correct answers, even though there has not been a prior tuning in the ROS 2 messages for autonomous agents’ navigation tasks.
+VI-C Correctness Evaluation Using LLM-as-a-judge
+Figure 10 shows the results of the correctness of the RAG evaluation, presenting both the overall results and the assessments based on predefined categories.
+The results demonstrate consistency across the three scenarios, obtaining a significant number of correct answers, even though there has not been prior tuning in the ROS 2 messages for autonomous agents’ navigation tasks.
 The slight percentage of incorrect answers remains within an acceptable threshold, given the specific context employed and the inclusion of general questions.
@@
 Regarding individual categories, the correctness in the category ‘Navigation Process Overview’ is relatively consistent across all scenarios, indicating a good understanding of events and progress related to navigation.
-However, the broad nature of the question "What has happened in this ROS 2 log regarding navigation?" initially yielded a correctness below 65% in the third scenario.
-In this instance, the cancellation of a goal resulted in the execution and failures of Nav2 Behavior Tree nodes, not included in the other two scenarios and focused on navigation recovery.
-Therefore, the incorporation in a dataset of information related to the cancellation of a goal results in LLM answers that usually diverge from the ground truth provided, as the model occasionally prioritized BTs information, other times focused on locations or timestamps.
+However, the broad nature of the question ”What has happened in this ROS 2 log regarding navigation?” initially yielded a correctness rate of less than 65% in the third scenario.
+In this instance, the cancellation of a goal resulted in the execution and failure of Nav2 Behavior Tree nodes, which were not included in the other two scenarios and focused on navigation recovery.
+Therefore, the incorporation in a dataset of information related to the cancellation of a goal results in LLM answers that usually diverge from the ground truth provided, as the model occasionally prioritized BTs information, other times focused on locations, or timestamps.
 This variability made it difficult to align with the ground truth provided, even if the obtained answers could be considered partially accurate.
-Introducing the prompt extension "Describe the sequence of events regarding navigation goals and include information about each of them." in Question 1 significantly increased correctness in the first category, achieving results around and exceeding 75% in all three scenarios.
-This fact underscores the effectiveness of prompt engineering when tackling ambiguous or broad queries.
+Introducing the prompt extension ”Describe the sequence of events regarding navigation goals and include information about each of them.” in Question 1 significantly increased the correctness in the first category, achieving results around and exceeding 75% in all three scenarios.
+This underscores the effectiveness of prompt engineering when tackling ambiguous or broad queries.
 The results also show consistently high correctness across all scenarios concerning questions related to trajectory planning and re-planning.
-The model can determine whether the modification of a predicted trajectory has been caused by the occurrence of an obstacle or other circumstance.
-Scenario 1 and Scenario 2 also exhibit relatively high correctness in questions related to goal completion and task status.
+The model can determine whether the modification of a predicted trajectory has been caused by the occurrence of an obstacle or by other circumstance.
+Scenarios 1 and 2 also exhibit relatively high correctness in questions related to goal completion and task status.
 However, the model encounters challenges in Scenario 3, occasionally failing to identify whether a navigation that starts correctly concludes with an abortion or cancellation.
-This issue may be attributed to misalignment during the retrieval process, specifically in the selection of the best chunk associated with a particular goal associated with numerous embeddings within a single collection from the vector database.
-Correctness in identifying specific targets and locations lacks some consistency, with results exceeding 50% in the first scenario, around 70% in the second and over 75% in the third.
-In particular, three scenarios encounter difficulties in identifying the initial position of the robot before starting the navigation, as the model misaligns the queried position with the one assumed by the robot when starting the navigation process.
-This fact may also be attributed to deficiencies in the retrieval process.
-In Scenario 1, understanding the sequence of events or the order of goals is a challenge for the model, leading to inaccuracies in specifying the second location or goal.
+This issue may be attributed to a misalignment during the retrieval process, specifically in the selection of the best chunk associated with a particular goal associated with numerous embeddings within a single collection from the vector database.
+Correctness in identifying specific targets and locations lacks consistency, with results exceeding 50% in the first scenario, around 70% in the second and over 75% in the third.
+In particular, three scenarios encounter difficulties in identifying the initial position of the robot before starting the navigation, as the model misaligns the queried position with that assumed by the robot when starting the navigation process.
+This may also be attributed to deficiencies in the retrieval process.
+In Scenario 1, understanding the sequence of events or order of goals is a challenge for the model, leading to inaccuracies in specifying the second location or goal.
 Nevertheless, this challenge is not present in Scenario 3 and is absent or reduced in Scenario 2, demonstrating how differences in data distribution can introduce varied responses even when using specific and unambiguous questions.
@@
 Particularly, the mistakes stem from the model’s confusion in determining whether a previously failed Nav2 BT node has recovered or finished its execution, as well as identifying all nodes that were executed during navigation.
-This can be attributed to the fact that the complexities of the recovery actions and the dynamic nature of BT executions during navigation recovery introduce complexities that the model may not fully generalize from the training data, underscoring the need for additional fine-tuning in such scenario-specific contexts in case the obtained correctness needs to be improved.
-6.4 Criteria-based Evaluation Using LLM-as-a-judge
+This can be attributed to the fact that the nuanced behavior of recovery actions and the dynamic execution of BTs during navigation introduce difficulties that the model may not adequately capture from the training data, underscoring the need for additional fine-tuning in such scenario-specific contexts in case the obtained correctness needs to be improved.
+VI-D Criteria-based Evaluation Using LLM-as-a-judge
 Figure 11 shows the performance of the selected LLM in the evaluated metrics, demonstrating consistency with slight variations in the scores.
@@
 Figure 11: (a) Overall criteria evaluation in the three scenarios.
-The language model’s adaptability to challenges like obstacles and goal cancellations regarding the described categories, is shown in Figure 12.
+The adaptability of the language model to challenges such as obstacles and goal cancellations regarding the described categories, is shown in Figure 12.
 Figure 12: LLM criteria-based evaluation by categories. Comparative visualizations for: (a) Scenario 1, (b) Scenario 2, and (c) Scenario 3.
 Concerning general questions about events and the progression of the navigation task, we observe variations in scores across Scenarios 2 and 3.
-While coherence and conciseness improve, Scenario 1, which involves direct navigation to three goals, scores higher in accuracy and helpfulness.
+While coherence and conciseness improve, Scenario 1, which involves direct navigation to the three goals, scores higher in accuracy and helpfulness.
 This fact suggests that effective communication on general queries benefits from the absence of obstacles or goal cancellations, enhancing overall understanding.
 In the category ‘Trajectory Planning and Re-planning’, questions are effectively addressed across all three scenarios, showing scores above 8 in every metric, even when obstacles or goal cancellations are introduced.
-These results highlight the system’s capability to manage trajectory-related queries effectively despite challenges.
-Regarding the category ‘Goal Completion and Navigation Task Status’, Scenarios 1 and 2 effectively convey information about goal completion and navigation task progress.
-Nevertheless, slightly lower scores in Scenario 3 indicate challenges in effectively communicating completion status when a goal is aborted.
+These results highlight the system’s capability to manage trajectory-related queries effectively despite these challenges.
+Regarding the category ‘Goal Completion and Navigation Task Status’ category, Scenarios 1 and 2 effectively convey information about goal completion and navigation task progress.
+Nevertheless, the slightly lower scores in Scenario 3 indicate challenges in effectively communicating the completion status when a goal is aborted.
 In the ‘Specifics about Goals and Locations’ category, the system provides valuable goal and location information, even in the presence of obstacles or goal cancellations.
@@
 The results regarding questions included in ‘Nav2 Behavior Tree and Node Status’ category demonstrate effectiveness in communication across all scenarios and metrics.
-This emphasizes the potential for BTs to complement language models in explainability, offering insight into the decision-making processes during navigation tasks.
-In summary, although results can be improved in accuracy for general questions, the language model shows adaptability across different navigation scenarios, including those with obstacles and task changes.
-The consistency in conveying information related to Nav2 BTs suggests a solid understanding of underlying navigation processes.
+This emphasizes the potential of BTs to complement language models in explainability, offering insight into decision-making processes during navigation tasks.
+In summary, although the results can be improved in accuracy for general questions, the language model shows adaptability across different navigation scenarios, including those with obstacles and task changes.
+Consistency in conveying information related to Nav2 BTs suggests a solid understanding of underlying navigation processes.
 The model also performs well when providing information about trajectory planning, goal completion, and specifics about goals and locations, regardless of the specific challenges introduced in each scenario.
-Nevertheless, while evaluation using LLMs excels in scalability and efficient resource usage, it is still recommended to test a subset of questions with human evaluators.
-This ensures a deeper understanding of how well the approach aligns with human expectations and effectively communicates, particularly in scenarios involving complex or nuanced details.
-6.5 Human Evaluation of LLM-generated Explanations
-This section presents an analysis of the results from the human evaluation conducted with 17 participants, as stated in Section 5.2.2.
-6.5.1 Distribution Analysis
+Nevertheless, although evaluation using LLMs excels in scalability and efficient resource usage, it is still recommended to test a subset of questions with human evaluators.
+This ensures a deeper understanding of how well the approach aligns with human expectations and communicates effectively, particularly in scenarios involving complex or nuanced details.
+VI-E Human Evaluation of LLM-generated Explanations
+This section presents an analysis of the results of the human evaluation conducted with 17 participants, as stated in Section V-B2.
+VI-E1 Distribution Analysis
 Figure 13 depicts the distribution of Likert scale responses for the five evaluation criteria across the three scenarios, showing the frequency of each response value.
-The highest ratings predominate across all scenarios and criteria, suggesting that participants generally found the explanations satisfactory.
-However, Scenario 2 and Scenario 3 exhibit greater variability, with an increase in mid-range scores for certain criteria, such as Usefulness.
+The highest ratings predominate across all scenarios and criteria, suggesting that participants generally found the explanations to be satisfactory.
+However, Scenarios 2 and 3 exhibit greater variability, with an increase in mid-range scores for certain criteria, such as Usefulness.
 These trends are particularly noticeable in broad questions, such as those in Category 1 regarding the navigation process overview, or when the model’s answers do not align with the ground truth.
-In such cases, participant opinions diverge significantly, even in the presence of factual inaccuracies, indicating a potential bias in how the utility is perceived when discrepancies with the ground truth arise.
+In such cases, participant opinions diverge significantly, even in the presence of factual inaccuracies, indicating a potential bias in how utility is perceived when discrepancies with the ground truth arise.
 Figure 13: Distribution of Likert scale responses across scenarios and evaluation criteria.
-6.5.2 Reliability Analysis
+VI-E2 Reliability Analysis
 The reliability analysis resulted in an overall Cronbach’s Alpha value of 0.9792, indicating good internal consistency and confirming that the evaluation criteria reliably capture the intended dimensions of the robot’s explanations.
 Additionally, we calculated Cronbach’s Alpha if an item is deleted for each criterion.
-This metric helps assess the impact of removing an individual item on the scale’s overall reliability.
-As summarized in Table 5, the "Alpha if Deleted" values are consistently close to the overall Cronbach’s Alpha.
-Specifically, these values range from 0.9699 (Usefulness) to 0.9798 (Informativeness), with all values slightly lower than the overall result.
-This suggests that removing any single item would only have a minimal impact on the reliability of the scale, indicating that each criterion contributes positively to the measurement of overall quality.
-Further, we examined the item-total correlations, which quantify the strength of the relationship between each criterion and the total score.
+This metric helps to assess the impact of removing an individual item on the scale’s overall reliability.
+As summarized in Table V, the ”Alpha if Deleted” values are consistently close to the overall Cronbach’s Alpha.
+Specifically, these values range from 0.9699 (Usefulness) to 0.9798 (Informativeness), with all values being slightly lower than the overall result.
+This suggests that removing any single item would have only a minimal impact on the reliability of the scale, indicating that each criterion contributes positively to the measurement of overall quality.
+Further, we examined item-total correlations, which quantify the strength of the relationship between each criterion and the total score.
 These correlations range from 0.9286 for the Informativeness criteria to 0.9660 for Usefulness.
 The high item-total correlations support the relevance of each item in assessing the key dimensions of the robot’s explanations.
-Table 5: Cronbach’s Alpha and item-wise "Alpha if Deleted" in human evaluation.
+TABLE V: Cronbach’s Alpha and item-wise ”Alpha if Deleted” in human evaluation.
 Criteria Cronbach’s Alpha if Deleted Item-Total Correlation
@@
 Usefulness 0.9699 0.9660
-Finally, we assessed the inter-item correlations, which provide insight into how strongly the criteria correlate with one another.
-The inter-item correlation matrix, presented in Table 6, shows values ranging from 0.878 (between Informativeness and Satisfaction) to 0.981 (between Satisfaction and Usefulness).
+Finally, we assessed inter-item correlations, which provide insight into how strongly the criteria correlate with one another.
+The inter-item correlation matrix, presented in Table VI, shows values ranging from 0.878 (between Informativeness and Satisfaction) to 0.981 (between Satisfaction and Usefulness).
 These high correlations indicate that the criteria are closely related but not redundant, as each captures a unique and complementary aspect of the evaluation process.
-Table 6: Inter-Item correlation matrix in human evaluation.
+TABLE VI: Inter-Item correlation matrix in human evaluation.
 Completeness Informativeness Satisfaction Understandability Usefulness
@@
 Usefulness 0.911 0.881 0.981 0.947 1.000
-6.5.3 Correlation Analysis
-Building on the findings from the reliability analysis, we further explored the relationships among the variables in our dataset through Pearson correlation matrices.
-While Cronbach’s Alpha results demonstrated the internal consistency of the evaluation criteria, the correlation analysis extends this by examining how these criteria relate to each other across different scenarios and explanation categories.
+VI-E3 Correlation Analysis
+Building on the findings of the reliability analysis, we further explored the relationships among the variables in our dataset using Pearson correlation matrices.
+While Cronbach’s Alpha results demonstrated the internal consistency of the evaluation criteria, the correlation analysis extends it by examining how these criteria relate to each other across different scenarios and explanation categories.
 Figure 14 presents the Pearson correlation matrix for participants’ responses across the three scenarios.
-It reveals strong associations, indicating a high level of consistency in participants’ assessments.
+This reveals strong associations, indicating a high level of consistency in the participants’ assessments.
 This consistency suggests that the explanations provided were clear and easily comparable, allowing participants to reliably evaluate them regardless of the specific context of the robot’s task.
-Figures 15 and 16 depict the relationships among evaluation criteria and question categories.
-Regarding the criteria, Completeness and Informativeness exhibit a strong positive correlation, as do Understandability and Usefulness.
-This indicates that participants are more likely to perceive explanations as useful when they are also clear.
+Figure 15 depicts the relationships among the evaluation criteria.
+Completeness and Informativeness exhibit a strong positive correlation, as do Understandability and Usefulness.
+This indicates that participants are more likely to perceive explanations as useful when they are clear.
 In contrast, the correlation between Informativeness and Usefulness is weaker (0.55), highlighting a divergence in participants’ evaluations.
-This divergence may stem from biases previously discussed, where participants find it challenging to judge the usefulness of an explanation that is not fully accurate.
-Regarding the question categories, the ‘Navigation Process Overview’ and ‘Specifics About Goals and Locations’ categories exhibit a strong correlation (0.99), indicating that participants tended to provide similar ratings for explanations in these areas.
-This suggests that these types of explanations are perceived similarly.
-In contrast, the ‘Trajectory Planning and Re-planning’ category shows weak or negative correlations with other categories.
-This discrepancy arises from differences in evaluating broad, open-ended explanations versus more specific ones.
+This divergence may stem from the biases previously discussed, where participants find it challenging to judge the usefulness of an explanation that is not fully accurate.
 Figure 14: Correlation matrix of responses by scenario.
 Figure 15: Correlation matrix of responses by criteria.
-Figure 16: Correlation matrix of responses by category.
-6.5.4 Inferential Statistics
-Statistical analysis often requires understanding the underlying data distribution to ensure the appropriate methods are applied.
-Since parametric tests assume normality, violations of this assumption can compromise the validity of the results.
+VI-E4 Inferential Statistics
+Statistical analysis often requires an understanding of the underlying data distribution to ensure that appropriate methods are applied.
+As parametric tests assume normality, violations of this assumption can compromise the validity of the results.
 To assess normality, we applied the Shapiro-Wilk test, a robust method for small to moderate-sized samples, across various groupings, including scenarios, categories, questions, and criteria.
-The test results consistently revealed significant deviations from normality, with $p$-values well below 0.05, leading to the rejection of the null hypothesis of normality for all groups.
+The test results consistently revealed significant deviations from normality, with $p$-values well below 0.05, leading to rejection of the null hypothesis of normality for all groups.
 These findings confirm that the dataset does not satisfy the assumptions required for parametric tests, such as ANOVA or t-tests.
 As a result, non-parametric methods were adopted in subsequent analyses to ensure the robustness and validity of statistical inferences.
-The Friedman test was used due to its ability to assess differences in related groups when the data does not meet the assumptions of normality required for parametric tests.
+The Friedman test was used because of its ability to assess differences in related groups when the data does not meet the assumptions of normality required for the parametric tests.
 The results of the Friedman test across the three scenarios reveal significant differences in participant ratings for several explanations, indicating varying levels of agreement across conditions.
-In Scenario 1, significant differences were observed for explanations to Questions 1 and 2, suggesting diverse responses.
-In contrast, most other questions in this scenario showed no significant differences ($p\geq 0.05$).
-Explanations 1 and 2 in all three scenarios addressed broad questions, whose answers diverged from the ground truth, as discussed in the GPT-4 evaluation in 6.3.
+In Scenario 1, significant differences were observed for the explanations of Questions 1 and 2, suggesting diverse responses.
+In contrast, most of the other questions in this scenario showed no significant differences ($p\geq 0.05$).
+Explanations 1 and 2 in all three scenarios addressed broad questions, whose answers diverged from the ground truth, as discussed in the GPT-4 evaluation in VI-C.
 Broad questions often result in responses that prioritize varied information, such as BTs, locations, or timestamps, making it difficult to align the answers with the ground truth.
-This fact justifies the significant variability in participant ratings for these explanations.
-In Scenario 2, significant differences were observed for Explanations 1, 7, 8, and 15, with $p$-values below the 0.05 threshold, while other questions did not show significant variations.
-Specifically, the explanations for Questions 7 and 8 were misaligned with the ground truth, leading to significant variability in participant assessments.
-In these cases, participants’ ratings lacked a clear trend across the evaluated criteria.
+This fact justifies the significant variability in the participant ratings for these explanations.
+In Scenario 2, significant differences were observed for Explanations 1, 7, 8, and 15, with $p$-values below the threshold of 0.05, while the other questions did not show significant variations.
+Specifically, the explanations for Questions 7 and 8 were misaligned with the ground truth, leading to significant variability in the participant assessments.
+In these cases, the participants’ ratings lacked a clear trend across the evaluated criteria.
 A similar pattern was observed in Explanation 15 of Scenario 2, where the model provided irrelevant information that did not directly address the user’s inquiry.
-This lack of relevance and alignment likely contributed to the observed variability in participant ratings.
-Finally, Scenario 3 showed significant differences for Explanations 1, 2, 7, and 16, indicating that the conditions for these explanations led to notable variations in participant ratings.
+This lack of relevance and alignment likely contributed to the observed variability in the participant ratings.
+Finally, Scenario 3 showed significant differences for Explanations 1, 2, 7, and 16, indicating that the conditions for these explanations led to notable variations in the participant ratings.
 Explanation 7, does not address the answer to the corresponding question, including no relevant information.
@@
 Figure 16: Friedman test results.
-The post-hoc Nemenyi test was conducted for explanations that showed significant differences in participant ratings based on the Friedman test results.
+A post-hoc Nemenyi test was conducted for explanations that showed significant differences in participant ratings based on the Friedman test results.
 This allowed us to identify specific participant pairs with notable divergences in their responses across the three scenarios.
-The Nemenyi test revealed that most participants were consistent in their evaluations, but certain pairs exhibited significant variability.
+The Nemenyi test revealed that most participants were consistent in their evaluations however, certain pairs exhibited significant variability.
 For example, in Scenario 1, the evaluation for Explanations 1 and 2 highlighted differences involving Participant 8 and several others.
-In Scenario 2, significant differences emerged for Explanation 1 among Participants 3, 6, 7, and 13.
-Similarly, in Scenario 3, variability was observed for Explanations 1 and 16, where selected participants differed significantly from others.
-These discrepancies are likely due to natural human biases, differences in interpretation, and subjective judgments that are inherent in human responses.
+In Scenario 2, significant differences emerged for Explanation 1 between Participants 3, 6, 7, and 13.
+Similarly, in Scenario 3, variability was observed for Explanations 1 and 16, where the selected participants differed significantly from the others.
+These discrepancies are likely due to natural human biases, differences in interpretation, and subjective judgments inherent in human responses.
 Overall, while the explanations were generally clear and consistent for most participants, the observed variability underscores the role of individual perspectives and biases in shaping evaluations, reflecting the inherent complexity of human judgment.
-6.6 Similarity Between Human and LLM Evaluations
-As described in 5.2, previous studies have demonstrated that LLMs can effectively approximate human judgment in various evaluation tasks, including those involving subjective interpretation.
-In this work, we conducted a focused validation of LLM-generated evaluations by comparing a subset of 48 explanations from one of six evaluation runs performed by the LLM against human assessments.
-This subset included 16 explanations per scenario, allowing us to validate the LLM’s judgment within a practical scope.
-Figures 18 and 19 depict the mean responses from 17 human participants and GPT-4, respectively, acting as a judge in this subset.
+VI-F Similarity Between Human and LLM Evaluations
+As described in V-B, previous studies have demonstrated that LLMs can effectively approximate human judgment in various evaluation tasks, including those involving subjective interpretation.
+In this study, we conducted a focused validation of LLM-generated evaluations by comparing a subset of 48 explanations from one of the six evaluation runs performed by the LLM against human assessments.
+This subset included 16 explanations per scenario, allowing us to validate the LLM judgment within a practical scope.
+Figures 17(a) and 17(b) depict the mean responses from the 17 human participants and GPT-4, respectively, acting as a judge in this subset.
 The comparison highlights notable similarities in scoring patterns across scenarios and explanations, suggesting a high degree of alignment between human and LLM evaluations.
-Both evaluations consistently identify high-quality explanations with scores clustered around 4.5–5, indicating shared agreement on the quality of certain explanations.
-Similarly, lower-rated explanations identified by humans are also flagged by the LLM, revealing a mutual recognition of weaker content.
-However, while the overall trends are similar, there are some differences in the exact ratings and outlier values between the two evaluations.
-These differences could be attributed to nuances in how the LLM interprets the explanations compared to human evaluators, who might take into account subtle contextual or subjective factors that an LLM might miss.
+Both evaluations consistently identify high-quality explanations with scores clustered around 4.5–5, indicating a shared agreement on the quality of certain explanations.
+Similarly, lower-rated explanations identified by humans are also flagged by the LLM, revealing mutual recognition of weaker content.
+However, although the overall trends are similar, there are some differences in the exact ratings and outlier values between the two evaluations.
+These differences could be attributed to nuances in how the LLM interprets the explanations compared to human evaluators, who might consider subtle contextual or subjective factors that an LLM might miss.
 Despite these slight variations, the visual correspondence between the human and LLM heatmaps indicates that the LLM provides a reliable approximation of human judgment, making it a promising tool for scaling evaluation processes without sacrificing much in terms of accuracy.
 This similarity could be valuable in demonstrating that LLMs are suitable alternatives for tasks like explanation evaluation, especially in large-scale studies where human evaluation may be impractical.
-Figure 18: Heatmap of mean responses from human evaluators Across Scenarios and Explanations.
-Figure 19: Heatmap of mean responses from the LLM-as-Judge evaluation across scenarios and explanations.
-7 Conclusions
-This work presents an accountability and explicability architecture, implemented for Robot Operating System (ROS)-based mobile robots.
-Our approach comprises a black box component for accountability that effectively addresses real-time recording challenges, ensuring message integrity and verifiability through blockchain technology, and an explainability component.
-This last component is based on the use of Retrieval Augmented Generation (RAG) and Large Language Models (LLMs) through a private question-answering system designed to use the information recorded by the first component as context.
-This solution demonstrates the feasibility of obtaining accurate, coherent, concise and helpful explanations for robot actions from raw data stored in Rosbag files, providing satisfactory results for users and thereby enhancing human-robot interaction.
-Additionally, both main components have been integrated through the development of a third main functionality in charge of filtering non-relevant information and making messages closer to the final user, due to their specific nature can introduce challenges in their comprehension process.
-Furthermore, this last component releases software developers from making direct calls to the logging Application Programming Interface (API), easing the development process and improving the scalability and adaptability of our solution for every navigation task performed by a ROS-based robot.
-The results achieved highlight the capability of our solution to guarantee the immutability of information generated by an agent without compromising its performance, a feature particularly important when trying to identify the causes behind an event.
-Moreover, our system improves efficiency and effectiveness in the deployment of autonomous agents, providing natural language explanations across key navigation aspects including goal achievement, obstacle appearance, trajectory re-planning, and objective cancellation or abortion.
-This fact strengthens trust and confidence in the use of autonomous agents across diverse applications and contexts.
-While our approach has shown promising results in navigation functionalities regarding several scenarios and circumstances, future work should involve adapting our solution to other applications by simply adapting the recording and the processing components to the relevant topics of the area or domain of use.
-Furthermore, the consistency and robustness of responses generated by our explainability approach could benefit from refinement across different scenarios, especially when categorizing questions under specific subjects.
-Therefore, future work should explore the impact of In-Context Learning techniques on prompt generation, in addition to fine-tuning the model for the specific context of autonomous agent navigation tasks.
-Additionally, our RAG approach could improve in terms of retrieving the most relevant chunk of information from the context.
-This enhancement may involve exploring the use of separate collections within the vector database or employing a hybrid approach, such as integrating Knowledge Graphs with vector indexes.
-Furthermore, by incorporating real-time explanation generation mechanisms into our solution, the responsiveness and adaptability of the system could be improved.
-These approaches can help to improve the accuracy, quality and overall consistency of the answers obtained across several scenarios, thereby enhancing the performance of our proposed system.
+(a) Human evaluation.
+(b) LLM-as-Judge evaluation.
+Figure 17: Heatmaps of mean responses across scenarios and explanations. (a) Human evaluators. (b) LLM-as-Judge evaluation.
+VII Conclusions
+This work presents an accountability and explainability architecture for Robot Operating System (ROS)-based mobile robots.
+The proposed system comprises a black box component that ensures the verifiability and integrity of logged messages using blockchain technology, and an explainability component based on Retrieval-Augmented Generation (RAG) and Large Language Models (LLMs).
+This private question-answering system leverages the data recorded by the accountability component to provide context for generating natural language explanations.
+The results demonstrate the feasibility of producing accurate, coherent, concise, and helpful explanations from raw Rosbag data, thereby supporting user understanding and enhancing human-robot interactions.
+Both components are integrated through a third main module in charge of filtering non-relevant information and transforming technical content into more accessible messages.
+This functionality reduces the complexity for developers by eliminating the need for direct interaction with the logging Application Programming Interface, thereby easing the development process and improving the adaptability and scalability of the system across diverse navigation tasks.
+Our evaluation confirms that the system guarantees data immutability without compromising performance, supporting post-hoc analysis and failure investigation.
+Specifically, hashing one of every 100 messages resulted in less than 5% message loss, even for high-frequency topics exceeding 300 Hz, with minimal resource consumption.
+These results validate the efficiency and practicality of accountability mechanisms.
+The system also provides effective explanations across key aspects of autonomous navigation, such as goal completion, obstacle detection, trajectory re-planning, and task cancellation or abortion.
+The generated explanations achieved high scores for accuracy, coherence, and user satisfaction.
+LLM-as-a-judge evaluations reported correctness rates at or above 75% across all scenarios, while human evaluations showed strong agreement, with a Cronbach’s Alpha of 0.9792, indicating high internal consistency.
+The alignment between human and LLM judgments suggests that LLMs can serve as reliable and scalable alternatives to human evaluation for assessing explanation quality.
+These findings highlight the potential of the proposed architecture to support the broader deployment of autonomous agents in safety-critical scenarios.
+By enabling reliable post-hoc inspection and producing accessible, high-quality explanations, our system bridges the gap between complex robotic behavior and human interpretation.
+This capability is particularly relevant in contexts that require transparency, positioning our approach as a practical foundation for responsible and explainable robots.
+While our solution has shown promising results in navigation functionalities across various scenarios, future work should extend its application to other domains by tailoring the recording and processing components to specific topics relevant to each context.
+Furthermore, the consistency and robustness of the responses generated by our explainability approach could benefit from refinement across different scenarios, especially when categorizing questions under specific subjects.
+To this end, future efforts should explore the impact of In-Context Learning techniques on prompt formulation, along with fine-tuning the model to better align with the unique demands of autonomous agent tasks.
+Additionally, our RAG approach could be further optimized to retrieve more relevant information from the available context.
+This improvement may involve structuring the vector database into separate collections tailored to specific information types or adopting a hybrid retrieval strategy that combines Knowledge Graphs with vector-based indexing.
+Incorporating real-time explanation generation into the architecture could also enhance the responsiveness and adaptability of the system.
+Together, these enhancements have the potential to significantly improve the accuracy, relevance, and consistency of the generated explanations across diverse scenarios, thereby strengthening the overall effectiveness of the proposed architecture.
 Appendix A Supplementary Materials
-A.1 Prompt Template
-[…]
+A-A Prompt Template
 Figure 18: Prompt template used to generate explanations with Llama 2 LLM.
-A.2 Explanations Obtained
+A-B Explanations Obtained
 The 288 questions and answers obtained through the explanation generation and evaluation process are available in the aforementioned mentioned Github repository  [^4] https://github.com/laurafbec/immutable_explainable_BBR/blob/main/
@@
 .
-Additionally, a Python notebook has been developed to interact with the content of the previous sheet.
-This notebook at the following link  [^5] https://github.com/laurafbec/immutable_explainable_BBR/blob/main/
+A Python notebook has been developed to interact with the content of the previous sheet.
+This notebook is available at the following link  [^5] https://github.com/laurafbec/immutable_explainable_BBR/blob/main/
 gpt4_eval_results/GPT4 Evaluation Results.ipynb
-[…]
-Declaration of Competing Interest
-The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
-Data Availability
-Data and tools are available in a public repository  [^6] https://github.com/laurafbec/immutable_explainable_BBR.
-Acknowledgments
+Acknowledgment
 This publication is part of the TESCAC project, financed “by European Union NextGeneration-EU, the Recovery Plan, Transformation and Resilience, through INCIBE”.